### v1

# Learning-based Multi-View Stereo: A Survey

Abstract
3D reconstruction aims to recover the dense 3D structure of a scene. It plays an essential role in various applications such as Augmented/Virtual Reality (AR/VR), autonomous driving and robotics. Leveraging multiple views of a scene captured from different viewpoints, Multi-View Stereo (MVS) algorithms synthesize a comprehensive 3D representation, enabling precise reconstruction in complex environments. Due to its efficiency and effectiveness, MVS has become a pivotal method for image-based 3D reconstruction. Recently, with the success of deep learning, many learning-based MVS methods have been proposed, achieving impressive performance against traditional methods. We categorize these learning-based methods as: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based, and large feed-forward methods. Among these, we focus significantly on depth map-based methods, which are the main family of MVS due to their conciseness, flexibility and scalability. In this survey, we provide a comprehensive review of the literature at the time of this writing. We investigate these learning-based methods, summarize their performances on popular benchmarks, and discuss promising future research directions in this area.
Index Terms:
Multi-View Stereo, 3D Reconstruction, Deep Learning.
3D reconstruction describes the general task of recovering the 3D structure of a scene. It is widely applied in augmented/virtual reality (AR/VR), autonomous driving, and robotics [1]. The advancement of 3D acquisition techniques has led to the increased affordability and reliability of depth sensors, such as depth cameras and LiDARs. These sensors are extensively utilized for real-time tasks, enabling rough estimations of the surrounding environment, such as simultaneous localization and mapping (SLAM)  [2, 3, 4, 5] or dense reconstruction  [6, 7, 8]. Nevertheless, depth maps captured by such sensors tend to be partial and sparse, resulting in incomplete 3D representations with limited geometric details. In addition, these sensors are active and usually consume a lot of power. In contrast, camera-based solutions, commonly found in edge devices like smartphones and AR/VR headsets, offer a more economically viable alternative for 3D reconstruction.
One fundamental technique in image-based 3D reconstruction is Multi-View Stereo (MVS). Given a set of calibrated images, MVS aims to reconstruct the dense 3D geometry for an observed scene. Based on the scene representations, traditional MVS methods can be divided into three categories: volumetric, point cloud, and depth map. Volumetric methods [9, 10, 11, 12] discretize the continuous 3D space into voxels and label each as inside or outside of the surface. They are limited to scenes of small scale because of the large memory consumption. Point cloud methods [13, 14] operate directly on 3D points and often employ propagation to gradually densify the reconstruction. As the propagation of point clouds is proceeded sequentially, these methods are difficult to be parallelized and thus take much time in processing [15]. In addition, the irregularity and large size of point cloud are also not very suitable for deep learning, especially in large-scale scenes. In contrast, methods relying on depth maps [16, 17, 18, 19, 20] use patch matching with photometric consistency to estimate the depth maps for individual images. Subsequently, these depth maps are fused into a dense representation, e.g., point cloud or mesh, as a post-processing step. Such design decouples the reconstruction problem into per-view depth estimation and depth fusion, which explicitly improves the flexibility and scalability. Although MVS has been studied extensively for several decades, traditional MVS methods rely on hand-crafted matching metrics and thus encounter challenges in handling various conditions, e.g., illumination changes, low-textured areas, and non-Lambertian surfaces [21, 22, 23, 24].
To overcome these challenges, recent works [15, 25] have turned to learning-based methods, leveraging the emergence of Convolutional Neural Networks (CNNs) that have shown remarkable success in various 2D vision tasks. These learning-based MVS techniques have achieved remarkable success and explicitly outperform traditional methods [18, 17] on various benchmarks [21, 22, 23, 26, 27].
In this survey, we categorize existing learning-based MVS methods based on their characteristics as follows: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. Voxel-based methods estimate the geometry with volumetric representation and implicit function, e.g., Signed Distance Functions (SDF). They are limited to small-scale scenes due to the high memory consumption. NeRF and 3D Gaussian Splatting-based methods adapt NeRF [28] and 3D Gaussian Splatting [29], which are originally used for novel view synthesis, to extract surface from the implicit field or point cloud. They typically need to optimize the geometry for each new scene, which needs lots of run-time and memory consumption. Large feed-forward methods typically use large transformer models to directly learn the 3D representation from given images. They require massive computation because of the huge network, and are mainly limited to object-level scenes [30]. Depth map-based methods utilize deep learning in depth estimation and then fuse depth maps with traditional fusion algorithm. Comparatively, depth map-based methods [15, 25] are the main family of learning-based methods since they inherit the advantages from those traditional methods [17, 18, 19, 20] by decoupling 3D reconstruction into depth estimation and depth fusion. Therefore, we focus more on the depth map-based methods and discuss them in more details in this survey.
For clarity, we further categorize depth map-based methods into online and offline methods, shown in Fig. 1. Specifically, online MVS methods [25, 31] usually perform multi-view depth estimation with a video sequence of low-resolution images, e.g., ScanNet [26]. They typically use small resolution inputs and a lightweight network structure to ensure quick inference considering speed and simplicity. The primary focus of online MVS is to provide instant reconstruction for applications where time efficiency is critical, e.g., augmented reality and live video processing. In contrast, offline MVS methods [15, 32] prioritize better quality reconstruction at the cost of computation. These methods excel in multi-view depth estimation and point cloud fusion using high-resolution image sets, e.g., DTU [21], Tanks and Temples [22] and ETH3D [23]. By operating on higher-resolution inputs and employing complex network designs, these approaches thoroughly analyze the multi-view information in the images. The emphasis on achieving high accuracy and capturing intricate details often comes at the cost of requiring considerable computational resources and time. As a result, offline MVS is frequently employed in applications demanding precise and photorealistic scene representations, including 3D modeling and archaeological reconstruction.
In summary, our survey covers the most recent literature on learning-based MVS methods until 2023, including four main families: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. We provide a comprehensive review and insights on different aspects, including the pipelines and algorithmic intricacies. Moreover, we summarize the performance of the reviewed methods on different benchmarks, and discuss the potential future directions for deep learning-based MVS methods.
Figure 1: An overall illustration of both online and offline depth map-based MVS pipelines. Online MVS usually deals with sequential data, e.g., video, and employs TSDF volumes as an intermediate representation for mesh extraction. Given a full set of images, offline MVS holds the global information of the captured scene, and usually fuses estimated depth maps into a point cloud with filtering.
I Preliminaries
Depth map-based MVS, including most traditional and learning-based methods, typically consists of several components: camera calibration, view selection, multi-view depth estimation and depth fusion. In this section, we introduce these components to provide readers a clear picture of the MVS problem. Note that camera calibration and view selection are components for other learning-based methods as well. Sec. I-A introduces camera calibration with Structure from Motion (SfM) or SLAM. Sec. I-B discusses how to select neighboring views to accurately reconstruct the geometry. Sec. I-C explains how to build cost volumes in learning-based MVS methods with plane sweep [33]. Sec. I-D introduces the typical depth fusion strategies after depth estimation. Sec. I-E lists common datasets and benchmarks for MVS and Sec. I-F summarizes the common evaluation metrics.
{forest}
for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, draw=black, rounded corners,align=left, minimum width=2.5em, minimum height=1.2em, s sep=6pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=8em,font=, where level=3font=, where level=4font=, where level=5font=, [MVS [Traditional Methods,text width=10 em [COLMAP [34], Furu [14], Gipuma [18]…,leaf,text width=15 em] ] [Datasets & Benchmarks, text width=10 em [ScanNet [26], DTU [21], Tanks and Temples [22]…, leaf,text width=15 em] ] [Learning-based Methods
with Depth Estimation,text width=10 em [Pipeline,text width=6 em [Camera Calibration, View Selection, Multi-View Depth Estimation, Depth Fusion,leaf,text width=25.3em] ] [Supervised,text width=6em [Online Methods [MVDepthNet [25], DeepvideoMVS [31], SimpleRecon [35]…,leaf,text width=18.8 em] ] [Offline Methods [Direct 3D CNN,text width=5 em [MVSNet [15], CIDER [36]…,leaf,text width=12 em] ] [RNN,text width=5 em [R-MVSNet [37], $D^{2}$-RMVSNet [38]…,leaf,text width=12 em] ] [Coarse-to-fine,text width=5 em [CasMVSNet [32], UCSNet [39]…,leaf,text width=12 em] ] [Iterative Update,text width=5 em [PatchmatchNet [40], IterMVS [41]…,leaf,text width=12 em] ] ] ] [Unsupervised,text width=6em [End-to-end,text width=5 em [ JDACS [42], RC-MVSNet [43]…,leaf,text width=10 em] ] [Multi-stage,text width=5 em [U-MVSNet [44], KD-MVS [45]…,leaf,text width=10 em] ] ] [Semi-Supervised,text width=6em [ SGT-MVSNet [46],leaf,text width=6.8 em] ] ] [Learning-based Methods
without Depth Estimation,text width=10em [Voxel-based,text width=10 em [Atlas [47], NeuralRecon [48]…, leaf,text width=13 em] ] [NeRF-based,text width=10 em [VolSDF [49], NeuS [50]…, leaf,text width=13 em] ] [3D Gaussian Splatting-based,text width=10 em [DreamGaussian [51], SuGaR [52]…, leaf,text width=13 em] ] [Large feed-forward,text width=10 em [LRM [53], DUSt3R [54]…, leaf,text width=13 em] ] ] [Future Research Directions, text width=9.2 em [Dataset & Benchmarks, View Selection, Depth Fusion, Features, Run-time & Memory Efficiency, Prior Assistance, leaf,text width=33.6 em] ] ]
Figure 2: Taxonomy of Multi-View Stereo.
I-A Camera Calibration
[…]
I-B View Selection
The selection of views is an important step for reconstruction. It is important to balance triangulation quality, matching accuracy, and view frustum overlap[31]. Currently, there are two main strategies for view selection.
First, for most online MVS depth estimation methods [25, 58, 31], a frame is selected as a keyframe when its pose has sufficient difference compared with the previous keyframe. Then each keyframe adopts several previous keyframes to estimate depth. GP-MVS [58] proposes a heuristic pose-distance measure as:
$\text{disc}(\mathbf{T}_{ij})=\sqrt{||\mathbf{t}_{ij}||^{2}+\frac{2}{3}\text{tr}(\mathbbm{I}-\mathbf{R}_{ij})},$ (1)
where $\mathbf{T}_{ij}=[\mathbf{R}_{ij}|\mathbf{t}_{ij}]$ is the relative transformation between view $i$ and $j$. This strategy is used by many following methods [31, 35].
Second, for most offline MVS methods [15, 37, 32], view selection is done with the sparse point cloud obtained by Structure-from-Motion [34, 56]. For a reference view $i$, MVSNet [15] computes a score $s(i,j)=\sum_{\mathbf{P}}\eta(\theta_{ij}(\mathbf{P}))$ for the neighboring view $j$, where $\mathbf{P}$ is a 3D point observed by both view $i$ and $j$. $\theta_{ij}(\mathbf{P})=(180/\pi)\arccos{((\mathbf{c}_{i})-\mathbf{P})\cdot((\mathbf{c}_{j})-\mathbf{P}))}$ represents the baseline angle for $\mathbf{P}$ and $\mathbf{c}_{i},\mathbf{c}_{j}$ are the camera centers. $\eta(\cdot)$ is a piece-wise Gaussian function [59] to favor a certain baseline angle $\theta_{0}$:
$\eta(\theta)=\left\{\begin{aligned} \exp{(-\frac{(\theta-\theta_{0})^{2}}{2\sigma_{1}^{2}})},\theta\leq\theta_{0}\\ \exp{(-\frac{(\theta-\theta_{0})^{2}}{2\sigma_{2}^{2}})},\theta>\theta_{0}\end{aligned}\right.$ (2)
where $\theta_{0}$, $\sigma_{0}$, $\sigma_{1}$ are hyper-parameters. Then the view selection is done by choosing neighboring views with highest scores. Almost all the following offline MVS methods [37, 32, 40] use the same strategy.
I-C Multi-view Depth Estimation with Plane Sweep
Figure 3: Illustration of plane sweep algorithm [33]. To estimate the depth map for reference image (cam_x), neighboring source images (cam_1, cam_2) are projected with homography to fronto-parallel planes of the reference view frustum.
To form a structured data format that is more suitable for convolution operations, most learning-based MVS methods rely on the plane sweep algorithm [33] to calculate matching costs. The plane sweep algorithm discretizes the depth space with a set of fronto-parallel planes along the depth direction. This practice is deeply inspired by learning-based binocular stereo methods [60, 61, 62], which assess matching costs for a set of disparity hypotheses and subsequently estimate the disparity.
In a nutshell, the plane sweep algorithm in MVS entails iteratively sweeping planes through the object space, computing homographies between images, and selecting depth values based on consensus among different views, ultimately facilitating accurate 3D reconstruction. The plane sweep algorithm discretizes the depth space using a series of parallel planes along the depth direction. It operates by sweeping a conceptual plane through the object space and evaluating the spatial distribution of geometric surfaces. It is worth noting that, for learning-based MVS, plane sweep is conducted purely on GPU, and instead of warping hand-crafted features or edges, the dense pixels are efficiently warped.
We take a toy example of three cameras viewing the same object as illustrated in Fig. 3. Plane sweep is then performed on the frustum of cam_x (termed as the reference camera) by creating a series of fronto-parallel hypothesized planes within a given range (typically $[d_{\textrm{min}},d_{\textrm{max}}]$), with each plane corresponding to a depth value w.r.t. cam_x. Let’s then examine two 3D points, $M$ and $M^{\prime}$, as examples of occupied and unoccupied positions on hypothesized planes, respectively. For $M$, all of the three cameras capture the identical point lying on the geometry surface with photometric consistency. In contrast, the photometric consistency of the observations of $M^{\prime}$ are poor, indicating that $M^{\prime}$ is an invalid hypothesis. To evaluate the similarity of the observations towards $M$ and $M^{\prime}$, the images of cam_1 and cam_2 (termed as the source cameras) are warped to each plane by homography. These warped images are then compared against the scaled images of cam_x. Depth hypotheses with high similarity measures are considered reliable.
In practice, we divide the depth range, which is either manually set [25] or estimated by SfM [15], into discrete samples and assign hypotheses at these values. To map coordinates at depth hypothesis $d$, homography transformation [15] $\mathbf{p}_{i}(d)\sim\mathbf{H}_{i}(d)\cdot\mathbf{p}$, is applied, where $\mathbf{p}_{i}(d)$ is the corresponding pixel in the $i$-th source view for pixel $\mathbf{p}$ of the reference view. The homography $\mathbf{H}_{i}(d)$ between the $i$-th source view and the reference view 0 can be computed as:
$\mathbf{H}_{i}(d)=\mathbf{K}_{i}\mathbf{R}_{i}\left(\mathbbm{I}-\frac{(-\mathbf{R}_{0}^{\top}\mathbf{t}_{0}+\mathbf{R}_{i}^{\top}\mathbf{t}_{i})\mathbf{n}^{\top}\mathbf{R}_{0}}{d}\right)\mathbf{R}_{0}^{\top}\mathbf{K}_{0}^{-1},$ (3)
where $\mathbf{K}_{0}$, $\mathbf{K}_{i}$ denote camera intrinsics, $[\mathbf{R}_{0}|\mathbf{t}_{0}]$, and $[\mathbf{R}_{i}|\mathbf{t}_{i}]$ denote camera extrinsics, $\mathbf{n}$ denote the principle axis of the reference view. Equivalently, we can also project a reference pixel into source views with depth hypothesis [36, 40, 41]. We compute $\mathbf{p}_{i}(d)$ in the $i$-th source view for pixel $\mathbf{p}$ in the reference and depth hypothesis $d$ as follows:
$\mathbf{p}_{i}(d)=\mathbf{K}_{i}\cdot\left(\mathbf{R}_{i}\mathbf{R}_{0}^{\top}\cdot(\mathbf{K}_{0}^{-1}\cdot\mathbf{p}\cdot d)-\mathbf{R}_{i}\mathbf{R}_{0}^{\top}\cdot\mathbf{t}_{0}+\mathbf{t}_{i}\right).$ (4)
The warped source feature maps are then obtained via differentiable interpolation.
Regarding selecting $D$ depth hypotheses from the triangulated depth range $[d_{\textrm{min}},d_{\textrm{max}}]$, there are two main schemes, namely the forward depth sampling and the inverse depth sampling. The naive forward sampling divides the depth range into $D-1$ depth intervals with identical lengths. Given a depth index $k$, we have
$d_{k}=d_{\textrm{min}}+\frac{k}{D-1}(d_{\textrm{max}}-d_{\textrm{min}}),k=0,\ldots,D-1,$ (5)
where the depth hypotheses distribute uniformly between the two ends. While for the inverse sampling scheme [36], we sample uniformly in the multiplicative inverse of $d$, such that
$\frac{1}{d_{k}}=\frac{1}{d_{\textrm{max}}}+\frac{k}{D-1}(\frac{1}{d_{\textrm{min}}}-\frac{1}{d_{\textrm{max}}}),k=0,\ldots,D-1.$ (6)
In this way, the distribution of sampled depth values becomes sparser with $d$ approaching to $d_{\textrm{max}}$. When reconstructing unbounded outdoor scenes, where the depth range is rather large, the inverse sampling will be a reasonable choice since it samples more densely at the foreground. In addition, as revealed in [63], the inverse sampling leads to a uniform sampling on the projected epipolar lines of source images.
So far, for each depth hypothesis made for cam_x, we have the warped source feature maps as well as the scaled reference feature map. The photometric similarity (or matching cost) is measured in a one-to-many manner between reference and warped source features for the following depth estimation, which will be discussed in Sec. II-B. By performing plane sweep, we obtain a cost volume with a regular spatial shape, making it easy for CNNs to process in parallel.
I-D Depth Fusion
For depth map-based MVS, after estimating all the depth maps, we need to fuse them into a dense 3D representation, e.g., point cloud or mesh. Online MVS methods [31, 35] usually adopt TSDF (Truncated Signed Distance Function) fusion [64, 65] to fuse the depth maps into a TSDF volume and then use Marching Cube [66] to extract the mesh. However, there usually exist outliers in the depth maps, which may reduce the reconstruction accuracy. To overcome this problem and improve the accuracy, offline MVS methods [15, 32, 40] typically filter the depth maps before fusing into a point cloud, which is motivated by Galliani et al. [18]. There are two main filtering steps: photometric consistency filtering and geometric consistency filtering [15]. For photometric consistency filtering, a per-pixel confidence, discussed in Sec. II-G, is estimated to measure the confidence of depth estimation, i.e., probability that the ground truth depth is within a small range near the estimation. Then a threshold can be set to filter depth values with low confidence. For geometric consistency filtering, the consistency of depth estimations are measured among multiple views. For a pixel $\mathbf{p}$ in the reference view 0, we project it to pixel $\mathbf{p}^{\prime}$ in its $i$-th neighboring view through its depth prediction $d_{0}(\mathbf{p})$. After looking up the depth for $\mathbf{p}^{\prime}$, $d_{i}(\mathbf{p}^{\prime})$, we re-project $\mathbf{p}^{\prime}$ back to the reference view at pixel $\mathbf{p}^{\prime\prime}$ and look up its depth $D_{0}(\mathbf{p}^{\prime\prime})$. We consider pixel $\mathbf{p}$ and its depth as consistent to the $i$-th neighboring view, if the distances, in image space and depth, between the original estimate and its re-projection satisfy:
$\xi_{d}=\|\mathbf{p}-\mathbf{p}^{\prime\prime}\|_{2}\leq\tau_{1},$ (7)
$\xi_{p}=\frac{\|d_{0}(\mathbf{p}^{\prime\prime})-d_{0}(\mathbf{p})\|}{d_{0}(\mathbf{p})}\leq\tau_{2},$ (8)
where $\tau_{1}$ and $\tau_{2}$ are thresholds. The pixels are considered to be reliable estimations if they are consistent in at least $N$ neighboring views. Recently, instead of using predefined thresholds for $\tau_{1}$ and $\tau_{2}$, dynamic consistency checking [38] is proposed to dynamically aggregate geometric matching error among all views and improve the robustness of reconstruction. Specifically, the dynamic multi-view geometric consistency $C_{\textrm{geo}}(\mathbf{p})$ is computed as:
$C_{\textrm{geo}}(\mathbf{p})=\sum_{i}\exp{(-(\xi_{p}+\lambda\xi_{d}))},$ (9)
where $\lambda$ is a weight to balance the reprojection error in two metrics. Then the outliers with $C_{\textrm{geo}}(\mathbf{p})$ smaller than a threshold are filtered.
I-E Datasets and Benchmarks
Tab. I is a brief summary of commonly used public MVS datasets and benchmarks for training and evaluation.
TABLE I: An overview of commonly used datasets and benchmarks for learning-based Multi-View Stereo.
Dataset Provided Ground Truth1 Synthetic Online Evaluation
Camera pose Depth Map Point Cloud Mesh Benchmark Target
ScanNet [26] Depth Map / Mesh
7-Scenes [27] Depth Map
DTU [21] Point Cloud
Tanks and Temples [22] Point Cloud
ETH3D [23] Point Cloud
BlendedMVS [24] Depth Map
  • 1
    For datasets with online benchmark, the point cloud ground truth of test set is not released.
ScanNet: ScanNet [26] is a large RGB-D dataset that contains 1613 indoor scenes with ground-truth camera poses, depth maps, surface reconstruction and semantic segmentation labels. Online MVS methods [25, 31, 35] mainly use ScanNet for training and testing.
7-Scenes: 7-Scenes dataset [27] is a RGB-D dataset captured in indoor scenes with a handheld Kinect RGB-D camera. Since it is relatively small, 7-Scenes is usually used to test the generalization performance of the models trained on ScanNet [26] without finetuning.
DTU: DTU dataset [21] is an object-centric MVS dataset collected under well-controlled laboratory conditions with known accurate camera trajectory. It contains 128 scans with 49 or 64 views under 7 different lighting conditions. Since DTU dataset officially provides scanned ground truth point clouds instead of depth maps, it is required to generate mesh models with surface reconstruction, e.g., screened Poisson surface reconstruction algorithm [67], and then render depth maps [15] for training.
Tanks and Temples: Tanks and Temples [22] is a large-scale benchmark captured in more complex real indoor and outdoor scenarios. It is divided into intermediate and advanced sets. Different scenes have different scales, surface reflection and exposure conditions. It is usually used to test the generalization performance. Note that Tanks and Temples does not provide ground truth camera parameters, which are usually estimated with Structure-from-Motion methods such as COLMAP [17] or OpenMVG [56].
ETH3D: ETH3D benchmark [23] contains 25 large-scale scenes, including indoor and outdoor, with high-resolution RGB images. The scenes typically contain many low-textured regions and non-Lambertian surfaces, e.g., white walls and reflective floor. In addition, the images are usually sparse and have strong viewpoint variations and occlusions. Therefore, ETH3D is considered as a very challenging benchmark. Similar to Tanks and Temples, ETH3D is usually used to test the generalization performance.
[…]
I-F Evaluation Metrics
As mentioned in Tab. I, based on the ground truth, e.g., depth maps or point clouds, evaluation metrics can be categorized into 2D metrics and 3D metrics.
I-F1 2D Metrics
2D depth-based metrics are commonly used by online MVS methods [25, 31] to evaluate the depth maps [68, 69]. We list the commonly used metrics as follows: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel), mean absolute inverse depth error (Abs Inv), and inlier ratio with threshold 1.25 ($\delta<1.25$):
[…]
where $d(\mathbf{p})$ denotes the ground truth depth, $\hat{d}(\mathbf{p})$ denotes the estimation, $N$ denotes the number of pixels with valid depth measurements and $\mathbbm{1}[\cdot]$ denotes the indicator function.
I-F2 3D Metrics
3D point cloud evaluation is widely used by offline MVS methods [15, 37]. Note that the reconstructed point clouds should be aligned to ground truth point clouds before evaluation, e.g., by Iterative Closest Point (ICP), if their extrinsics are differently calibrated.
Precision/Accuracy: Precision/Accuracy measures the percentage of predicted points that can be matched to the ground truth point cloud. Considering a point $\mathbf{P}_{p}$ in the predicted point cloud, it is considered to have a good match in the ground truth point cloud $\{\mathbf{P}_{g}\}$ if
$\|\mathbf{P}_{p}-\mathop{\arg\min}\limits_{\mathbf{P}\in\{\mathbf{P}_{g}\}}\|\mathbf{P}-\mathbf{P}_{p}\|_{2}\|_{2}\leq\lambda,$ (11)
where $\lambda$ is a scene-dependent threshold assigned by datasets. Usually $\lambda$ is set to a large value for large-scale scenes. Note that in some datasets, instead of being measured by percentage [22, 23], precision/accuracy is measured by mean or median absolute distance [21].
Recall/Completeness: Recall/Completeness measures the percentage of ground truth points that can be matched to the predicted point cloud. For a point $\mathbf{P}_{g}$ in the ground truth point cloud, it is considered a good match in the predicted point cloud $\{\mathbf{P}_{p}\}$ if
$\|\mathbf{P}_{g}-\mathop{\arg\min}\limits_{\mathbf{P}\in\{\mathbf{P}_{p}\}}\|\mathbf{P}-\mathbf{P}_{g}\|_{2}\|_{2}\leq\lambda.$ (12)
Similar to precision/accuracy, recall/completeness is sometimes measured by the percentage of points [22, 23] and sometimes measured by mean or median absolute distance [21], whose definition is similar to Chamfer distance.
F-Score: The two aforementioned metrics measure the accuracy and completeness of predicted point clouds. However, each of these metrics alone cannot present the overall performance since different MVS methods adopt different assumptions. A stronger assumption usually leads to higher accuracy but lower completeness. If only precision/accuracy is reported, it would favor MVS algorithms that only include estimated points of high certainty. On the other hand, if only recall/completeness is reported it would favor MVS algorithms that include everything, regardless of accuracy. Therefore, F-score, an integrated metric, is introduced [22, 23]. F-score is the harmonic mean of precision and recall. It is sensitive to extremely small values and tends to get more affected by smaller values, which means that F-score does not encourage imbalanced results. However, in most cases, F-score still suffers from unfairness due to the limitations of ground truth, e.g., sparse and incomplete point clouds may penalize for filling in the areas that are not present in the ground truth [47].
II Supervised Method with Depth Estimation
This section mainly introduces supervised learning-based MVS methods with depth estimation. A typical depth map-based MVS pipeline mainly consists of the feature extraction (Sec. II-A), cost volume construction (Sec. II-B), cost volume regularization (Sec. II-C) and depth estimation (Sec. II-E). The pipelines of MVDepthNet [25] and MVSNet [15] are illustrated in Fig. 4 and Fig. 5, which are the representative methods of online and offline MVS methods respectively.
II-A Feature Extraction
Considering efficiency, most methods use simple CNN structures to extract deep features from images, e.g., ResNet [70], U-Net [71] and FPN [72]. For online MVS methods, feature extraction networks are usually used in line with the real-time operation goal. DeepVideoMVS [31] combines MNasNet [73], which is lightweight and has low latency, with FPN. SimpleRecon [35] proposes utilizing the first two blocks from ResNet18 [70] and an EfficientNet-v2 [74] encoder, which maintains efficiency and yields a sizeable improvement in depth map accuracy. For offline MVS methods, MVSNet [15] uses a stacked eight-layer 2D CNN to extract deep features for all the images. Coarse-to-fine methods further extract multi-scale features for estimation on multiple scales with FPN [32, 39, 40] or multi-scale RGB images [75, 76]. Recently, many following works have been paying more attention to feature extraction to improve the representation power of deep features. [77, 78, 79] introduce deformable convolutions to adaptively learn receptive fields for areas with varying richness of texture. CDS-MVSNet [80] uses a dynamic scale network, CDSFNet, to extract the discriminative features by analyzing the normal curvature of the image surface. GeoMVSNet [76] proposes a geometry fusion network that fuses the image features with coarse depth maps. With FPN as the main feature extractor, WT-MVSNet [81] and MVSFormer [82], further introduces Vision Transformers [83, 84, 85] for image feature enhancement. ET-MVSNet [86] integrates the Epipolar Transformer into the feature extraction, which performs non-local feature aggregation on the epipolar lines.
II-B Cost Volume Construction
For both online and offline MVS methods, the cost volume is constructed with the plane sweep algorithm as discussed in Sec. I-C.
II-B1 Online MVS
To reduce computation and improve the efficiency for online applications, online MVS methods usually construct 3D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D}$, which store a single value as matching cost for each pixel $\mathbf{p}$ and depth sample $d_{i}$. MVDepthNet [25] and GP-MVS [58] compute the per-pixel intensity difference between the reference and each source view as matching cost. If there is more than one source view, the cost volumes are averaged. Instead of intensity difference, Neural RGB-D [87] computes deep feature difference. DeepVideoMVS [31] and MaGNet [88] calculate the cost from pixel-wise correlation between the reference feature and the warped source feature. In addition to the dot product between reference image features and warped source image features, SimpleRecon [35] further adds metadata, e.g., ray direction and relative pose, into the 4D cost volume. Then an MLP is used to reduce the dimension of 4D cost volume to a 3D cost volume.
II-B2 Offline MVS
Offline MVS methods mainly focus on reconstructing high-quality dense geometry with high-resolution images. To encode more matching information and improve the quality, offline methods [15, 32] usually construct 4D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D\times C}$, where each pixel $\mathbf{p}$ and depth sample $d_{i}$ is associated with a matching cost of dimension $C$. Since there may exist an arbitrary number of source views and serious occlusion [23], robustly aggregating matching information from all the source views is an important step. MVSNet [15] proposes a variance-based cost metric as follows:
$\mathbf{C}={\text{Var}}\left(\mathbf{V}_{0},\cdots,\mathbf{V}_{N-1}\right)=\frac{\sum_{i=0}^{N-1}\left(\mathbf{V}_{i}-\overline{\mathbf{V}}\right)^{2}}{N},$ (13)
where $\{\mathbf{V}_{i}\}_{i=0}^{N-1}$ are (warped) feature volumes, $\overline{\mathbf{V}}$ is the average feature volume. To further reduce dimension, CIDER [36] adopts group-wise correlation [89] to compute a lightweight cost volume between reference and each warped source view. Then $N-1$ cost volumes are averaged for regularization. Without considering occlusions, these methods consider all source views equally with the averaging operation.
Nonetheless, it is essential to emphasize that occlusions are crucial to consider, as they pose common challenges in MVS, frequently leading to invalid matching and inaccurate estimations [34]. Incorporating visibility information to aggregate matching details from source views can substantially bolster robustness against occlusions, thereby enhancing the accuracy of the reconstruction process [34].
To estimate view weights for source views, PVA-MVSNet [90] applies gated convolution [91] to adaptively aggregate cost volumes. View aggregation tends to give occluded areas smaller weights and the reweighting map is yielded according to the volume itself. Vis-MVSNet [92] aggregates pair-wise cost volumes by weighted sum, where the weight is negatively related to the uncertainty of depth probability distribution.  [63, 40, 41] utilize pixel-wise view weight networks to learn view weights from the pair-wise cost volumes without supervision.
II-C Cost Volume Regularization
Usually, the raw cost volume constructed from image features may be noisy and should be incorporated with smoothness constraint for depth estimation [15]. Therefore, cost volume regularization is an important step to refine the raw cost volume and aggregate matching information from a large receptive field.
II-C1 Online MVS
A 2D encoder-decoder architecture is commonly used to aggregate information. MVDepthNet [25] and GP-MVS [58] concatenate the reference image and cost volume and send them to an encoder-decoder architecture with skip connections. In addition, GP-MVS [58] uses Gaussian process to fuse information from previous views. Neural-RGBD [87] accumulates depth probability volumes over time with a Bayesian filtering framework to effectively reduce depth uncertainty and improve robustness and temporal stability. DeepVideoMVS [31] applies 2D U-Net [71] on the cost volume and adds skip connections between the image encoder and cost volume encoder at all resolutions. It further uses ConvLSTM [93] to propagate past information with a small overhead of computation time and memory consumption. Long et al. [94] use MatchNet to regularize the cost volume and ContextNet to learn 2D context information from the reference view. By concatenating the cost volume and context feature, a hybrid volume is fed into the Epipolar Spatio-Temporal transformer to aggregate temporal information. SimpleRecon [35] fuses multi-scale image features, extracted from the pretrained EfficientNetv2 [74], into the cost volume encoder to improve the performance.
Figure 4: Pipeline of MVDepthNet [25]. Multiple image frames are encoded in the cost volume. MVDepthNet takes the reference image and the cost volume as the input and outputs inverse depth maps of four different resolutions. Skip connections between the encoder and decoder of the same resolution are used for the per-pixel depth estimation.
Figure 5: Pipeline of MVSNet [15]. Reference and source images go through the feature extraction network, followed by the differentiable homograph warping to construct the cost volume. A 3D U-Net is used to regularize the cost volume into a probability volume. The final depth map is estimated from the probability volume and refined with the reference image features.
II-C2 Offline MVS
Among most learning-based offline MVS methods that use 4D cost volumes, there are three main categories to perform cost volume regularization: direct 3D CNN, coarse-to-fine and RNN. Fig. 6 illustrates these three main schemes.
Figure 6: Illustration of typical cost regularization schemes for offline MVS methods that use 4D cost volumes. (a) direct 3D CNN regularization [15, 95, 96, 36] applies a 3D CNN to aggregate contextual information; (b) coarse-to-fine regularization [32, 39, 75, 63] constructs multi-scale cost volumes based on coarse prediction and uses 3D CNN regularization on each scale; (c) RNN regularization [37, 38, 77] sequentially regularizes 2D slices of the cost volume to reduce memory consumption.
Direct 3D CNN: Similar to stereo estimation [60, 61, 62], 3D CNN is widely used for cost volume regularization in MVS [15, 97, 95, 96, 36]. As the blueprint of learning-based offline MVS methods, MVSNet [15] adopts a 3D U-Net [71] to regularize the cost volume, which aggregates context information from a large receptive field with relatively low computation cost. It is found that the 3D regularization can capture better geometry structures, perform photometric matching in 3D space, and alleviate the influence of image distortion caused by perspective transformation and occlusions [95]. However, since 3D CNN is memory and run-time consuming, many offline methods [15, 95, 96, 36] use limited depth hypotheses and estimate depth maps at low resolution.
RNN: Instead of 3D CNN, R-MVSNet [37] sequentially regularizes 2D slices of the cost volume along the depth dimension with a convolutional GRU [98], which is able to gather spatial and uni-directional context information in the depth dimension. $D^{2}$-RMVSNet [38] augments R-MVSNet [37] with a complex convolutional LSTM [93]. AA-RMVSNet [77] further introduce an intra-view feature aggregation module for feature extraction and an inter-view cost volume aggregation module to adaptively aggregate cost volumes of different views. With sequential process of 2D slices, these methods improve the scalability for high-resolution reconstruction as well as large-scale scenes and reduce memory, however, at the cost of run-time.
Coarse-to-Fine: Predicting depth in a coarse-to-fine manner [32, 39, 75, 40, 78, 63, 99] is another solution of reducing both memory consumption and running-time. A coarse depth map is first predicted and then upsampled and refined during finer stages to construct fine details. CasMVSNet [32] constructs cascade cost volumes by warping features with reduced depth ranges around the previous coarse maps. Based on Eq. 3, the homography of differential warping at stage $k+1$ is
$\mathbf{H}_{i}^{(k+1)}(d^{(k)}+\Delta^{(k+1)})=(d^{(k)}+\Delta^{(k+1)})\mathbf{K}_{0}\mathbf{T}_{0}\mathbf{T}^{-1}_{i}\mathbf{K}^{-1}_{i},$ (14)
where $d^{(k)}$ is the estimated depth value at stage $k$ and $\Delta^{(k+1)}$ is the residual depth to be determined in the current stage. Following MVSNet [15], 3D CNN is used on each stage to regularize the cost volume. UCS-Net [39] estimates uncertainty from the coarse prediction to adaptively adjust search ranges in finer stages. CVP-MVSNet [75] constructs the cost volume with a proposed optimal depth resolution of half pixel to narrow depth range in finer stages. EPP-MVSNet [100] introduces an epipolar-assembling module to assemble high-resolution information into cost volume and an entropy-based process to adjust depth range. TransMVSNet [79] uses a Feature Matching Transformer for robust long-range global context aggregation within and across images. WT-MVSNet [81] uses a window-based Epipolar Transformer for enhanced patch-to-patch matching, and a window-based Cost Transformer to better aggregate global information.
II-D Iterative Update
Diverging from conventional approaches, certain methods [40, 41, 101, 102] adopt iterative updates to gradually refine depth maps. Iterative methods introduce a dynamic approach to depth map estimation, enabling iterative refinement of the reconstruction process. By iteratively updating the depth maps based on successive iterations, these methods can progressively improve the accuracy and consistency of the reconstructed 3D scene. This iterative refinement is particularly advantageous in scenarios where initial estimations may be coarse or inaccurate, as it allows for finer adjustments to be made to the depth maps over multiple iterations. Moreover, the ability to control the number of iterations provides users with the flexibility to prioritize either computational efficiency or reconstruction quality, depending on specific application requirements.
Some methods [40, 103] combine iterative PatchMatch [104] with deep learning. PatchMatch algorithm is widely used in many traditional MVS methods [105, 18, 34, 19]. The PatchMatch algorithm mainly consists of: random initialization, propagation of hypotheses to neighbors, and evaluation for choosing best solutions. After initialization, the approach iterates between propagation and evaluation until convergence. Traditional methods usually design fixed patterns for propagation, e.g., propagation of Gipuma [18] as shown in Fig. 7. Recently, PatchmatchNet [40] proposes learned adaptive propagation and cost aggregation modules, which enables PatchMatch to converge faster and deliver more accurate depth maps. Instead of propagating depth samples naively from a static set of neighbors as done traditionally [18], PatchmatchNet adaptively gathers samples from pixels of the same surface with the guidance of image features, shown in Fig. 8. It iteratively updates the depth maps in a coarse-to-fine structure. Without 3D CNN, PatchmatchNet demonstrates a remarkably low computation time and memory than other coarse-to-fine methods [32, 39, 75]. Another PatchMatch-based method, PatchMatch-RL [103], jointly estimates depth, normal and visibility with a coarse-to-fine structure. Considering argmax based hard decisions/sampling of PatchMatch is non-differentiable, PatMatch-RL adopts reinforcement learning in the training process.
Figure 7: Traditional chessboard propagation scheme of Galliani et al. [18]. Planes from a local neighborhood (red points) serve as candidates to update a given pixel (black). (a) Standard pattern with 20 neighbors. (b) Simplified pattern with 8 neighbors.
Figure 8: Adaptive propagation scheme of Wang et al. [40]. Pixels located at the object boundary (yellow) and a textureless region (red) receive depth hypotheses from sampled neighbors (green and orange). (a) Reference image. (b) Fixed sampling locations of classic propagation. (c) Adaptive sampling locations with adaptive propagation. The grayscale image in (b) and (c) is the ground truth depth map.
Recently, RAFT [106] estimates optical flow by iteratively updating a motion field through a GRU and achieves state-of-the-art performance. The idea is further adopted in stereo [107], scene flow [108] and SfM [109]. In MVS, several methods [41, 102, 101] also adopted this approach to enhance efficiency and flexibility. IterMVS [41] proposes a lightweight GRU-based probability estimator that encodes the per-pixel probability distribution of depth in its hidden state. In each iteration, multi-scale matching information is injected to update the pixel-wise depth distribution. CER-MVS [101] and Effi-MVS [102] both embed the RAFT module in a coarse-to-fine structure. With the injection of matching information, the hidden state is updated and outputs a residual that is added to the previous depth estimation. DELS-MVS [110] proposes Epipolar Residual Network to search for the corresponding point in the source image directly along the corresponding epipolar line and follows an iterative manner to narrow down the search space. Based on IterMVS [41], RIAV-MVS [111] iteratively refines index fields with a 3-level GRUs. MaGNet [88] iteratively updates the Gaussian distribution of depth for each pixel with the matching scores. IGEV-MVS [112] iteratively updates a disparity map regressed from geometry encoding cost volumes and pairs of group-wise correlation volumes.
II-E Depth Estimation
II-E1 Online MVS
Many methods [25, 58, 31, 35] apply encoder-decoder on the cost volume $\mathbf{C}$ and reduce the feature channel to 1, i.e., $\mathbf{C}^{\prime}\in\mathbb{R}^{H\times W\times 1}$. Then the sigmoid activation $\sigma$ is applied for normalization. Together with the predefined depth range $[d_{\textrm{min}},d_{\textrm{max}}]$ (mostly manually set), the depth map can be computed. For example, DeepVideoMVS [31] estimates depth as:
$\hat{d}(\mathbf{p})=\left(\left(\frac{1}{d_{\textrm{min}}}-\frac{1}{d_{\textrm{max}}}\right)\cdot\sigma(\mathbf{C}^{\prime}(\mathbf{p}))+\frac{1}{d_{\textrm{max}}}\right)^{-1},$ (15)
where $\mathbf{p}$ is the pixel coordinate. For other methods, Neural-RGBD [87] directly models the cost volume as probability volume $\mathbf{P}$ and adopts soft argmax  [60] to predict depth, Eq. 16. MaGNet [88] takes the mean of Gaussian distribution at last iteration as depth.
II-E2 Offline MVS
For a cost volume $\mathbf{C}\in\mathbb{R}^{H\times W\times D\times C}$, a probability volume $\mathbf{P}\in\mathbb{R}^{H\times W\times D}$ is usually generated after cost volume regularization, which is then used for depth estimation. Currently, almost all the learning-based MVS methods use exclusively either regression (soft argmax) or classification (argmax) to predict depth.
Following GCNet [60], MVSNet [15] uses soft argmax to regress the depth map with sub-pixel precision. Specifically, the expectation value along the depth direction of probability volume $\mathbf{P}$ is computed as the final prediction:
$\hat{d}(\mathbf{p})=\sum_{i=1}^{D}d_{i}\cdot\mathbf{P}(i,\mathbf{p}),$ (16)
where $\mathbf{p}$ is the pixel coordinate, $d_{i}$ is $i$-th depth sample and $P(i,\mathbf{p})$ is the predicted depth probability. For coarse-to-fine methods [32, 39, 75], soft argmax is applied on each stage to regress the depth maps. On the other hand, some methods [36, 40] compute the expectation value of inverse depth samples since this sampling is more suitable for complex and large-scale scenes [36]. The depth prediction is computed as:
$\hat{d}(\mathbf{p})=\left(\sum_{i=1}^{D}\frac{1}{d_{i}}\cdot\mathbf{P}(i,\mathbf{p})\right)^{-1}.$ (17)
In contrast, methods [37, 38, 77] that use RNN for cost volume regularization mainly adopt argmax operation. They choose the depth sample with the highest probability as the final prediction, which is similar to classification. Since the argmax operation adopted by winner-take-all cannot produce depth estimations with sub-pixel accuracy, the depth map may be refined in post-processing [37].
Comparatively, soft argmax operation corresponds to measuring the distance of the expectation to the ground truth depth. While the expectation can take any continuous value, the measure cannot handle multiple modes in $\mathbf{P}$ and strictly prefers unimodal distributions [41]. The argmax operation corresponds to measuring the Kullback-Leibler divergence between a one-hot encoding of the ground truth and $\mathbf{P}$, but cannot achieve sub-pixel precision [41]. Recently, Wang et al. [41] propose a simple yet effective hybrid strategy that combines regression and classification. Similar to [113], it is robust to multi-modal distributions but also achieves sub-pixel precision. First, similar to R-MVSNet [37], the index $\mathbf{X}(\mathbf{p})$ with the highest probability for pixel $\mathbf{p}$ from probability $\mathbf{P}$ is found. Then the expectation in the local inverse range is computed as the depth estimate $\hat{d}(\mathbf{p})$:
$\hat{d}(\mathbf{p})\!\!=\!\!\left(\!\!\frac{1}{\sum_{j=\mathbf{X}(\mathbf{p})-r}^{\mathbf{X}(\mathbf{p})+r}\mathbf{P}(j,\mathbf{p})}\!\sum_{j=\mathbf{X}(\mathbf{p})-r}^{\mathbf{X}(\mathbf{p})+r}\!\!\frac{1}{d_{j}}\!\cdot\!\mathbf{P}(j,\mathbf{p})\!\!\right)^{\!\!\!-1}\!\!\!\!\!,$ (18)
where $d_{j}$ is the $j$-th depth sample. Concurrently, Peng et al. [114] propose a different strategy to unify regression and classification. They first use classification to get the optimal hypothesis and then regress the proximity for it.
DMVSNet [115] estimates two depth maps on multi-stages and composes the final depth map by alternating between selecting the maximum and minimum predicted depth values, which generates an oscillating depth map for improved geometry.
II-F Depth Refinement
Given that the raw depth estimation from MVS may be noisy, refinement is usually used to improve the accuracy. R-MVSNet [37] enforces multi-view photo-consistency to alleviate the stair effect and achieve sub-pixel precision. Point-MVSNet [95] and VA-Point-MVSNet [116] use PointFlow to refine the point cloud iteratively by estimating the residual between the depth of the current iteration and that of the ground truth. Fast-MVSNet [117] adopts an efficient Gauss-Newton layer to optimize the depth map by minimizing the feature residuals. PatchmatchNet [40] refines the final upsampled depth map with a depth residual network [118] and reference image feature.  [41, 102, 101] use the mask upsampling module from RAFT [106] to upsample and refine the depth map to full resolution by computing the weighted sum of depth values in a window based on the mask. Based on a coarse depth map, RayMVSNet [119, 120] aggregates multi-view image features with an epipolar transformer and use a 1D implicit field to estimate the SDF of the sampled points and the location of the zero-crossing point. GeoMVSNet [76] filters the depth map by geometry enhancement in the frequency domain. EPNet [121] uses a hierarchical edge-preserving residual learning module to refine multi-scale depth estimation with image context features.
II-G Confidence Estimation
As discussed in Sec. I-D, photometric confidence is important to filter out unreliable estimations during depth fusion for offline MVS methods. Following MVSNet [15], most offline MVS methods take the probability of the estimation [37, 38, 77] or the probability sum over several samples near the estimation [15, 32, 40] from the probability volume as confidence. Fig. 9 provides a set of RGB images, estimated depth maps, and the visualized confidence. In stereo matching, some methods learn confidence from disparity [122], RGB image [123, 124] or matching costs [125] and obtain confidence scores in $[0,1]$ interval. Motivated by this, some traditional MVS methods [126, 127] based on classical PatchMatch [105] propose to estimate the confidence with deep learning and use this to refine the results from PatchMatch. Recently, IterMVS [41] estimates confidence from the pixel-wise depth probability distributions, which are encoded by the hidden state of a convolutional GRU. A 2D CNN followed by a sigmoid is applied to the hidden state to predict the confidence. DELS-MVS [110] feeds pixel-wise entropy of the partition probabilities, which is computed for each source image, into a confidence network to learn a confidence map. The confidence map is used to guide the fusion of multiple depth maps.
Figure 9: Visualization of confidence on DTU [21]. The regions with low confidence typically correspond to occlusions and low texture.
II-H Loss Function
II-H1 Online MVS
Many methods [25, 58, 31] compute the regression loss of estimated inverse depth maps for training:
$L=\sum_{\mathbf{p}}\|\frac{1}{d(\mathbf{p})}-\frac{1}{\hat{d}(\mathbf{p})}\|_{1},$ (19)
where $\mathbf{p}$ denotes the pixel coordinate, $d(\mathbf{p})$ denotes the ground truth depth, $\hat{d}(\mathbf{p})$ denotes the estimation and $||\cdot||_{1}$ denotes the $L_{1}$ loss. For other methods, Neural-RGBD [87] uses Negative-Log Likelihood (NLL) over the depth with its depth probability volume. Similarly, MaGNet [88] uses NLL loss since the per-pixel depth is modeled as Gaussian distribution. SimpleRecon [35] computes the regression loss with log-depth. To improve performance, SimpleRecon further uses gradient loss on depth, normal loss where normal is computed with depth and intrinsics, and multi-view regression loss.
II-H2 Offline MVS
Based on the depth estimation strategy as discussed in Sec. II-E, loss functions can be mainly categorized into regression and classification. For methods [15, 36] that predict depth with soft argmax [60], (smooth) $L_{1}$ loss is usually adopted as the loss function, which is stated as:
$L=\sum_{\mathbf{p}}\|d(\mathbf{p})-\hat{d}(\mathbf{p})\|_{1}.$ (20)
For coarse-to-fine methods [32, 39, 75, 40], the $L_{1}$ loss is computed on each stage for multi-stage supervision.
For methods[37, 38, 77] that predict depth with an argmax operation, cross entropy loss is commonly used for the loss function since the problem is multi-class classification. The cross entropy loss function is defined as:
$L=\sum_{\mathbf{p}}\left(\sum_{i=1}^{D}-G(i,\mathbf{p})\cdot\log[P(i,\mathbf{p})]\right),$ (21)
where $G(i,\mathbf{p})$ is the ground truth one-hot vector of depth at pixel $\mathbf{p}$, and $P(i,\mathbf{p})$ is the predicted depth probability.
For methods that predict depth with a hybrid strategy of classification and regression, Wang et al. [41] adopt both $L_{1}$ loss and cross entropy loss to supervise the regression and classification respectively, while Peng et al. [114] use focal loss [128].
III Unsupervised & Semi-supervised Methods with Depth Estimation
Figure 10: Ambiguous supervision from photometric consistency assumption [43]. (a) Complicated lighting condition caused by varying camera exposures. (b) Invisibility of specific area caused by occlusion.
In this section, we introduce unsupervised learning-based MVS methods [129, 130, 43, 44, 42] and semi-supervised method [46]. Supervised MVS methods mentioned in Sec. II depend extensively on the availability of accurate ground truth depth maps obtained through depth-sensing equipment. This requirement not only complicates the data collection process, making it labor-intensive and costly but also restricts these methods to limited datasets and primarily indoor settings. To make MVS practical in more general real-world scenarios, it is vital to consider alternative unsupervised learning-based methods that can provide competitive accuracy compared to the supervised ones without any ground truth depth. Existing unsupervised methods are built upon the assumption of photometric consistency (Sec. III-A), which indicates that corresponding pixels from different views of the same 3D point share similar features. These approaches are categorized into end-to-end ones (Sec. III-B) and multi-stage ones (Sec. III-C). SGT-MVSNet [46] is the only semi-supervised method so far, and we introduce it in Sec. III-D. We also analyze the primary challenges in unsupervised MVS in Sec. III-B.
III-A Photometric Consistency Assumption
In the realm of unsupervised depth map prediction, extant methods [131, 44, 132, 133] endeavor to establish photometric consistency between reference and source perspectives. This pivotal notion revolves around the augmentation of similarity between the reference image $\mathbf{I}_{0}$ and individual source images $\mathbf{I}_{i}$ after their warping to align with the reference view.
In relation to the depth estimation denoted as $\hat{d}$ for the initial image $\mathbf{I}_{0}$, the process involves the projection of reference pixels into the subsequent image $\mathbf{I}_{i}$ using the formulation presented in Eq. 4.
Subsequently, the distorted version of the source image denoted as $\hat{\mathbf{I}}_{0}^{i}$ is created by interpolating the RGB values of the source image at the displaced pixel positions through bilinear sampling. This interpolation is carried out at the locations where the pixels have been transformed due to the warping process. Additionally, alongside the warped image $\hat{\mathbf{I}}_{0}^{i}$, a binary mask $M_{i}$ is commonly generated. This mask is employed to exclude pixels that have been projected beyond the boundaries of the image and are, thus, considered invalid.
The photometric consistency loss $L_{\textrm{PC}}$ can be written as:
$\begin{split}L_{\textrm{PC}}=&\sum_{i=1}^{N-1}\frac{1}{\left\|M_{i}\right\|_{1}}(\left\|\left(\hat{\mathbf{I}}_{0}^{i}-\mathbf{I}_{0}\right)\odot M_{i}\right\|_{2}\\ &+\left\|\left(\nabla\hat{\mathbf{I}}_{0}^{i}-\nabla\mathbf{I}_{0}\right)\odot M_{i}\right\|_{2}),\end{split}$ (22)
where $\nabla$ denotes the gradient at the pixel level, while $\odot$ symbolizes element-wise Hadamard multiplication. In most cases [131, 42, 44, 132, 43], the incorporation of structural similarity loss and depth smoothness loss into the computation is commonplace. This practice aims to enhance the training process’s stability and speed up the convergence.
The computation of the structural similarity loss occurs between a synthesized image and the initial reference image, aiming to uphold contextual congruity. More precisely, the assessment of contextual similarity often involves the utilization of the Structural Similarity Index (SSIM) proposed by [134], a metric commonly employed for quantifying contextual consistency. This metric is defined as follows:
$L_{\textrm{SSIM}}=\sum_{i=1}^{N-1}\left[1-\operatorname{SSIM}\left(\mathbf{I}_{0},\hat{\mathbf{I}}_{0}^{i}\right)\right]\odot M_{i},$ (23)
$\operatorname{SSIM}(x,y)=\frac{\left(2\mu_{x}\mu_{y}+c_{1}\right)\left(2\sigma_{xy}+c_{2}\right)}{\left(\mu_{x}^{2}+\mu_{y}^{2}+c_{1}\right)\left(\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2}\right)},$ (24)
where $\mu$, $\sigma^{2}$ represent the mean and variance of the images, $c_{1},c_{2}$ are constants to avoid numerical issues.
The incorporation of a smoothness loss term serves to promote the continuity of depth information within the context of image and depth disparity alignment. This continuity is evaluated based on the color intensity gradient present in the input reference image. The computation of the smoothness loss, $L_{\textrm{SM}}$, is defined as follows:
$L_{\textrm{SM}}=\sum_{\mathbf{x}}\left|\nabla_{u}\tilde{d}(\mathbf{x})\right|e^{-\left|\nabla_{u}\mathbf{I}_{0}(\mathbf{x})\right|}+\left|\nabla_{v}\tilde{d}(\mathbf{x})\right|e^{-\left|\nabla_{v}\mathbf{I}_{0}(\mathbf{x})\right|},$ (25)
where $\nabla_{u}$ and $\nabla_{v}$ refer to the gradient along x and y axis, $\tilde{d}=d/\bar{d}$ is the mean-normalized inverse depth, $M$ represents the set of valid pixels in the reference image.
III-B End-to-end Unsupervised Methods
End-to-end methods [131, 42, 43, 135, 136] are those methods that are trained from scratch with the same input information as supervised methods (Sec. II), while not using ground truth depth for supervision. Instead, they usually adopt photometric consistency, structural similarity and smoothness constraint as parts of their loss terms. Unsup_MVS [131] proposes to inherit the supervision signal of view synthesis, as discussed in Sec. III-A, and dynamically select $X$ “best” (lowest loss) values out of $Y$ loss maps.
However, the bottleneck in unsupervised MVS is to find such accurate photometric correspondences. As illustrated in Fig. 10, in the real scenarios, non-Lambertian surfaces, varying camera exposures, and occlusions will make the assumption of photometric consistency invalid. Hence, such an invalid assumption will lead to “ambiguous supervision”.
To alleviate the ambiguous supervision, JDACS [42] introduces semantic consistency in addition to photometric consistency. It proposes to extract semantic features with a pre-trained network and create semantic classification map by applying non-negative matrix factorization. Then pixel-wise cross-view segmentation consistency is calculated. The difference here is that the semantic map is warped, instead of the image. In this way, the semantic classification maps are supervised by cross-entropy loss between the warped map $S_{i}$ and reference map $S_{0}$, both encoded as pixel-wise one-hot vector. However, the proposed semantic cross-view consistency loss is unstable to converge during training and such high-level semantic information excludes details and boundaries in MVS.
RC-MVSNet [43] digs into providing reliable and occlusion-aware supervision by introducing neural rendering. It takes the advantage of both strong representation ability of Neural Radiance Field [28] and strong generalization ability of cost volume. First, the unsupervised CasMVSNet [32] backbone in RC-MVSNet yields initial depth map supervised by photometric consistency. In addition, it generates depth priors for the Gaussian-Uniform mixture sampling in the rendering-consistency (RC) network. Then after volumetric rendering of the reference view, the depth map is supervised by an extra reference view synthesis loss and a depth rendering consistency loss. In this way, occlusion-aware neural rendering with mixture sampling is utilized to alleviate the invisibility phenomenon and ambiguous supervision from different lighting condition is solved by reference view synthesis.
ElasticMVS [135] introduces an architecture of part-aware patchmatch to address limitations in photometric loss-based geometry information, which tends to have missing data and artifacts in certain areas, particularly textureless regions. The proposed framework incorporates an elastic part representation to encode geometric details for guiding piecewise-smooth depth map prediction. The initial step involves the formal definition of part-aware representation. Building upon this learned representation, enhancements are proposed for both depth hypotheses propagation and evaluation within the context of part-aware patchmatch. Finally, the network is optimized through the incorporation of a dense contrastive loss for self-supervised training, coupled with the integration of a spatial concentration loss [137] designed to promote the isotropic isolation of all pixel embeddings.
CL-MVSNet [136] proposes a framework that aims to enhance proximity among positive pairs by ensuring contrastive consistency between a regular branch and two contrastive branches. The regular branch is a CasMVSNet [32] structure that has been used in  [43, 135, 42] as well, while contrastive branches incorporate an image-level branch and a scene-level branch. These branches enforce contrastive consistency guided by the confidence mask estimated from the regular branch. Additionally, the $L0.5$ photometric consistency is proposed to enhance the accuracy of the reconstruction.
Note that these end-to-end methods are all trained from scratch without any pre-processing which saves training time and reduces the complexity of application in real scenarios.
III-C Multi-stage Unsupervised Methods
Multi-stage methods require either pre-training of a specific module or pre-processing on the training data. These methods are built upon the idea of pseudo-label generation.
Self-supervised CVP-MVSNet [132] proposes to generate pseudo ground truth of depth by training the CVP-MVSNet [75] backbone with photometric consistency loss (Sec. III-A). The pseudo ground truth depth maps are refined with high-resolution images, filtered with cross-view depth consistency check and then fused in a point cloud [15]. Then the mesh is reconstructed with the screened Poisson surface reconstruction algorithm [67] to render complete pseudo depth maps for training in the next iteration. The whole pipeline consists of several iterations to improve the performance. However, [132] cannot provide a satisfying reconstruction of the 3D objects since it heavily relies on the photometric consistency assumption. Additionally, the self-training stage of the whole pipeline takes few days to complete.
U-MVSNet [44] introduces a flow-depth consistency loss by pre-training an optical-flow estimation network, PWC-Net [138], in an unsupervised manner. The dense 2D optical flow correspondences are used to generate pseudo labels for cross-view flow-depth consistency, which alleviates the ambiguous supervision in the foreground. Then the uncertainty-aware self-training consistency further reduces the invalid supervision in the background utilizing the generated pseudo labels and uncertainty map.
Recently, KD-MVS [45] achieves superior performance by leveraging a typical teacher-student scheme of knowledge distillation [139, 140, 141]. It trains a teacher MVS model, e.g., MVSNet [15] or CasMVSNet [32], in a self-supervised manner, by enforcing both photometric and feature-metric consistency between the reconstructed and original images. Then the teacher model produces pseudo ground-truth labels of the training set by cross-view consistency check and probabilistic encoding. In addition to the reliable estimation by the teacher model, the pseudo labels also contain probabilistic knowledge with estimation uncertainty. By training MVS networks [15, 32] with the pseudo labels, the knowledge of the teacher model is distilled to the student models.
III-D Semi-supervised Methods
SGT-MVSNet [46] proposes to use only a few sparse ground truth 3D points to estimate the depth map of the reference view. 3D point consistency loss is used to minimizes the difference between the 3D points back-projected from the corresponding pixels and the ground truth. To tackle the problem of inaccurate estimation on the edge and boundary, a coarse-to-fine reliable depth propagation module rectifies the erroneous predictions.
IV Learning-based MVS without Depth Estimation
Though the learning-based methods that predict individual depth maps with plane-sweep are the main family of learning-based multi-view stereo, there are many methods of other families that achieve impressive 3D reconstruction quality in the recent years. In this section, We discuss four main families: voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
IV-A Voxel-based Methods
These methods [47, 48, 142, 143] estimate the scene geometry with volumetric representation by leveraging implicit function, e.g., SDF. Specifically, Atlas [47] and NeuralRecon [48] attempt to predict the TSDF volume from the 3D feature volume constructed by lifting 2D image features. As shown in Fig. 11, Atlas uses 3D CNN to regress the TSDF volume based on the feature volume accumulated from all images of the scene, which exhibits great completeness of reconstruction. The TSDF reconstruction is supervised using $L_{1}$ loss to the ground truth TSDF vaules. Since dense feature volume brings lots of computational overhead, NeuralRecon further improves the efficiency by incrementally reconstructing the scene in a fragment-wise and coarse-to-fine manner. The 3D features from different fragments are passed through the whole incremental reconstruction process with a RNN. TransformerFusion [142] fuses coarse and fine image features in a voxel grid with two transformers and then predicts an occupancy field to represent the scene geometry. VoRTX [143] uses a similar design to TransformerFusion and the scene geometry is obtained by passing three different-level features output from transformers through a 3D CNN.
Figure 11: Pipeline of Atlas [47]. 2D image features are back-projected into 3D volumes, which are aggregated and passed through a 3D CNN to directly regress a TSDF volume.
IV-B NeRF-based Methods
In novel view synthesis, Neural Radiance Field (NeRF) [28] has kicked off a new emerging representation of 3D, which offers a differentiable volume-rendering scheme to supervise a 3D radiance-based representation with 2D image-level losses. The pipeline of NeRF is shown in Fig. 12. NeRF employs multi-layer perceptron (MLP) to map a position $(x,y,z)$ and the normalized view direction $(\theta,\phi)$ to the corresponding color $\mathbf{c}$ and volume density $\sigma$. For a specific ray at a novel viewpoint, NeRF uses approximated numerical volume rendering to compute the accumulated color as:
$\mathbf{C}=\sum_{i=1}^{N}T_{i}(1-\exp(-\sigma_{i}\delta_{i}))\mathbf{c}_{i},$ (26)
where $i$ is the index of sample, $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ is the accumulated transmittance, and $\delta_{i}=t_{i+1}-t_{i}$ is the distance between adjacent samples. The model is trained by minimizing the loss between the predicted and ground truth color:
$\mathcal{L}_{\text{color}}=\mathbb{E}[(||\mathbf{C}-\mathbf{C}_{gt}||^{2}].$ (27)
Many subsequent endeavors [144, 145, 146, 147, 148, 149, 150, 151, 152] further improve NeRF in quality, fast training, memory efficiency and real-time rendering.
Figure 12: Pipeline of NeRF [28]. Given a 3D position and 2D viewing direction (a), an MLP produces the color and volume density (b). Then volume rendering is used to composite these values into an image (c). The optimization is minimizing the rendering loss (d).
Though the initial purpose of NeRF is to perform novel view synthesis, VolSDF [49] and NeuS [50] integrate NeRF with SDF for surface reconstruction. The SDF, denoted as $f$, is transformed into the density $\sigma$ for volume rendering. Specifically, for a point $\mathbf{p}(t)$, VolSDF computes the volume density $\sigma(\mathbf{p}(t))$ from the signed distance $f(\mathbf{p}(t))$ as:
$\sigma(\mathbf{p}(t))=\frac{1}{\beta}\Psi_{\beta}(-f(\mathbf{p}(t))),$ (28)
where $\Psi_{\beta}$ denotes the Cumulative Distribution Function (CDF) of a zero-mean Laplace distribution with learnable scale parameter $\beta>0$. NeuS computes the density in a different way as:
$\sigma(\mathbf{p}(t))=\max(\frac{-f^{\prime}(\mathbf{p}(t))\Phi^{\prime}_{s}(f(\mathbf{p}(t)))}{\Phi_{s}(f(\mathbf{p}(t)))},0),$ (29)
where $\Phi_{s}$ is the sigmoid function with learnable scale parameter $s$. After training, the mesh can be extracted from the SDF field with Marching Cubes [66].
Note that without ground truth geometry supervision, e.g., depth map or TSDF volume, these methods are trained in a self-supervised way as NeRF, Eq. 27. Since training with rendering loss only like NeRF has ambiguity in geometry [153], some following methods improve the reconstruction with explicit geometry supervision, e.g., monocular depth/normal priors [154, 155] and sparse SfM point cloud [156]. Motivated by Instant-NGP [146] that accelerates training with hash grids, other methods [157, 158, 159, 160] use hash grids to speed up training and improve surface details. In addition, to overcome the drawback that NeRF typically needs to be trained for each new scene, some recent methods [161, 162, 163] propose generalizable pipelines for implicit reconstruction even under sparse-view settings. Specifically, these methods project 3D points on the image planes and aggregate the corresponding image features as a feature volume like Atlas [47], which is used to estimate the surface location.
Recall that the depth map-based MVS methods predict depth with photometric consistency across multiple views. However, this assumption fails for glossy surfaces with reflections and thus they cannot reconstruct them accurately. In contrast, some NeRF methods can handle reflections well. Recently, Ref-NeRF [164] reparameterizes the appearance with separate diffuse and reflective components by using the reflected view direction, which improves the rendering of specular surfaces. Therefore, recent methods [165, 166, 167, 168, 160] adopt this representation in reconstruction and can successfully reconstruct specular surfaces. Specifically, based on VolSDF / NeuS, these methods mainly replace the view direction with the reflected view direction following Ref-NeRF.
IV-C 3D Gaussian Splatting-based Methods
Unlike implicit representations with a coordinate-based MLP such as NeRF [169], 3D Gaussian Splatting [170] (3DGS) explicity represents the scene with point primitives, each of which is parameterized as a scaled Gaussian with 3D covariance matrix $\mathbf{\Sigma}$ and mean $\mathbf{\mu}$:
$G(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{\mu})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{\mu})},$ (30)
where $\mathbf{x}$ is an arbitrary position. $\mathbf{\Sigma}$ is formulated with a scaling matrix $\mathbf{S}$ and rotation matrix $\mathbf{R}$ as:
$\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}.$ (31)
In addition, each Gaussian contains the color $\mathbf{c}$ modeled by Spherical Harmonics and an opacity $\alpha$. Different from NeRF that uses volume rendering, 3DGS efficiently renders the scene via tile-based rasterization. After projecting 3D Gaussian $G(\mathbf{x})$ into the 2D Gaussian $G^{\prime}(x)$ on the image plane [170], a tile-based rasterizer efficiently sorts the 2D Gaussians and employs $\alpha$-blending for rendering:
$\mathbf{C}(x)=\sum_{i\in N}\mathbf{c}_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\sigma_{j}),\quad\sigma_{i}=\alpha_{i}G^{\prime}(x),$ (32)
where $x$ is the pixel location, $N$ is the number of sorted 2D Gaussians. During training, 3DGS minimizes the rendering loss like NeRF, as in Eq. 27. 3DGS can be initialized with SfM [171, 17] or MVS [34], which performs better than random initialization [170].
Motivated by the NeRF-based MVS methods, researchers try to adapt 3DGS for reconstruction. Though 3DGS achieves high-quality novel-view synthesis, it is challenging to recover high quality geometry since no explicit geometry constraint is used and 3D Gaussians do not correspond well to the actual surface because of the 3D covariance [52]. As one of the earliest methods, DreamGaussian [51] follows NeRF [28] to compute a dense density grid with the 3D Gaussians and then extracts mesh with Marching Cubes [66]. SuGaR [52] introduces a geometry regularization term encouraging the 3D Gaussians to be well-aligned over the scene surfaces so that the Gaussians can contribute to better scene geometry. To reconstruct the mesh, SuGaR samples 3D points on a level set of the density computed from 3D Gaussians and then runs Poisson Reconstruction [172] on these points. NeuSG [173] jointly trains Neuralangelo [158], a NeRF-based method, and 3DGS. During training, the normals from the SDF field of Neuralangelo regularize the rotation of 3D Gaussians, while the SDF field is regularized to ensure that the SDF values at 3D Gaussians’ positions are close to zero. However, retrieving good geometry from Gaussian-based representations still needs to be explored.
IV-D Large Feed-forward Methods
In 3D reconstruction and 3D generation, a recent prominent trend is to directly learn the 3D representation from large-scale 3D datasets, such as Objaverse [30]. These methods typically adopt large-scale transformers [174] to generate a 3D representation based on various inputs, including single image [53, 175, 176], textual description [177], posed multi-view images [178], and un-posed multi-view images [179, 54]. LRM [53] uses a transformer-based feed-forward network to regress the 3D object from a single image. Specifically, LRM predicts features of a tri-plane representation [180], which is subsequently converted into an implicit field like NeRF [53]. Then the surface is extracted with Marching Cubes [66].
Recently, DUSt3R [54] reconstructs the scene as a point cloud given a pair of un-calibrated and un-posed images. After extracting image token representations with Vision Transformer [181], two transformer decoders exchange information of the token representations via cross-attention and then regress the per-pixel 3D points in the coordinate frame of the first image. Many downstream tasks can be performed with the point clouds, e.g., point matching, localization, intrinsic and relative pose estimation.
V Discussions
In this section, we summarize and discuss the performance of learning-based MVS methods, including supervised online methods with depth estimation (Sec. V-A), supervised offline methods with depth estimation (Sec. V-B), unsupervised methods with depth estimation (Sec. V-C) and methods without depth estimation (Sec. V-D). Moreover, we discuss potential directions for future research (Sec. V-E).
TABLE II: Quantitative results of supervised online MVS methods on ScanNet [26] and 7-Scenes [27].
Methods ScanNet [26] 7-Scenes [27]
abs $\downarrow$ abs-rel $\downarrow$ abs-inv $\downarrow$ $\eta<1.25$uparrow abs $\downarrow$ abs-rel $\downarrow$ abs-inv $\downarrow$ $\eta<1.25$uparrow
MVDepthNet [25] 0.167 0.087 0.054 0.925 0.201 0.117 0.071 0.877
DPSNet [97] 0.219 0.119 0.071 0.868 0.249 0.149 0.085 0.826
Neural-RGBD [87] 0.236 0.122 0.075 0.850 0.214 0.131 0.076 0.865
GP-MVS [58] 0.149 0.076 0.049 0.940 0.174 0.100 0.064 0.903
Long et al. [94] 0.151 0.081 - 0.931 0.253 0.147 - 0.804
DeepVideoMVS [31] 0.119 0.060 0.038 0.965 0.145 0.038 0.054 0.938
MaGNet [88] 0.147 0.081 - 0.930 0.213 0.126 - 0.855
RIAV-MVS [111] 0.139 0.075 - 0.938 0.178 0.100 - 0.897
SimpleRecon [35] 0.089 0.043 - 0.981 0.105 0.058 - 0.974
TABLE III: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for unsupervised depth map-based MVS methods.“Finetuned with BlendedMVS” denotes whether the methods are finetuned on BlendedMVS [24] before evaluating on Tanks and Temples [22].
Methods DTU [21] Finetuned with Tanks and Temples [22]
Acc. $\downarrow$ Comp. $\downarrow$ Overall $\downarrow$ BlendedMVS [24] Intermediate $F_{1}$uparrow Advanced $F_{1}$uparrow
End-to-end Unsup_MVSNet [131] 0.881 1.073 0.977 - -
MVS2 [129] 0.760 0.515 0.637 37.21 -
M3VSNet [130] 0.636 0.531 0.583 37.67 -
JDACS-MS [42] 0.398 0.318 0.358 45.48 -
RC-MVSNet[43] 0.396 0.295 0.345 55.08 30.82
ElasticMVS[135] 0.374 0.325 0.349 57.88 37.81
CL-MVSNet[136] 0.375 0.283 0.329 59.39 37.03
Multi-stage Self-supervised CVP-MVSNet [132] 0.308 0.418 0.363 43.48 -
U-MVSNet [44] 0.354 0.354 0.354 57.15 30.97
KD-MVS [45] 0.359 0.295 0.327 64.14 37.96
TABLE IV: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for supervised offline MVS methods. “Finetuned with BlendedMVS” denotes whether the learning-based methods are finetuned on BlendedMVS [24] before evaluating on Tanks and Temples [22] and ETH3D [23].
Methods DTU [21] Finetuned with Tanks and Temples [22] ETH3D [23]
Acc. $\downarrow$ Comp. $\downarrow$ Overall $\downarrow$ BlendedMVS [24] Intermediate $F_{1}$uparrow Advanced $F_{1}$uparrow Training $F_{1}$uparrow Test $F_{1}$uparrow
Direct 3D CNN MVSNet [15] 0.396 0.527 0.462 43.48 - - -
P-MVSNet [182] 0.406 0.434 0.420 55.62 - - -
CIDER [36] 0.417 0.437 0.427 46.76 23.12 - -
PointMVSNet [95] 0.342 0.411 0.376 - - - - -
VA-Point-MVSNet [116] 0.359 0.358 0.359 48.70 - - -
PVA-MVSNet [90] 0.379 0.336 0.357 54.46 - - -
Fast-MVSNet [117] 0.336 0.403 0.370 47.39 - - -
RNN R-MVSNet [37] 0.383 0.452 0.417 48.40 24.91 - -
$D^{2}$-RMVSNet [38] 0.395 0.378 0.386 59.20 - - -
AA-RMVSNet [77] 0.376 0.339 0.357 61.51 - - -
BH-RMVSNet [183] 0.368 0.303 0.335 61.96 34.81 - 79.61
Coarse-to-fine CasMVSNet [32] 0.325 0.385 0.355 56.84 - - -
CVP-MVSNet [75] 0.296 0.406 0.351 54.03 - - -
UCS-Net [39] 0.338 0.349 0.344 54.83 - - -
AttMVS [184] 0.383 0.329 0.356 60.05 37.34 - -
Vis-MVSNet [92] 0.369 0.361 0.365 60.03 - - -
EPP-MVSNet [100] 0.413 0.296 0.355 61.68 35.72 74.00 83.40
CDS-MVSNet [80] 0.351 0.278 0.315 61.58 - - -
TransMVSNet [79] 0.321 0.289 0.305 63.52 37.00 - -
GBi-Net [78] 0.315 0.262 0.289 61.42 37.32 - -
UniMVSNet [114] 0.352 0.278 0.315 64.36 38.96 - -
NP-CVP-MVSNet [185] 0.356 0.275 0.315 59.64 - - -
MVSTER [99] 0.340 0.266 0.303 60.92 37.53 72.06 79.01
PVSNet [63] 0.337 0.315 0.326 59.11 35.51 76.57 82.62
MVSFormer [82] 0.327 0.251 0.289 66.37 40.87 - -
IS-MVSNet [186] 0.351 0.359 0.355 62.82 34.87 73.33 83.15
HR-MVSNet [187] 0.332 0.310 0.321 63.12 34.27 - -
EPNet [121] 0.299 0.323 0.313 63.68 40.52 79.08 83.72
GeoMVSNet [76] 0.331 0.259 0.295 65.89 41.52 - -
RA-MVSNet [188] 0.326 0.268 0.297 65.72 39.93 - -
DMVSNet [115] 0.338 0.272 0.305 64.66 41.17 - -
ET-MVSNet [86] 0.329 0.253 0.291 65.49 40.41 - -
Iterative update PatchmatchNet [40] 0.427 0.277 0.352 53.15 32.31 64.21 73.12
PatchMatch-RL [103] - - - 51.80 31.80 67.80 72.40
IterMVS [41] 0.373 0.354 0.363 56.94 34.17 71.69 80.09
Effi-MVS [102] 0.321 0.313 0.317 56.88 34.39 - -
CER-MVS [101] 0.359 0.305 0.332 64.82 40.19 - -
IGEV-MVS [112] 0.331 0.326 0.324 - - - - -
V-A Supervised Online MVS with Depth Estimation
Typically, online MVS methods are trained on ScanNet [26] and then evaluated on ScanNet [26] and 7-Scenes [27]. We summarize the quantitative results of online MVS methods in Tab. II. By incorporating temporal information in depth estimation,  [58, 94, 31] achieve better performance than MVDepthNet [25]. However, this usually increase the network complexity. Instead of temporal fusion, SimpleRecon [35] injects cheaply available metadata into the cost volume and achieves state-of-the-art performance on both ScanNet and 7-Scenes.
V-B Supervised Offline MVS with Depth Estimation
Offline MVS methods are usually trained on DTU [21] and then evaluated on DTU [21], Tanks and Temples [22] and ETH3D [23]. Recently, many methods further finetune the DTU-pretrained model on BlendedMVS [24] before evaluating on Tanks and Temples and ETH3D since it contains large-scale scenes.
V-B1 Benchmark Performance
Tab. IV summarizes quantitative results of offline MVS methods. Compared with those methods that use direct 3D CNN for regularization, the methods that use RNN and coarse-to-fine regularization perform much better. Since RNN based methods have bad time efficiency, we can find that coarse-to-fine methods become main stream of the community because of their impressive performance and high efficiency in both memory and run-time. In addition, iterative methods based on traditional PatchMatch [40, 103] or RAFT [41, 102, 101] also become popular since they achieve comparable performance as state-of-the-art methods with more lightweight structure. For benchmarks, most learning-based methods perform explicitly better than traditional methods on DTU and Tanks and Temples. However, on ETH3D, traditional methods [189, 127] perform better than all existing learning-based methods. We conjecture that the strong viewpoint variations and large textureless regions in human-made environments still make learning-based methods struggle. Therefore, it is valuable to pay more attention to ETH3D to evaluate the robustness of learning-based methods in real-world scenes.
V-B2 Memory Consumption and Run-time
Low run-time and memory consumption are essential in most industrial applications with limited computational power and storage, e.g., autonomous driving and AR/VR. Because of the low-resolution input and the simplicity of network structures, online MVS methods [25, 31, 35] usually achieve high efficiency in both memory and run-time. In contrast, offline MVS methods focus on high-resolution images and usually use computationally expensive network modules, e.g., 3D convolutions, to improve reconstruction quality. These increase run-time and GPU memory. Recently, many researchers try to improve the efficiency while maintaining the reconstruction accuracy as other offline MVS methods. We compare and summarize the efficiency of existing efficient methods [117, 32, 39, 40, 102, 41, 99] in Fig. 13. All the methods share similar coarse-to-fine structures to improve efficiency. To further reduce run-time and memory, [40, 102, 41, 99] further get rid of costly 3D convolution and use 2D convolution instead. It can be found many methods [40, 102, 41, 99] can achieve near real-time estimation (about 10 Hz) with relatively high resolution images, e.g., $1152\times 864$. In conclusion, it is promising to further improve both the performance and efficiency for practical applications.
Figure 13: GPU memory consumption and run-time of efficient offline methods [117, 32, 39, 40, 102, 41, 99] on DTU dataset [21]. The image resolution and the number of images are set to $1152\times 864$ and 5 respectively. Experiments are done on a NVIDIA 2080Ti GPU.
V-C Unsupervised MVS with Depth Estimation
We summarize the results of unsupervised MVS methods in Tab. III. Based on the conventional photometric consistency loss (Sec. III-A), unsupervised MVS methods may have difficulties with many challenging situations, e.g., varying lighting conditions or occlusions, and suffer from downgraded reconstruction results. Therefore, many methods propose to employ different strategies to improve the robustness, e.g., segmentation consistency [42], optical flow consistency [44], neural rendering consistency [43], pseudo-label generation with geometric filtering [132, 45], multi-iteration training [132, 45] and featuremetric loss [45]. It is inspiring that KD-MVS [45] already outperforms its backbone, CasMVSNet [32], and many other fully supervised methods. However, since unsupervised methods mainly test on relatively simple scenes [21, 22], their scalability to more complex and large-scale scenes, e.g., ETH3D [23], remains a question. This good generalization capability across various scenes and datasets is crucial for practical applications. Therefore, evaluating unsupervised MVS methods on various scenes is an important task for future research. Most unsupervised methods use simple MVS backbones [15, 32, 75] and mainly focus on the training strategy. Employing state-of-the-art architectures is also a potential direction that can further boost the performance.
V-D Learning-based MVS without Depth Estimation
V-D1 Voxel-based Methods
We summarize the quantitative results on ScanNet [26] in Tab. V. Since the voxel representation increases the computation overhead explicitly when the scale of scene increases, voxel-based methods mainly focus on indoor scenes, e.g., ScanNet [26]. We find that depth map-based MVS methods, e.g., SimpleRecon [35], fuse depth maps with traditional TSDF fusion [64] and can achieve comparable performance as these voxel-based methods with lower complexity. This further shows the advantages of depth map-based methods over voxel-based methods.
TABLE V: Quantitative results of voxel-based MVS methods on ScanNet [26].
Methods Prec. $\uparrow$ Recall $\uparrow$ F-score $\uparrow$
Atlas [47] 0.675 0.605 0.636
NeuralRecon [48] 0.630 0.612 0.619
TransformerFusion [142] 0.728 0.600 0.655
VoRTX [143] 0.767 0.651 0.703
TABLE VI: Quantitative results of NeRF-based MVS methods on DTU [21].
Methods VolSDF [49] NeuS [50] NeuralWarp [133] HF-NeuS [190] RegSDF [191] PET-NeuS [192]
Overall $\downarrow$ 0.86 0.87 0.68 0.77 0.72 0.71
Methods GeoNeuS [156] Voxurf [193] NeuS2 [157] PermutoSDF [159] Neuralangelo [158] UniSDF [160]
Overall $\downarrow$ 0.51 0.72 0.70 0.68 0.61 0.64
V-D2 NeRF-based Methods
The quantitative results of nerf-based methods on DTU dataset [21] are summarized in Tab. VI. We visualize the qualitative reconstructions of two representative methods, NeuS [50] and VolSDF [49], in Fig. 14. The most outstanding features of NeuS [50] and VolSDF [49] are that they are self-supervised, i.e., without depth supervision, and produce smooth and complete mesh surfaces. However, many recent methods [156, 154, 155] find that incorporating more explicit geometric supervision explicitly improves the reconstruction quality. Compared with depth map-based MVS methods [25, 15], the main drawback of NeRF-based methods is that they are mostly per-scene optimization problems and need lots of time to train the model on each new scene. Recently, some methods [161, 162, 163] try to make the pipeline generalizable so that there is no need to train the model for new scenes. However, they perform worse than depth map-based methods [15]. Another way to improve training time efficiency is to employ hash grids [146], which also improves surface details. However, these methods [157, 158, 159, 160] need lots of GPU memory and space to train and store the large model. In addition, the scalability of NeRF-based methods remains a problem when attempting to apply on large-scale scenes [23].
Figure 14: Reconstruction results (rendered surface normal map and surface mesh) of NeuS [50] and VolSDF [49] on DTU dataset [21]. NeuS and VolSDF apply different approximation of density with SDF.
V-D3 3D Gaussian Splatting-based Methods
Because of its highly efficient rasterization, 3D Gaussian Splatting [170] becomes popular in novel view synthesis and gradually replaces NeRF [28] in the last half year. Recently, some researchers start exploring the potential of 3D Gaussian Splatting in surface reconstruction. However, there are many challenges. For example, the 3D Gaussians do not correspond well to the actual surface of the scene since the 3D Gaussian has a 3D covariance matrix [52]. Moreover, the surface may contain noisy undulations [52]. However, it is worthwhile to further explore in this direction since 3D Gaussian Splatting is much faster to train than those NeRF-based MVS methods. Moreover, instead of directly extracting the surface from 3D space [52], rendering accurate depth maps and then fuse them like depth map-based MVS methods is a potential direction.
V-D4 Large Feed-forward Methods
As a new trend of 3D reconstruction, large feed-forward methods demonstrate their impressive performance in challenging settings that other MVS methods cannot handle, e.g., single image input and un-posed sparse images. Employing a large model that is able to get scaled up, these methods manage to learn strong priors of multi-view geometry from large datasets [30]. However, these methods [53, 179, 54] face limitations when it comes to extending the scale of the scene. For example, the 3D representation employed in LRM [53] is a tri-plane representation [180], which has a small resolution and struggles with scaling to larger sizes due to memory constraints. In addition, the tri-plane representation is difficult to model more complex geometry, e.g., with strong self-occlusions. In addition to tri-plane-based 3D representations, new methods [194, 195] have been proposed that directly reconstruct 3D Gaussians in a feed-forward manner. These new methods demonstrate improved capabilities in reconstructing both objects and scenes, but require posed sparse images as inputs. It is worth noting that the generalization ability of these methods is questionable since they learn priors of multi-view geometry completely from the training dataset, while the depth map-based methods utilize feature matching from plane sweep algorithm from traditional methods. Moreover, these methods typically need high-end GPUs since they are usually GPU memory consuming. It is worth noting that, since large feed-forward methods usually do not enforce geometric constraint explicitly, e.g., epipolar geometry, the reconstructed geometry is by default placed in a local coordinate system without absolute scales, making it not feasible for applications requiring precise measurement.
V-E Future Research Directions
V-E1 Datasets and Benchmarks
For learning-based MVS methods, ScanNet [26] and DTU [21] are two main training datasets, while ScanNet [26], DTU [21], Tanks & Temples [22] and ETH3D [23] are the main evaluation benchmarks.
For training, the scene scale of ScanNet and DTU is relatively small (room-scale for ScanNet and object-scale for DTU) and their quality is not satisfactory. For example, the camera calibration of ScanNet is not very accurate. For DTU, the ground truth depth maps are rendered from mesh, which contain some outliers and holes. This is because the mesh is reconstructed from the sparse ground truth point cloud with Screened Poisson surface reconstruction [67] and thus contains incomplete regions [184]. Therefore, improving the scalability and quality of training datasets is an important research direction.
Recently, there are many researchers trying to solve this problem. TartanAir dataset [196] is a large-scale synthetic dataset that is collected in photo-realistic simulation environments. Precise multi-modal sensor data and ground truth labels are provided, including the stereo RGB image, depth image, segmentation, optical flow, camera poses, and LiDAR point cloud. BlendedMVS [24] introduces more large-scale scenes and improves the performance of many MVS methods on real-world scenes, shown in Tab. IV. ArKitScenes [197] consists of 5,048 RGB-D sequences, which is more than three times the size of the current largest available indoor dataset, ScanNet. As an extension of ScanNet, ScanNet++ [198] is a large-scale dataset that captures high-quality geometry and color of indoor scenes with high-end laser scanner, DSLR camera, and RGB-D streams from an iPhone. Motivated by ImageNet [199] that drives a remarkable trend of learning from large-scale data in 2D visual tasks, Yu et al. [200] propose MVImgNet, a large-scale dataset of multi-view images collected by shooting videos of real-world objects. MVImgNet demonstrates the potential of various 3D visual tasks, including radiance field reconstruction, multi-view stereo, and view-consistent image understanding. Similarly, Objaverse [30] is a large dataset of objects with 800K+ (and growing) 3D models with descriptive captions, tags, and animations. Both MVImgNet [199] and Objaverse [30] can be used to train large feed-forward models, e.g., LRM [53], for reconstruction, since these models need massive data to learn the strong prior. However, both of them mainly focus on object-level scenes. Large datasets for larger-scale scenes, e.g., rooms, are valuable to be explored.
For evaluation, the main benchmarks [26, 21, 22, 23] are all introduced before 2018. Some of these benchmarks, e.g.,  [21, 22], are currently saturated and the learning-based methods are difficult to further improve the performance on them. Therefore, it is meaningful to introduce new benchmarks to evaluate the robustness and performance of learning-based methods in challenging real-world scenes. For example, LaMAR [201] is a recent large-scale dataset captured using multiple modern AR devices in diverse environments. It contains challenging short-term appearance and structural changes and high quality LiDAR point cloud ground truth. These make LaMAR a potential benchmark for MVS.
V-E2 View Selection
As discussed in Sec. I-B, view selection is crucial for triangulation quality. Picking neighboring views that are suitable for triangulation can not only improve the reconstruction accuracy but also reduce useless computation for the bad views, e.g., views with strong occlusions. However, view selection is often overlooked and not well studied. For example, all offline depth map-based MVS methods [32, 40, 79, 99] follow MVSNet [15] and use the same simple heuristic strategy to compute scores for neighboring views and sort them. Online depth map-based MVS methods [58, 31, 35] and voxel-based methods [48] also adopt heuristic strategies to choose views with enough pose-distance. Though it is intractable to incorporate non-differentiable view selection into deep learning, it is worth exploring new view selection strategies since it may improve the reconstruction without changing the model design of existing methods. Current view selection strategies simply pick the same set of neighboring views for all the pixels in the reference image, and it is probable that different reference pixels have different optimal choices of neighboring views. If it is possible to choose the best neighboring views for each pixel, reconstruction quality can be further improved and computation reduced. The learned pixel-wise view weight [63, 40, 41] is one solution for this since it weights the source views differently for each pixel based on its visibility across neighboring views.
V-E3 Depth Fusion
As discussed in Sec. I-D, both online and offline MVS methods with plane-sweep use traditional TSDF fusion or depth filtering methods to reconstruct mesh or point cloud from the estimated depth maps. However, TSDF fusion [64, 65] is not robust enough to the outliers in the depth maps and may have memory issues due to the dense volumetric representation. Depth filtering following Galliani et al. [18] introduces lots of hyper-parameters. Researchers may carefully finetune these hyper-parameters for each scene since it may have a great impact on the evaluation metrics. Therefore, it is valuable to improve the depth fusion step so that the reconstruction quality can be further improved. Recently, there have been some learning-based depth map fusion methods [202, 203]. However, they are limited to small scenes due to the dense volume representation and large memory consumption.
V-E4 Features
One not well-studied topic is what kind of feature extractors are suitable for MVS, as mentioned in Sec. II-A. So far, most of the offline depth map-based methods [15, 32, 78] mainly apply simple 2D CNN and FPN [72] structure as a feature extractor. To improve the receptive fields flexibly, Deformation Convolution [204] is used  [78, 77, 40, 79]). The attention mechanism is also applied in feature learning, where works such as  [90, 79, 77] use intra-attention or inter-attention to capture long-range feature dependencies. Furthermore, MVSFormer [82] digs deeper into using patch-wise ViT [181] as a feature extractor in MVS. The insight is that ViT works better to formulate global feature correlations, and FPN can learn detailed ones. In contrast, online depth map-based methods [31, 35] usually adopt efficient pretrained backbones [73, 70, 74]. Recently, large feed-forward methods [53, 54, 175] adopt large pretrained feature encoders, e.g., ViT [181], DINO [83], to learn strong prior from large-amount of data. However, using a powerful feature extraction network usually increases the computation overhead and reduces the efficiency. Finding a good balance between performance and efficiency is an interesting direction, e.g., distilling the large ViT into a relatively small transformer [205].
V-E5 Real-time and Memory Efficient Models
Though reconstruction accuracy is considered the most important factor in MVS, it is also critical to have models that can run in near real-time and with low memory, e.g., AR/VR and autonomous driving. Comparatively, depth map-based methods are most suitable for achieving efficiency because of their conciseness, flexibility, and scalability. Though online MVS methods can achieve real-time estimation with images of low resolutions, they may have issues with memory consumption since many large backbones are usually used for feature extraction [31, 35]. In addition, the efficiency in both run-time and memory will drop when image resolution increases. For offline MVS methods, as discussed in Sec. V-B2, some recent methods try to improve efficiency with carefully simplified network architectures. However, this usually limits the performance when compared with state-of-the-art methods. There are many other directions for further improvement. For example, using model compression techniques and knowledge distillation are potential directions to not only keep the good performance and also improve efficiency.
V-E6 Prior Assistance
MVS mainly relies on measurements of local photometric consistency to find the optimal matching across reference and source images. Accordingly, it usually encounters difficulties when estimating the geometry for regions where the photometric measurement becomes unreliable or invalid, e.g., textureless areas and non-Lambertian surfaces, which are common in human-made environments, e.g., ETH3D [23]. Therefore, leveraging prior information to guide the MVS algorithm in these challenging regions is a promising research direction. We elaborate several typical examples of prior assistance as follows.
  • Surface Normal: It is a non-local representation of the geometry compared with the depth map. To enforce the constraints of normal maps in depth estimation, Kusupati et al. [206] integrate the multi-view normal estimation network into the MVS pipeline. Long et al. [207] introduce a Combined Normal Map, estimated with PlaneCNN [208] to enforce the normal consistency on the depth map. Liu et al. [209] perform graph-based depth map optimization as post-processing. Since normal monocular normal predictions [210] provide high-quality priors for the full scene, [154, 155] use monocular normal to improve the optimization of neural implicit surfaces.
  • Shape Prior: For indoor scenes where common textureless areas, e.g., walls, planes are suitable choices of the geometric primitives and are exploited in traditional methods [20]. Recently, PlaneMVS [211] and PlanarRecon [212] explicitly estimate plane parameters for depth refinement or holistic reconstruction. In addition, predicting object-level attributes simultaneously when estimating depth [213] is also a feasible solution for specific applications like urban modeling.
  • Semantic Segmentation: Intuitively, points assigned with the same semantic labels may be more likely to lie on the same 3D plane. Some attempts [214, 215] employ a rule-based protocol to utilize semantic segmentation for better matching and depth quality. Manhattan-SDF [216] adopts Manhattan world priors to handle low-textured planar regions corresponding to walls and floors with 2D segmentation.
VI Conclusion
In this survey, we have provided a comprehensive review of the learning-based MVS methods until 2023, which are categorized into: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. Particularly, we have devoted significant attention to depth map-based methods because of their conciseness, flexibility and scalability. We explain key aspects such as datasets utilized, general working pipelines, and algorithmic intricacies. Furthermore, we have summarized and compared the quantitative performance of these methods across popular benchmarks, providing valuable insights into their efficacy and applicability. Finally, our discourse extends to an in-depth exploration of potential research directions for the future of MVS, highlighting avenues for further investigation and innovation in the field.

### v2

@@
 Index Terms:
-Multi-View Stereo, 3D Reconstruction, Deep Learning.
+3D Reconstruction, Multi-View Stereo, Deep Learning.
+1 Introduction
 3D reconstruction describes the general task of recovering the 3D structure of a scene.
-It is widely applied in augmented/virtual reality (AR/VR), autonomous driving, and robotics [1].
+It is widely employed in augmented/virtual reality (AR/VR), autonomous driving, and robotics [1].
 The advancement of 3D acquisition techniques has led to the increased affordability and reliability of depth sensors, such as depth cameras and LiDARs.
@@
 One fundamental technique in image-based 3D reconstruction is Multi-View Stereo (MVS).
-Given a set of calibrated images, MVS aims to reconstruct the dense 3D geometry for an observed scene.
+Given a set of calibrated images, MVS aims to reconstruct dense 3D geometry for an observed scene.
 Based on the scene representations, traditional MVS methods can be divided into three categories: volumetric, point cloud, and depth map.
-Volumetric methods [9, 10, 11, 12] discretize the continuous 3D space into voxels and label each as inside or outside of the surface.
-They are limited to scenes of small scale because of the large memory consumption.
+Volumetric methods [9, 10, 11, 12] discretize the 3D space into voxels and label each as inside or outside of the surface.
+They are limited to scenes of small scale due to the large memory consumption.
 Point cloud methods [13, 14] operate directly on 3D points and often employ propagation to gradually densify the reconstruction.
-As the propagation of point clouds is proceeded sequentially, these methods are difficult to be parallelized and thus take much time in processing [15].
+As point cloud propagation occurs sequentially, these methods are difficult to parallelize, leading to longer processing times.
+[15].
 In addition, the irregularity and large size of point cloud are also not very suitable for deep learning, especially in large-scale scenes.
@@
 Although MVS has been studied extensively for several decades, traditional MVS methods rely on hand-crafted matching metrics and thus encounter challenges in handling various conditions, e.g., illumination changes, low-textured areas, and non-Lambertian surfaces [21, 22, 23, 24].
-To overcome these challenges, recent works [15, 25] have turned to learning-based methods, leveraging the emergence of Convolutional Neural Networks (CNNs) that have shown remarkable success in various 2D vision tasks.
-These learning-based MVS techniques have achieved remarkable success and explicitly outperform traditional methods [18, 17] on various benchmarks [21, 22, 23, 26, 27].
+To overcome these challenges, recent works [15, 25] have shifted towards learning-based approaches, have adopted learning-based approaches, using convolutional neural networks (CNNs) that have excelled in 2D vision tasks.
+These methods have significantly outperformed traditional methods on various benchmarks [21, 22, 23, 26].
 In this survey, we categorize existing learning-based MVS methods based on their characteristics as follows: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
@@
 Comparatively, depth map-based methods [15, 25] are the main family of learning-based methods since they inherit the advantages from those traditional methods [17, 18, 19, 20] by decoupling 3D reconstruction into depth estimation and depth fusion.
-Therefore, we focus more on the depth map-based methods and discuss them in more details in this survey.
+Therefore, we focus more on the depth map-based methods and discuss them in details.
 For clarity, we further categorize depth map-based methods into online and offline methods, shown in Fig.
 1.
-Specifically, online MVS methods [25, 31] usually perform multi-view depth estimation with a video sequence of low-resolution images, e.g., ScanNet [26].
-They typically use small resolution inputs and a lightweight network structure to ensure quick inference considering speed and simplicity.
-The primary focus of online MVS is to provide instant reconstruction for applications where time efficiency is critical, e.g., augmented reality and live video processing.
+Specifically, online MVS methods [25, 30] usually perform multi-view depth estimation with a video sequence of low-resolution images, e.g., ScanNet [26].These methods typically rely on low-resolution inputs and lightweight network structures to ensure fast inference, prioritizing speed and simplicity.
+The main goal of online MVS is to deliver real-time reconstruction for time-sensitive applications such as augmented reality and live video processing.
 In contrast, offline MVS methods [15, 31] prioritize better quality reconstruction at the cost of computation.
 These methods excel in multi-view depth estimation and point cloud fusion using high-resolution image sets, e.g., DTU [21], Tanks and Temples [22] and ETH3D [23].
-By operating on higher-resolution inputs and employing complex network designs, these approaches thoroughly analyze the multi-view information in the images.
+By operating on higher-resolution inputs and employing complex network, they thoroughly analyze the multi-view information in the images.
 The emphasis on achieving high accuracy and capturing intricate details often comes at the cost of requiring considerable computational resources and time.
 As a result, offline MVS is frequently employed in applications demanding precise and photorealistic scene representations, including 3D modeling and archaeological reconstruction.
-In summary, our survey covers the most recent literature on learning-based MVS methods until 2023, including four main families: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
+In summary, our survey covers the most recent literature on learning-based MVS methods, including four main families: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
 We provide a comprehensive review and insights on different aspects, including the pipelines and algorithmic intricacies.
@@
 Given a full set of images, offline MVS holds the global information of the captured scene, and usually fuses estimated depth maps into a point cloud with filtering.
-I Preliminaries
+2 Preliminaries
 Depth map-based MVS, including most traditional and learning-based methods, typically consists of several components: camera calibration, view selection, multi-view depth estimation and depth fusion.
-In this section, we introduce these components to provide readers a clear picture of the MVS problem.
+In this section, we introduce these components to provide readers a clear picture of the MVS.
 Note that camera calibration and view selection are components for other learning-based methods as well.
 Sec.
-I-A introduces camera calibration with Structure from Motion (SfM) or SLAM.
+2.1 introduces camera calibration with Structure from Motion (SfM) or SLAM.
 Sec.
-I-B discusses how to select neighboring views to accurately reconstruct the geometry.
+2.2 discusses how to select neighboring views to reconstruct the geometry.
 Sec.
-I-C explains how to build cost volumes in learning-based MVS methods with plane sweep [33].
+2.3 explains how to build cost volumes in learning-based MVS methods with plane sweep [32].
 Sec.
-I-D introduces the typical depth fusion strategies after depth estimation.
+2.4 introduces the typical depth fusion strategies after depth estimation.
 Sec.
-I-E lists common datasets and benchmarks for MVS and Sec.
-I-F summarizes the common evaluation metrics.
+2.5 lists common datasets and benchmarks for MVS and Sec.
+2.6 summarizes the common evaluation metrics.
 {forest}
-for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, draw=black, rounded corners,align=left, minimum width=2.5em, minimum height=1.2em, s sep=6pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=8em,font=, where level=3font=, where level=4font=, where level=5font=, [MVS [Traditional Methods,text width=10 em [COLMAP [34], Furu [14], Gipuma [18]…,leaf,text width=15 em] ] [Datasets & Benchmarks, text width=10 em [ScanNet [26], DTU [21], Tanks and Temples [22]…, leaf,text width=15 em] ] [Learning-based Methods
-with Depth Estimation,text width=10 em [Pipeline,text width=6 em [Camera Calibration, View Selection, Multi-View Depth Estimation, Depth Fusion,leaf,text width=25.3em] ] [Supervised,text width=6em [Online Methods [MVDepthNet [25], DeepvideoMVS [31], SimpleRecon [35]…,leaf,text width=18.8 em] ] [Offline Methods [Direct 3D CNN,text width=5 em [MVSNet [15], CIDER [36]…,leaf,text width=12 em] ] [RNN,text width=5 em [R-MVSNet [37], $D^{2}$-RMVSNet [38]…,leaf,text width=12 em] ] [Coarse-to-fine,text width=5 em [CasMVSNet [32], UCSNet [39]…,leaf,text width=12 em] ] [Iterative Update,text width=5 em [PatchmatchNet [40], IterMVS [41]…,leaf,text width=12 em] ] ] ] [Unsupervised,text width=6em [End-to-end,text width=5 em [ JDACS [42], RC-MVSNet [43]…,leaf,text width=10 em] ] [Multi-stage,text width=5 em [U-MVSNet [44], KD-MVS [45]…,leaf,text width=10 em] ] ] [Semi-Supervised,text width=6em [ SGT-MVSNet [46],leaf,text width=6.8 em] ] ] [Learning-based Methods
-without Depth Estimation,text width=10em [Voxel-based,text width=10 em [Atlas [47], NeuralRecon [48]…, leaf,text width=13 em] ] [NeRF-based,text width=10 em [VolSDF [49], NeuS [50]…, leaf,text width=13 em] ] [3D Gaussian Splatting-based,text width=10 em [DreamGaussian [51], SuGaR [52]…, leaf,text width=13 em] ] [Large feed-forward,text width=10 em [LRM [53], DUSt3R [54]…, leaf,text width=13 em] ] ] [Future Research Directions, text width=9.2 em [Dataset & Benchmarks, View Selection, Depth Fusion, Features, Run-time & Memory Efficiency, Prior Assistance, leaf,text width=33.6 em] ] ]
+for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, draw=black, rounded corners,align=left, minimum width=2.5em, minimum height=1.2em, s sep=6pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=8em,font=, where level=3font=, where level=4font=, where level=5font=, [MVS [Traditional Methods,text width=10 em [COLMAP [33], P-MVS [14], Gipuma [18]…,leaf,text width=16.3 em] ] [Datasets & Benchmarks, text width=10 em [ScanNet [26], DTU [21], Tanks and Temples [22]…, leaf,text width=16.3 em] ] [Learning-based Methods
+with Depth Estimation,text width=10 em [Pipeline,text width=6 em [Camera Calibration, View Selection, Multi-View Depth Estimation, Depth Fusion,leaf,text width=27.4 em] ] [Supervised,text width=6em [Online Methods [MVDepthNet [25], DeepvideoMVS [30], SimpleRecon [34]…,leaf,text width=20.2 em] ] [Offline Methods [Direct 3D CNN,text width=5.4 em [MVSNet [15], CIDER [35]…,leaf,text width=13 em] ] [RNN,text width=5.4 em [R-MVSNet [36], $D^{2}$-RMVSNet [37]…,leaf,text width=13 em] ] [Coarse-to-fine,text width=5.4 em [CasMVSNet [31], UCSNet [38]…,leaf,text width=13 em] ] [Iterative Update,text width=5.4 em [PatchmatchNet [39], IterMVS [40]…,leaf,text width=13 em] ] ] ] [Unsupervised,text width=6em [End-to-end,text width=5 em [ JDACS [41], RC-MVSNet [42]…,leaf,text width=10.5 em] ] [Multi-stage,text width=5 em [U-MVSNet [43], KD-MVS [44]…,leaf,text width=10.5 em] ] ] [Semi-Supervised,text width=6em [ SGT-MVSNet [45],leaf,text width=6.2 em] ] ] [Learning-based Methods
+without Depth Estimation,text width=10em [Voxel-based,text width=10 em [Atlas [46], NeuralRecon [47]…, leaf,text width=13 em] ] [NeRF-based,text width=10 em [VolSDF [48], NeuS [49], Geo-NeuS [50]…, leaf,text width=13 em] ] [3D Gaussian Splatting-based,text width=10 em [ SuGaR [51], 2DGS [52], PGSR [53]…, leaf,text width=13 em] ] [Large feed-forward,text width=10 em [LRM [54], DUSt3R [55], MASt3R [56]…, leaf,text width=13 em] ] ] [Future Research Directions, text width=10 em [Dataset & Benchmarks, View Selection, Depth Fusion, Features, Run-time & Memory Efficiency, Prior Assistance, leaf,text width=37.5 em] ] ]
 Figure 2: Taxonomy of Multi-View Stereo.
-I-A Camera Calibration
+2.1 Camera Calibration
 […]
-I-B View Selection
+2.2 View Selection
 The selection of views is an important step for reconstruction.
@@
 Almost all the following offline MVS methods [36, 31, 39] use the same strategy.
-I-C Multi-view Depth Estimation with Plane Sweep
-Figure 3: Illustration of plane sweep algorithm [33].
-To estimate the depth map for reference image (cam_x), neighboring source images (cam_1, cam_2) are projected with homography to fronto-parallel planes of the reference view frustum.
+2.3 Multi-view Depth Estimation with Plane Sweep
 To form a structured data format that is more suitable for convolution operations, most learning-based MVS methods rely on the plane sweep algorithm [32] to calculate matching costs.
@@
 This practice is deeply inspired by learning-based binocular stereo methods [62, 63, 64], which assess matching costs for a set of disparity hypotheses and subsequently estimate the disparity.
-In a nutshell, the plane sweep algorithm in MVS entails iteratively sweeping planes through the object space, computing homographies between images, and selecting depth values based on consensus among different views, ultimately facilitating accurate 3D reconstruction.
+In a nutshell, the plane sweep algorithm entails iteratively sweeping planes through the object space, computing homographies between images, and selecting depth values based on consensus among different views, ultimately facilitating accurate 3D reconstruction.
 The plane sweep algorithm discretizes the depth space using a series of parallel planes along the depth direction.
 It operates by sweeping a conceptual plane through the object space and evaluating the spatial distribution of geometric surfaces.
-It is worth noting that, for learning-based MVS, plane sweep is conducted purely on GPU, and instead of warping hand-crafted features or edges, the dense pixels are efficiently warped.
-We take a toy example of three cameras viewing the same object as illustrated in Fig.
-3.
-Plane sweep is then performed on the frustum of cam_x (termed as the reference camera) by creating a series of fronto-parallel hypothesized planes within a given range (typically $[d_{\textrm{min}},d_{\textrm{max}}]$), with each plane corresponding to a depth value w.r.t.
-cam_x.
-Let’s then examine two 3D points, $M$ and $M^{\prime}$, as examples of occupied and unoccupied positions on hypothesized planes, respectively.
-For $M$, all of the three cameras capture the identical point lying on the geometry surface with photometric consistency.
-In contrast, the photometric consistency of the observations of $M^{\prime}$ are poor, indicating that $M^{\prime}$ is an invalid hypothesis.
-To evaluate the similarity of the observations towards $M$ and $M^{\prime}$, the images of cam_1 and cam_2 (termed as the source cameras) are warped to each plane by homography.
-These warped images are then compared against the scaled images of cam_x.
-Depth hypotheses with high similarity measures are considered reliable.
 In practice, we divide the depth range, which is either manually set [25] or estimated by SfM [15], into discrete samples and assign hypotheses at these values.
@@
 The warped source feature maps are then obtained via differentiable interpolation.
-Regarding selecting $D$ depth hypotheses from the triangulated depth range $[d_{\textrm{min}},d_{\textrm{max}}]$, there are two main schemes, namely the forward depth sampling and the inverse depth sampling.
-The naive forward sampling divides the depth range into $D-1$ depth intervals with identical lengths.
-Given a depth index $k$, we have
-$d_{k}=d_{\textrm{min}}+\frac{k}{D-1}(d_{\textrm{max}}-d_{\textrm{min}}),k=0,\ldots,D-1,$ (5)
-where the depth hypotheses distribute uniformly between the two ends. While for the inverse sampling scheme [36], we sample uniformly in the multiplicative inverse of $d$, such that
-$\frac{1}{d_{k}}=\frac{1}{d_{\textrm{max}}}+\frac{k}{D-1}(\frac{1}{d_{\textrm{min}}}-\frac{1}{d_{\textrm{max}}}),k=0,\ldots,D-1.$ (6)
-In this way, the distribution of sampled depth values becomes sparser with $d$ approaching to $d_{\textrm{max}}$.
-When reconstructing unbounded outdoor scenes, where the depth range is rather large, the inverse sampling will be a reasonable choice since it samples more densely at the foreground.
-In addition, as revealed in [63], the inverse sampling leads to a uniform sampling on the projected epipolar lines of source images.
-So far, for each depth hypothesis made for cam_x, we have the warped source feature maps as well as the scaled reference feature map.
+To better elaborate the process of plane sweep, we include a toy example in the supplementary.
 The photometric similarity (or matching cost) is measured in a one-to-many manner between reference and warped source features for the following depth estimation, which will be discussed in Sec.
-II-B.
+3.2.
 By performing plane sweep, we obtain a cost volume with a regular spatial shape, making it easy for CNNs to process in parallel.
-I-D Depth Fusion
+2.4 Depth Fusion
 For depth map-based MVS, after estimating all the depth maps, we need to fuse them into a dense 3D representation, e.g., point cloud or mesh.
@@
 There are two main filtering steps: photometric consistency filtering and geometric consistency filtering [15].
-For photometric consistency filtering, a per-pixel confidence, discussed in Sec.
-II-G, is estimated to measure the confidence of depth estimation, i.e., probability that the ground truth depth is within a small range near the estimation.
-Then a threshold can be set to filter depth values with low confidence.
-For geometric consistency filtering, the consistency of depth estimations are measured among multiple views.
-For a pixel $\mathbf{p}$ in the reference view 0, we project it to pixel $\mathbf{p}^{\prime}$ in its $i$-th neighboring view through its depth prediction $d_{0}(\mathbf{p})$.
-After looking up the depth for $\mathbf{p}^{\prime}$, $d_{i}(\mathbf{p}^{\prime})$, we re-project $\mathbf{p}^{\prime}$ back to the reference view at pixel $\mathbf{p}^{\prime\prime}$ and look up its depth $D_{0}(\mathbf{p}^{\prime\prime})$.
-We consider pixel $\mathbf{p}$ and its depth as consistent to the $i$-th neighboring view, if the distances, in image space and depth, between the original estimate and its re-projection satisfy:
-$\xi_{d}=\|\mathbf{p}-\mathbf{p}^{\prime\prime}\|_{2}\leq\tau_{1},$ (7)
-$\xi_{p}=\frac{\|d_{0}(\mathbf{p}^{\prime\prime})-d_{0}(\mathbf{p})\|}{d_{0}(\mathbf{p})}\leq\tau_{2},$ (8)
-where $\tau_{1}$ and $\tau_{2}$ are thresholds.
-The pixels are considered to be reliable estimations if they are consistent in at least $N$ neighboring views.
-Recently, instead of using predefined thresholds for $\tau_{1}$ and $\tau_{2}$, dynamic consistency checking [38] is proposed to dynamically aggregate geometric matching error among all views and improve the robustness of reconstruction.
-Specifically, the dynamic multi-view geometric consistency $C_{\textrm{geo}}(\mathbf{p})$ is computed as:
-$C_{\textrm{geo}}(\mathbf{p})=\sum_{i}\exp{(-(\xi_{p}+\lambda\xi_{d}))},$ (9)
-where $\lambda$ is a weight to balance the reprojection error in two metrics. Then the outliers with $C_{\textrm{geo}}(\mathbf{p})$ smaller than a threshold are filtered.
-I-E Datasets and Benchmarks
-Tab. I is a brief summary of commonly used public MVS datasets and benchmarks for training and evaluation.
-TABLE I: An overview of commonly used datasets and benchmarks for learning-based Multi-View Stereo.
-Dataset Provided Ground Truth1 Synthetic Online Evaluation
-Camera pose Depth Map Point Cloud Mesh Benchmark Target
-ScanNet [26] Depth Map / Mesh
-7-Scenes [27] Depth Map
-DTU [21] Point Cloud
-Tanks and Temples [22] Point Cloud
-ETH3D [23] Point Cloud
-BlendedMVS [24] Depth Map
-  • 1
-    For datasets with online benchmark, the point cloud ground truth of test set is not released.
-ScanNet: ScanNet [26] is a large RGB-D dataset that contains 1613 indoor scenes with ground-truth camera poses, depth maps, surface reconstruction and semantic segmentation labels.
-Online MVS methods [25, 31, 35] mainly use ScanNet for training and testing.
-7-Scenes: 7-Scenes dataset [27] is a RGB-D dataset captured in indoor scenes with a handheld Kinect RGB-D camera.
-Since it is relatively small, 7-Scenes is usually used to test the generalization performance of the models trained on ScanNet [26] without finetuning.
-DTU: DTU dataset [21] is an object-centric MVS dataset collected under well-controlled laboratory conditions with known accurate camera trajectory.
-It contains 128 scans with 49 or 64 views under 7 different lighting conditions.
-Since DTU dataset officially provides scanned ground truth point clouds instead of depth maps, it is required to generate mesh models with surface reconstruction, e.g., screened Poisson surface reconstruction algorithm [67], and then render depth maps [15] for training.
-Tanks and Temples: Tanks and Temples [22] is a large-scale benchmark captured in more complex real indoor and outdoor scenarios.
-It is divided into intermediate and advanced sets.
-Different scenes have different scales, surface reflection and exposure conditions.
-It is usually used to test the generalization performance.
-Note that Tanks and Temples does not provide ground truth camera parameters, which are usually estimated with Structure-from-Motion methods such as COLMAP [17] or OpenMVG [56].
-ETH3D: ETH3D benchmark [23] contains 25 large-scale scenes, including indoor and outdoor, with high-resolution RGB images.
-The scenes typically contain many low-textured regions and non-Lambertian surfaces, e.g., white walls and reflective floor.
-In addition, the images are usually sparse and have strong viewpoint variations and occlusions.
-Therefore, ETH3D is considered as a very challenging benchmark.
-Similar to Tanks and Temples, ETH3D is usually used to test the generalization performance.
-[…]
-I-F Evaluation Metrics
-As mentioned in Tab. I, based on the ground truth, e.g., depth maps or point clouds, evaluation metrics can be categorized into 2D metrics and 3D metrics.
-I-F1 2D Metrics
-2D depth-based metrics are commonly used by online MVS methods [25, 31] to evaluate the depth maps [68, 69].
-We list the commonly used metrics as follows: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel), mean absolute inverse depth error (Abs Inv), and inlier ratio with threshold 1.25 ($\delta<1.25$):
-[…]
-[…]
-[…]
-[…]
-where $d(\mathbf{p})$ denotes the ground truth depth, $\hat{d}(\mathbf{p})$ denotes the estimation, $N$ denotes the number of pixels with valid depth measurements and $\mathbbm{1}[\cdot]$ denotes the indicator function.
-I-F2 3D Metrics
-3D point cloud evaluation is widely used by offline MVS methods [15, 37].
+For more details, please refer to the supplementary.
+2.5 Datasets and Benchmarks
+Along with this survey, we also include a brief summary of commonly used public MVS datasets and benchmarks for training and evaluation. For more details, please refer to supplementary.
+2.6 Evaluation Metrics
+Based on the ground truth, e.g., depth maps or point clouds, evaluation can be categorized into 2D and 3D metrics.
+2D Metrics: 2D depth-based metrics are commonly used by online MVS methods [25, 30] to evaluate the depth maps [68, 69].
+The common metrics are: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel) and inlier ratio with threshold 1.25 ($\delta<1.25$).
+For more details, please refer to supplementary.
+3D Metrics: 3D point cloud evaluation is widely used by offline MVS methods [15, 36].
 Note that the reconstructed point clouds should be aligned to ground truth point clouds before evaluation, e.g., by Iterative Closest Point (ICP), if their extrinsics are differently calibrated.
-Precision/Accuracy: Precision/Accuracy measures the percentage of predicted points that can be matched to the ground truth point cloud.
-Considering a point $\mathbf{P}_{p}$ in the predicted point cloud, it is considered to have a good match in the ground truth point cloud $\{\mathbf{P}_{g}\}$ if
-$\|\mathbf{P}_{p}-\mathop{\arg\min}\limits_{\mathbf{P}\in\{\mathbf{P}_{g}\}}\|\mathbf{P}-\mathbf{P}_{p}\|_{2}\|_{2}\leq\lambda,$ (11)
-where $\lambda$ is a scene-dependent threshold assigned by datasets.
-Usually $\lambda$ is set to a large value for large-scale scenes.
-Note that in some datasets, instead of being measured by percentage [22, 23], precision/accuracy is measured by mean or median absolute distance [21].
-Recall/Completeness: Recall/Completeness measures the percentage of ground truth points that can be matched to the predicted point cloud.
-For a point $\mathbf{P}_{g}$ in the ground truth point cloud, it is considered a good match in the predicted point cloud $\{\mathbf{P}_{p}\}$ if
-$\|\mathbf{P}_{g}-\mathop{\arg\min}\limits_{\mathbf{P}\in\{\mathbf{P}_{p}\}}\|\mathbf{P}-\mathbf{P}_{g}\|_{2}\|_{2}\leq\lambda.$ (12)
-Similar to precision/accuracy, recall/completeness is sometimes measured by the percentage of points [22, 23] and sometimes measured by mean or median absolute distance [21], whose definition is similar to Chamfer distance.
-F-Score: The two aforementioned metrics measure the accuracy and completeness of predicted point clouds.
-However, each of these metrics alone cannot present the overall performance since different MVS methods adopt different assumptions.
-A stronger assumption usually leads to higher accuracy but lower completeness.
-If only precision/accuracy is reported, it would favor MVS algorithms that only include estimated points of high certainty.
-On the other hand, if only recall/completeness is reported it would favor MVS algorithms that include everything, regardless of accuracy.
-Therefore, F-score, an integrated metric, is introduced [22, 23].
-F-score is the harmonic mean of precision and recall.
-It is sensitive to extremely small values and tends to get more affected by smaller values, which means that F-score does not encourage imbalanced results.
-However, in most cases, F-score still suffers from unfairness due to the limitations of ground truth, e.g., sparse and incomplete point clouds may penalize for filling in the areas that are not present in the ground truth [47].
-II Supervised Method with Depth Estimation
+The metrics are Precision/Accuracy, Recall/Completeness and F-score.
+For more details, please refer to supplementary.
+3 Supervised Method with Depth Estimation
 This section mainly introduces supervised learning-based MVS methods with depth estimation.
 A typical depth map-based MVS pipeline mainly consists of the feature extraction (Sec.
-II-A), cost volume construction (Sec.
-II-B), cost volume regularization (Sec.
-II-C) and depth estimation (Sec.
-II-E).
+3.1), cost volume construction (Sec.
+3.2), cost volume regularization (Sec.
+3.3) and depth estimation (Sec.
+3.5).
 The pipelines of MVDepthNet [25] and MVSNet [15] are illustrated in Fig.
 3 and Fig.
-5, which are the representative methods of online and offline MVS methods respectively.
-II-A Feature Extraction
+4, which are the representative methods of online and offline MVS methods respectively.
+3.1 Feature Extraction
 Considering efficiency, most methods use simple CNN structures to extract deep features from images, e.g., ResNet [70], U-Net [71] and FPN [72].
@@
 SimpleRecon [34] proposes utilizing the first two blocks from ResNet18 [70] and an EfficientNet-v2 [74] encoder, which maintains efficiency and yields a sizeable improvement in depth map accuracy.
-For offline MVS methods, MVSNet [15] uses a stacked eight-layer 2D CNN to extract deep features for all the images.
+For offline MVS methods, MVSNet [15] uses a stacked 8-layer 2D CNN to extract deep features for all the images.
 Coarse-to-fine methods further extract multi-scale features for estimation on multiple scales with FPN [31, 38, 39] or multi-scale RGB images [75, 76].
@@
 ET-MVSNet [86] integrates the Epipolar Transformer into the feature extraction, which performs non-local feature aggregation on the epipolar lines.
-II-B Cost Volume Construction
-For both online and offline MVS methods, the cost volume is constructed with the plane sweep algorithm as discussed in Sec. I-C.
-II-B1 Online MVS
+3.2 Cost Volume Construction
+For both online and offline MVS methods, the cost volume is constructed with the plane sweep algorithm as discussed in Sec. 2.3.
+3.2.1 Online MVS
 To reduce computation and improve the efficiency for online applications, online MVS methods usually construct 3D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D}$, which store a single value as matching cost for each pixel $\mathbf{p}$ and depth sample $d_{i}$.
@@
 Then an MLP is used to reduce the dimension of 4D cost volume to a 3D cost volume.
-II-B2 Offline MVS
-Offline MVS methods mainly focus on reconstructing high-quality dense geometry with high-resolution images.
+DoubleTake [89] further includes depth hint, the depth map rendered from the previously predicted geometry, into the cost volume.
+3.2.2 Offline MVS
+Offline MVS methods focus on reconstructing high-quality dense geometry with high-resolution images.
 To encode more matching information and improve the quality, offline methods [15, 31] usually construct 4D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D\times C}$, where each pixel $\mathbf{p}$ and depth sample $d_{i}$ is associated with a matching cost of dimension $C$.
@@
 [94, 39, 40] utilize pixel-wise view weight networks to learn view weights from the pair-wise cost volumes without supervision.
-II-C Cost Volume Regularization
+3.3 Cost Volume Regularization
 Usually, the raw cost volume constructed from image features may be noisy and should be incorporated with smoothness constraint for depth estimation [15].
 Therefore, cost volume regularization is an important step to refine the raw cost volume and aggregate matching information from a large receptive field.
-II-C1 Online MVS
+3.3.1 Online MVS
 A 2D encoder-decoder architecture is commonly used to aggregate information.
@@
 It further uses ConvLSTM [95] to propagate past information with a small overhead of computation time and memory consumption.
-Long et al.
-[94] use MatchNet to regularize the cost volume and ContextNet to learn 2D context information from the reference view.
-By concatenating the cost volume and context feature, a hybrid volume is fed into the Epipolar Spatio-Temporal transformer to aggregate temporal information.
 SimpleRecon [34] fuses multi-scale image features, extracted from the pretrained EfficientNetv2 [74], into the cost volume encoder to improve the performance.
@@
 The final depth map is estimated from the probability volume and refined with the reference image features.
-II-C2 Offline MVS
+3.3.2 Offline MVS
 Among most learning-based offline MVS methods that use 4D cost volumes, there are three main categories to perform cost volume regularization: direct 3D CNN, coarse-to-fine and RNN.
-Fig.
-6 illustrates these three main schemes.
-Figure 6: Illustration of typical cost regularization schemes for offline MVS methods that use 4D cost volumes.
-(a) direct 3D CNN regularization [15, 95, 96, 36] applies a 3D CNN to aggregate contextual information; (b) coarse-to-fine regularization [32, 39, 75, 63] constructs multi-scale cost volumes based on coarse prediction and uses 3D CNN regularization on each scale; (c) RNN regularization [37, 38, 77] sequentially regularizes 2D slices of the cost volume to reduce memory consumption.
 Direct 3D CNN: Similar to stereo estimation [62, 63, 64], 3D CNN is widely used for cost volume regularization in MVS [15, 96, 97, 98, 35].
@@
 WT-MVSNet [81] uses a window-based Epipolar Transformer for enhanced patch-to-patch matching, and a window-based Cost Transformer to better aggregate global information.
-II-D Iterative Update
-Diverging from conventional approaches, certain methods [40, 41, 101, 102] adopt iterative updates to gradually refine depth maps.
-Iterative methods introduce a dynamic approach to depth map estimation, enabling iterative refinement of the reconstruction process.
+3.4 Iterative Update
+Diverging from conventional approaches, certain methods [39, 102, 40, 103, 104] adopt iterative updates to gradually refine depth maps.
 By iteratively updating the depth maps based on successive iterations, these methods can progressively improve the accuracy and consistency of the reconstructed 3D scene.
-This iterative refinement is particularly advantageous in scenarios where initial estimations may be coarse or inaccurate, as it allows for finer adjustments to be made to the depth maps over multiple iterations.
 Moreover, the ability to control the number of iterations provides users with the flexibility to prioritize either computational efficiency or reconstruction quality, depending on specific application requirements.
@@
 After initialization, the approach iterates between propagation and evaluation until convergence.
-Traditional methods usually design fixed patterns for propagation, e.g., propagation of Gipuma [18] as shown in Fig.
-7.
+Traditional methods usually design fixed patterns for propagation.
 Recently, PatchmatchNet [39] proposes learned adaptive propagation and cost aggregation modules, which enables PatchMatch to converge faster and deliver more accurate depth maps.
-Instead of propagating depth samples naively from a static set of neighbors as done traditionally [18], PatchmatchNet adaptively gathers samples from pixels of the same surface with the guidance of image features, shown in Fig.
-8.
-It iteratively updates the depth maps in a coarse-to-fine structure.
-Without 3D CNN, PatchmatchNet demonstrates a remarkably low computation time and memory than other coarse-to-fine methods [32, 39, 75].
-Another PatchMatch-based method, PatchMatch-RL [103], jointly estimates depth, normal and visibility with a coarse-to-fine structure.
-Considering argmax based hard decisions/sampling of PatchMatch is non-differentiable, PatMatch-RL adopts reinforcement learning in the training process.
-Figure 7: Traditional chessboard propagation scheme of Galliani et al.
-[18].
-Planes from a local neighborhood (red points) serve as candidates to update a given pixel (black).
-(a) Standard pattern with 20 neighbors.
-(b) Simplified pattern with 8 neighbors.
-Figure 8: Adaptive propagation scheme of Wang et al.
-[40].
-Pixels located at the object boundary (yellow) and a textureless region (red) receive depth hypotheses from sampled neighbors (green and orange).
-(a) Reference image.
-(b) Fixed sampling locations of classic propagation.
-(c) Adaptive sampling locations with adaptive propagation.
-The grayscale image in (b) and (c) is the ground truth depth map.
-Recently, RAFT [106] estimates optical flow by iteratively updating a motion field through a GRU and achieves state-of-the-art performance.
-The idea is further adopted in stereo [107], scene flow [108] and SfM [109].
+Instead of propagating depth samples naively from a static set of neighbors as done traditionally [18], PatchmatchNet adaptively gathers samples from pixels of the same surface with the guidance of image features.
+PatchMatch-RL [102], jointly estimates depth, normal and visibility with a coarse-to-fine structure.
+Considering argmax based hard decisions/sampling of PatchMatch is non-differentiable, PatMatch-RL adopts reinforcement learning in training.
+Recently, RAFT [107] estimates optical flow by iteratively updating a motion field with GRU.
 In MVS, several methods [40, 104, 103] also adopted this approach to enhance efficiency and flexibility.
 IterMVS [40] proposes a lightweight GRU-based probability estimator that encodes the per-pixel probability distribution of depth in its hidden state.
-In each iteration, multi-scale matching information is injected to update the pixel-wise depth distribution.
-CER-MVS [101] and Effi-MVS [102] both embed the RAFT module in a coarse-to-fine structure.
-With the injection of matching information, the hidden state is updated and outputs a residual that is added to the previous depth estimation.
+CER-MVS [103] and Effi-MVS [104] both embed the RAFT module in a coarse-to-fine structure, which outputs a residual that is added to the previous depth.
+MaGNet [88] iteratively updates the Gaussian distribution of depth for each pixel with the matching scores.
 DELS-MVS [108] proposes Epipolar Residual Network to search for the corresponding point in the source image directly along the corresponding epipolar line and follows an iterative manner to narrow down the search space.
 Based on IterMVS [40], RIAV-MVS [109] iteratively refines index fields with a 3-level GRUs.
-MaGNet [88] iteratively updates the Gaussian distribution of depth for each pixel with the matching scores.
-IGEV-MVS [112] iteratively updates a disparity map regressed from geometry encoding cost volumes and pairs of group-wise correlation volumes.
-II-E Depth Estimation
-II-E1 Online MVS
+IGEV-MVS [110] iteratively updates a disparity map regressed from geometry encoding cost volumes and pairs of correlation volumes.
+3.5 Depth Estimation
+3.5.1 Online MVS
 Many methods [25, 60, 30, 34] apply encoder-decoder on the cost volume $\mathbf{C}$ and reduce the feature channel to 1, i.e., $\mathbf{C}^{\prime}\in\mathbb{R}^{H\times W\times 1}$.
@@
 For other methods, Neural-RGBD [87] directly models the cost volume as probability volume $\mathbf{P}$ and adopts soft argmax [62] to predict depth, Eq.
-16.
+8.
 MaGNet [88] takes the mean of Gaussian distribution at last iteration as depth.
-II-E2 Offline MVS
+3.5.2 Offline MVS
 For a cost volume $\mathbf{C}\in\mathbb{R}^{H\times W\times D\times C}$, a probability volume $\mathbf{P}\in\mathbb{R}^{H\times W\times D}$ is usually generated after cost volume regularization, which is then used for depth estimation.
@@
 On the other hand, some methods [35, 39] compute the expectation value of inverse depth samples since this sampling is more suitable for complex and large-scale scenes [35].
-The depth prediction is computed as:
-$\hat{d}(\mathbf{p})=\left(\sum_{i=1}^{D}\frac{1}{d_{i}}\cdot\mathbf{P}(i,\mathbf{p})\right)^{-1}.$ (17)
 In contrast, methods [36, 37, 77] that use RNN for cost volume regularization mainly adopt argmax operation.
@@
 Since the argmax operation adopted by winner-take-all cannot produce depth estimations with sub-pixel accuracy, the depth map may be refined in post-processing [36].
-Comparatively, soft argmax operation corresponds to measuring the distance of the expectation to the ground truth depth.
-While the expectation can take any continuous value, the measure cannot handle multiple modes in $\mathbf{P}$ and strictly prefers unimodal distributions [41].
-The argmax operation corresponds to measuring the Kullback-Leibler divergence between a one-hot encoding of the ground truth and $\mathbf{P}$, but cannot achieve sub-pixel precision [41].
 Recently, Wang et al.
-[41] propose a simple yet effective hybrid strategy that combines regression and classification.
-Similar to [113], it is robust to multi-modal distributions but also achieves sub-pixel precision.
-First, similar to R-MVSNet [37], the index $\mathbf{X}(\mathbf{p})$ with the highest probability for pixel $\mathbf{p}$ from probability $\mathbf{P}$ is found.
-Then the expectation in the local inverse range is computed as the depth estimate $\hat{d}(\mathbf{p})$:
-$\hat{d}(\mathbf{p})\!\!=\!\!\left(\!\!\frac{1}{\sum_{j=\mathbf{X}(\mathbf{p})-r}^{\mathbf{X}(\mathbf{p})+r}\mathbf{P}(j,\mathbf{p})}\!\sum_{j=\mathbf{X}(\mathbf{p})-r}^{\mathbf{X}(\mathbf{p})+r}\!\!\frac{1}{d_{j}}\!\cdot\!\mathbf{P}(j,\mathbf{p})\!\!\right)^{\!\!\!-1}\!\!\!\!\!,$ (18)
-where $d_{j}$ is the $j$-th depth sample.
-Concurrently, Peng et al.
-[114] propose a different strategy to unify regression and classification.
-They first use classification to get the optimal hypothesis and then regress the proximity for it.
+[40] propose a hybrid strategy to combine regression and classification.
+Similarly, Peng et al.
+[111] first use classification to get the optimal hypothesis and then regress the proximity for it.
 DMVSNet [112] estimates two depth maps on multi-stages and composes the final depth map by alternating between selecting the maximum and minimum predicted depth values, which generates an oscillating depth map for improved geometry.
-II-F Depth Refinement
+3.6 Depth Refinement
 Given that the raw depth estimation from MVS may be noisy, refinement is usually used to improve the accuracy.
 R-MVSNet [36] enforces multi-view photo-consistency to alleviate the stair effect and achieve sub-pixel precision.
-Point-MVSNet [95] and VA-Point-MVSNet [116] use PointFlow to refine the point cloud iteratively by estimating the residual between the depth of the current iteration and that of the ground truth.
+Point-MVSNet [97] uses PointFlow to refine the point cloud iteratively by estimating the residual between the depth of the current iteration and that of the ground truth.
 Fast-MVSNet [113] adopts an efficient Gauss-Newton layer to optimize the depth map by minimizing the feature residuals.
@@
 EPNet [117] uses a hierarchical edge-preserving residual learning module to refine multi-scale depth estimation with image context features.
-II-G Confidence Estimation
+3.7 Confidence Estimation
 As discussed in Sec.
-I-D, photometric confidence is important to filter out unreliable estimations during depth fusion for offline MVS methods.
+2.4, photometric confidence is important to filter out unreliable estimations during depth fusion.
 Following MVSNet [15], most offline MVS methods take the probability of the estimation [36, 37, 77] or the probability sum over several samples near the estimation [15, 31, 39] from the probability volume as confidence.
-Fig.
-9 provides a set of RGB images, estimated depth maps, and the visualized confidence.
 In stereo matching, some methods learn confidence from disparity [118], RGB image [119, 120] or matching costs [121] and obtain confidence scores in $[0,1]$ interval.
 Motivated by this, some traditional MVS methods [122, 123] based on classical PatchMatch [106] propose to estimate the confidence with deep learning and use this to refine the results from PatchMatch.
-Recently, IterMVS [41] estimates confidence from the pixel-wise depth probability distributions, which are encoded by the hidden state of a convolutional GRU.
+Recently, IterMVS [40] estimates confidence from the hidden state of a convolutional GRU.
 A 2D CNN followed by a sigmoid is applied to the hidden state to predict the confidence.
-DELS-MVS [110] feeds pixel-wise entropy of the partition probabilities, which is computed for each source image, into a confidence network to learn a confidence map.
-The confidence map is used to guide the fusion of multiple depth maps.
-Figure 9: Visualization of confidence on DTU [21]. The regions with low confidence typically correspond to occlusions and low texture.
-II-H Loss Function
-II-H1 Online MVS
+DELS-MVS [108] feeds pixel-wise entropy of the partition probabilities, which is computed for each source image, into a confidence network to learn confidence.
+The confidence is used to guide the fusion of multiple depth maps.
+3.8 Loss Function
+3.8.1 Online MVS
 Many methods [25, 60, 30] compute the regression loss of estimated inverse depth maps for training:
@@
 To improve performance, SimpleRecon further uses gradient loss on depth, normal loss where normal is computed with depth and intrinsics, and multi-view regression loss.
-II-H2 Offline MVS
+3.8.2 Offline MVS
 Based on the depth estimation strategy as discussed in Sec.
-II-E, loss functions can be mainly categorized into regression and classification.
+3.5, loss functions can be mainly categorized into regression and classification.
 For methods [15, 35] that predict depth with soft argmax [62], (smooth) $L_{1}$ loss is usually adopted as the loss function, which is stated as:
@@
 [111] use focal loss [124].
-III Unsupervised & Semi-supervised Methods with Depth Estimation
-Figure 10: Ambiguous supervision from photometric consistency assumption [43].
-(a) Complicated lighting condition caused by varying camera exposures.
-(b) Invisibility of specific area caused by occlusion.
+4 Unsupervised & Semi-supervised Methods with Depth Estimation
 In this section, we introduce unsupervised learning-based MVS methods [125, 126, 42, 43, 41] and semi-supervised method [45].
 Supervised MVS methods mentioned in Sec.
-II depend extensively on the availability of accurate ground truth depth maps obtained through depth-sensing equipment.
-This requirement not only complicates the data collection process, making it labor-intensive and costly but also restricts these methods to limited datasets and primarily indoor settings.
+3 depend extensively on the availability of accurate ground truth depth maps obtained through depth-sensing equipment.
 To make MVS practical in more general real-world scenarios, it is vital to consider alternative unsupervised learning-based methods that can provide competitive accuracy compared to the supervised ones without any ground truth depth.
 Existing unsupervised methods are built upon the assumption of photometric consistency (Sec.
-III-A), which indicates that corresponding pixels from different views of the same 3D point share similar features.
-These approaches are categorized into end-to-end ones (Sec.
-III-B) and multi-stage ones (Sec.
-III-C).
+4.1), and are categorized into end-to-end (Sec.
+4.2) and multi-stage (Sec.
+4.3).
 SGT-MVSNet [45] is the only semi-supervised method so far, and we introduce it in Sec.
-III-D.
-We also analyze the primary challenges in unsupervised MVS in Sec.
-III-B.
-III-A Photometric Consistency Assumption
+4.4.
+4.1 Photometric Consistency Assumption
 In the realm of unsupervised depth map prediction, extant methods [127, 43, 128, 129] endeavor to establish photometric consistency between reference and source perspectives.
@@
 The computation of the structural similarity loss occurs between a synthesized image and the initial reference image, aiming to uphold contextual congruity.
-More precisely, the assessment of contextual similarity often involves the utilization of the Structural Similarity Index (SSIM) proposed by [134], a metric commonly employed for quantifying contextual consistency.
-This metric is defined as follows:
+More precisely, the assessment of contextual similarity often involves the Structural Similarity Index (SSIM) proposed [130], defined as:
 $L_{\textrm{SSIM}}=\sum_{i=1}^{N-1}\left[1-\operatorname{SSIM}\left(\mathbf{I}_{0},\hat{\mathbf{I}}_{0}^{i}\right)\right]\odot M_{i},$ (13)
@@
 $L_{\textrm{SM}}=\sum_{\mathbf{x}}\left|\nabla_{u}\tilde{d}(\mathbf{x})\right|e^{-\left|\nabla_{u}\mathbf{I}_{0}(\mathbf{x})\right|}+\left|\nabla_{v}\tilde{d}(\mathbf{x})\right|e^{-\left|\nabla_{v}\mathbf{I}_{0}(\mathbf{x})\right|},$ (15)
-where $\nabla_{u}$ and $\nabla_{v}$ refer to the gradient along x and y axis, $\tilde{d}=d/\bar{d}$ is the mean-normalized inverse depth, $M$ represents the set of valid pixels in the reference image.
-III-B End-to-end Unsupervised Methods
-End-to-end methods [131, 42, 43, 135, 136] are those methods that are trained from scratch with the same input information as supervised methods (Sec.
-II), while not using ground truth depth for supervision.
-Instead, they usually adopt photometric consistency, structural similarity and smoothness constraint as parts of their loss terms.
-Unsup_MVS [131] proposes to inherit the supervision signal of view synthesis, as discussed in Sec.
-III-A, and dynamically select $X$ “best” (lowest loss) values out of $Y$ loss maps.
-However, the bottleneck in unsupervised MVS is to find such accurate photometric correspondences.
-As illustrated in Fig.
-10, in the real scenarios, non-Lambertian surfaces, varying camera exposures, and occlusions will make the assumption of photometric consistency invalid.
-Hence, such an invalid assumption will lead to “ambiguous supervision”.
-To alleviate the ambiguous supervision, JDACS [42] introduces semantic consistency in addition to photometric consistency.
-It proposes to extract semantic features with a pre-trained network and create semantic classification map by applying non-negative matrix factorization.
-Then pixel-wise cross-view segmentation consistency is calculated.
-The difference here is that the semantic map is warped, instead of the image.
-In this way, the semantic classification maps are supervised by cross-entropy loss between the warped map $S_{i}$ and reference map $S_{0}$, both encoded as pixel-wise one-hot vector.
-However, the proposed semantic cross-view consistency loss is unstable to converge during training and such high-level semantic information excludes details and boundaries in MVS.
-RC-MVSNet [43] digs into providing reliable and occlusion-aware supervision by introducing neural rendering.
-It takes the advantage of both strong representation ability of Neural Radiance Field [28] and strong generalization ability of cost volume.
-First, the unsupervised CasMVSNet [32] backbone in RC-MVSNet yields initial depth map supervised by photometric consistency.
-In addition, it generates depth priors for the Gaussian-Uniform mixture sampling in the rendering-consistency (RC) network.
-Then after volumetric rendering of the reference view, the depth map is supervised by an extra reference view synthesis loss and a depth rendering consistency loss.
-In this way, occlusion-aware neural rendering with mixture sampling is utilized to alleviate the invisibility phenomenon and ambiguous supervision from different lighting condition is solved by reference view synthesis.
-ElasticMVS [135] introduces an architecture of part-aware patchmatch to address limitations in photometric loss-based geometry information, which tends to have missing data and artifacts in certain areas, particularly textureless regions.
-The proposed framework incorporates an elastic part representation to encode geometric details for guiding piecewise-smooth depth map prediction.
-The initial step involves the formal definition of part-aware representation.
-Building upon this learned representation, enhancements are proposed for both depth hypotheses propagation and evaluation within the context of part-aware patchmatch.
-Finally, the network is optimized through the incorporation of a dense contrastive loss for self-supervised training, coupled with the integration of a spatial concentration loss [137] designed to promote the isotropic isolation of all pixel embeddings.
-CL-MVSNet [136] proposes a framework that aims to enhance proximity among positive pairs by ensuring contrastive consistency between a regular branch and two contrastive branches.
-The regular branch is a CasMVSNet [32] structure that has been used in  [43, 135, 42] as well, while contrastive branches incorporate an image-level branch and a scene-level branch.
-These branches enforce contrastive consistency guided by the confidence mask estimated from the regular branch.
-Additionally, the $L0.5$ photometric consistency is proposed to enhance the accuracy of the reconstruction.
-Note that these end-to-end methods are all trained from scratch without any pre-processing which saves training time and reduces the complexity of application in real scenarios.
-III-C Multi-stage Unsupervised Methods
-Multi-stage methods require either pre-training of a specific module or pre-processing on the training data. These methods are built upon the idea of pseudo-label generation.
-Self-supervised CVP-MVSNet [132] proposes to generate pseudo ground truth of depth by training the CVP-MVSNet [75] backbone with photometric consistency loss (Sec.
-III-A).
-The pseudo ground truth depth maps are refined with high-resolution images, filtered with cross-view depth consistency check and then fused in a point cloud [15].
-Then the mesh is reconstructed with the screened Poisson surface reconstruction algorithm [67] to render complete pseudo depth maps for training in the next iteration.
-The whole pipeline consists of several iterations to improve the performance.
-However, [132] cannot provide a satisfying reconstruction of the 3D objects since it heavily relies on the photometric consistency assumption.
-Additionally, the self-training stage of the whole pipeline takes few days to complete.
-U-MVSNet [44] introduces a flow-depth consistency loss by pre-training an optical-flow estimation network, PWC-Net [138], in an unsupervised manner.
-The dense 2D optical flow correspondences are used to generate pseudo labels for cross-view flow-depth consistency, which alleviates the ambiguous supervision in the foreground.
-Then the uncertainty-aware self-training consistency further reduces the invalid supervision in the background utilizing the generated pseudo labels and uncertainty map.
-Recently, KD-MVS [45] achieves superior performance by leveraging a typical teacher-student scheme of knowledge distillation [139, 140, 141].
-It trains a teacher MVS model, e.g., MVSNet [15] or CasMVSNet [32], in a self-supervised manner, by enforcing both photometric and feature-metric consistency between the reconstructed and original images.
-Then the teacher model produces pseudo ground-truth labels of the training set by cross-view consistency check and probabilistic encoding.
-In addition to the reliable estimation by the teacher model, the pseudo labels also contain probabilistic knowledge with estimation uncertainty.
-By training MVS networks [15, 32] with the pseudo labels, the knowledge of the teacher model is distilled to the student models.
-III-D Semi-supervised Methods
-SGT-MVSNet [46] proposes to use only a few sparse ground truth 3D points to estimate the depth map of the reference view.
-3D point consistency loss is used to minimizes the difference between the 3D points back-projected from the corresponding pixels and the ground truth.
-To tackle the problem of inaccurate estimation on the edge and boundary, a coarse-to-fine reliable depth propagation module rectifies the erroneous predictions.
-IV Learning-based MVS without Depth Estimation
+where $\nabla_{u}$ and $\nabla_{v}$ refer to the gradient along x and y axis, $\tilde{d}=d/\bar{d}$ is the mean-normalized inverse depth, and $M$ represents the set of valid pixels in the reference image.
+4.2 End-to-end Unsupervised Methods
+End-to-end methods [127, 41, 42, 131, 132] train from scratch with the same input as supervised methods (Sec.
+3) but without ground truth depth.
+They are based on photometric consistency, structural similarity, and smoothness constraints for loss terms.
+Unsup_MVS [127] adopts view synthesis supervision and dynamically selects the $X$ best (lowest loss) values out of $Y$ loss maps.
+However, photometric correspondences can be inaccurate due to non-Lambertian surfaces, camera exposure variations, and occlusions, leading to ambiguous supervision (Sec.
+4.1).
+JDACS [41] introduces semantic consistency to address these challenges, using a pre-trained network to generate semantic maps and enforcing cross-view segmentation consistency.
+However, this approach can struggle to converge and lacks detailed information.
+RC-MVSNet [42] enhances supervision through neural rendering by combining Neural Radiance Fields and cost volumes.
+It uses a CasMVSNet backbone to generate depth maps supervised by photometric consistency, along with rendering-consistency and view synthesis losses to handle occlusions and varying lighting.
+ElasticMVS [131] addresses issues in photometric loss-based geometry by introducing a part-aware patch-match framework, using an elastic part representation to guide the depth map prediction.
+The network is optimized with contrastive and spatial concentration losses to promote pixel isolation.
+CL-MVSNet [132] improves proximity between positive pairs through contrastive consistency between a regular CasMVSNet branch and two contrastive branches.
+It also introduces $L0.5$ photometric consistency to improve the precision of the reconstruction.
+These end-to-end methods train from scratch without pre-processing, reducing training time and complexity in real-world applications.
+4.3 Multi-stage Unsupervised Methods
+Multi-stage methods require pre-training or data pre-processing and are based on pseudo-label generation.
+Self-supervised CVP-MVSNet [128] generates pseudo ground truth depth using photometric consistency, refining depth maps with cross-view consistency checks and point cloud fusion [15].
+The meshes are then reconstructed for further training iterations.
+U-MVS [43] pre-trains an optical-flow network, PWC-Net [133], and uses flow-depth consistency to generate pseudo labels, reducing supervision ambiguity in foregrounds and backgrounds with an uncertainty-aware self-training consistency.
+KD-MVS [44] employs knowledge distillation to increase performance.
+A teacher MVS model generates pseudo ground-truth labels through photometric and feature-metric consistency loss, with uncertainty encoding.
+These labels are then used to train the student models.
+4.4 Semi-supervised Methods
+SGT-MVSNet [45] uses sparse 3D points to estimate depth maps.
+A 3D point consistency loss minimizes differences between back-projected 3D points and ground truth.
+A coarse-to-fine depth propagation module improves accuracy at edges and boundaries.
+5 Learning-based MVS without Depth Estimation
 Though the learning-based methods that predict individual depth maps with plane-sweep are the main family of learning-based multi-view stereo, there are many methods of other families that achieve impressive 3D reconstruction quality in the recent years.
 In this section, We discuss four main families: voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
-IV-A Voxel-based Methods
+5.1 Voxel-based Methods
 These methods [46, 47, 134, 135] estimate the scene geometry with volumetric representation by leveraging implicit function, e.g., SDF.
 Specifically, Atlas [46] and NeuralRecon [47] attempt to predict the TSDF volume from the 3D feature volume constructed by lifting 2D image features.
-As shown in Fig.
-11, Atlas uses 3D CNN to regress the TSDF volume based on the feature volume accumulated from all images of the scene, which exhibits great completeness of reconstruction.
+Atlas uses 3D CNN to regress the TSDF volume based on the feature volume accumulated from all images of the scene, which exhibits great completeness of reconstruction.
 The TSDF reconstruction is supervised using $L_{1}$ loss to the ground truth TSDF vaules.
@@
 VoRTX [135] uses a similar design to TransformerFusion and the scene geometry is obtained by passing three different-level features output from transformers through a 3D CNN.
-Figure 11: Pipeline of Atlas [47]. 2D image features are back-projected into 3D volumes, which are aggregated and passed through a 3D CNN to directly regress a TSDF volume.
-IV-B NeRF-based Methods
+5.2 NeRF-based Methods
 In novel view synthesis, Neural Radiance Field (NeRF) [27] has kicked off a new emerging representation of 3D, which offers a differentiable volume-rendering scheme to supervise a 3D radiance-based representation with 2D image-level losses.
-The pipeline of NeRF is shown in Fig.
-12.
 NeRF employs multi-layer perceptron (MLP) to map a position $(x,y,z)$ and the normalized view direction $(\theta,\phi)$ to the corresponding color $\mathbf{c}$ and volume density $\sigma$.
@@
 Many subsequent endeavors [136, 137, 138, 139, 140, 141, 142, 143, 144] further improve NeRF in quality, fast training, memory efficiency and real-time rendering.
-Figure 12: Pipeline of NeRF [28].
-Given a 3D position and 2D viewing direction (a), an MLP produces the color and volume density (b).
-Then volume rendering is used to composite these values into an image (c).
-The optimization is minimizing the rendering loss (d).
 Though the initial purpose of NeRF is to perform novel view synthesis, VolSDF [48] and NeuS [49] integrate NeRF with SDF for surface reconstruction.
@@
 Note that without ground truth geometry supervision, e.g., depth map or TSDF volume, these methods are trained in a self-supervised way as NeRF, Eq.
-27.
+17.
 Since training with rendering loss only like NeRF has ambiguity in geometry [145], some following methods improve the reconstruction with explicit geometry supervision, e.g., monocular depth/normal priors [146, 147] and sparse SfM point cloud [50].
@@
 Specifically, based on VolSDF/NeuS, these methods mainly replace the view direction with the reflected view direction following Ref-NeRF.
-IV-C 3D Gaussian Splatting-based Methods
+5.3 3D Gaussian Splatting-based Methods
 Unlike implicit representations with a coordinate-based MLP such as NeRF [27], 3D Gaussian Splatting [160] (3DGS) explicity represents the scene with point primitives, each of which is parameterized as a scaled Gaussian with 3D covariance matrix $\mathbf{\Sigma}$ and mean $\mathbf{\mu}$:
@@
 During training, 3DGS minimizes the rendering loss like NeRF, as in Eq.
-27.
+17.
 3DGS can be initialized with SfM [161, 17] or MVS [33], which performs better than random initialization [160].
@@
 Though 3DGS achieves high-quality novel-view synthesis, it is challenging to recover high quality geometry since no explicit geometry constraint is used and 3D Gaussians do not correspond well to the actual surface because of the 3D covariance [51].
-As one of the earliest methods, DreamGaussian [51] follows NeRF [28] to compute a dense density grid with the 3D Gaussians and then extracts mesh with Marching Cubes [66].
 SuGaR [51] introduces a geometry regularization term encouraging the 3D Gaussians to be well-aligned over the scene surfaces so that the Gaussians can contribute to better scene geometry.
@@
 During training, the normals from the SDF field of Neuralangelo regularize the rotation of 3D Gaussians, while the SDF field is regularized to ensure that the SDF values at 3D Gaussians’ positions are close to zero.
+2DGS [52] further improves surface accuracy by modeling the 3D primitives as surfels ( tangent plane clipped) and fuses rendered depth maps into mesh with TSDF fusion.
+PGSR [53] elicits unbiased depth maps as strong geometry prior from rendered distances and rendered normals, and enforces multi-view geometric consistency by regularizing such a prior with local plane assumption and edge constraint.
+RaDe-GS [164] also improves the quality of geometry by precisely modeling the depth and normal, instead of the error-prone approximate projection in the original 3DGS.
 However, retrieving good geometry from Gaussian-based representations still needs to be explored.
-IV-D Large Feed-forward Methods
-In 3D reconstruction and 3D generation, a recent prominent trend is to directly learn the 3D representation from large-scale 3D datasets, such as Objaverse [30].
-These methods typically adopt large-scale transformers [174] to generate a 3D representation based on various inputs, including single image [53, 175, 176], textual description [177], posed multi-view images [178], and un-posed multi-view images [179, 54].
+5.4 Large Feed-forward Methods
+In 3D reconstruction and 3D generation, a recent prominent trend is to directly learn the 3D representation from large-scale 3D datasets, e.g.Objaverse [29].
+These methods typically adopt large-scale transformers [165] to generate a 3D representation based on various inputs, including single image [54, 166, 167], textual description [168, 169], posed multi-view images [168, 170, 171, 172, 173], and un-posed multi-view images [174, 55].
 LRM [54] uses a transformer-based feed-forward network to regress the 3D object from a single image.
-Specifically, LRM predicts features of a tri-plane representation [180], which is subsequently converted into an implicit field like NeRF [53].
-Then the surface is extracted with Marching Cubes [66].
-Recently, DUSt3R [54] reconstructs the scene as a point cloud given a pair of un-calibrated and un-posed images.
+Specifically, it predicts features of a tri-plane representation [175], which is subsequently converted into a radiance field similar to NeRF.
+Large feed-forward models can also be combined with 3D Gaussian Splatting, where they directly generate 3D Gaussian points from posed multi-view inputs [168, 170, 171, 172, 173, 176].
+Instead of predicting 3D Gaussian locations directly, which is a highly ill-posed task, the common practice is to use per-pixel aligned Plücker ray embeddings and predict the depth for each pixel [176] of input views.
+As such, this design requires the input views to cover a large range of the object.
+Recently, DUSt3R [55] predicts pixel-aligned point coordinates directly from uncalibrated image pairs.
 After extracting image token representations with Vision Transformer [177], two transformer decoders exchange information of the token representations via cross-attention and then regress the per-pixel 3D points in the coordinate frame of the first image.
 Many downstream tasks can be performed with the point clouds, e.g., point matching, localization, intrinsic and relative pose estimation.
-V Discussions
+Directly extended from DUSt3R, MASt3R [56] focuses on the specific application of feature description and matching across multi-view images.
+It leverages an additional matching head on top of DUSt3R’s architecture to boost the performance of feature-based matching.
+With the dense correspondences established, MASt3R manages to achieve a comparable reconstruction quality as MVSNet [15].
+Spann3R [178] manages an external spatial memory to keep track of all previous relevant 3D information yielded by DUSt3R and then queries this spatial memory to predict the 3D structure.
+6 Discussions
 In this section, we summarize and discuss the performance of learning-based MVS methods, including supervised online methods with depth estimation (Sec.
-V-A), supervised offline methods with depth estimation (Sec.
-V-B), unsupervised methods with depth estimation (Sec.
-V-C) and methods without depth estimation (Sec.
-V-D).
+6.1), supervised offline methods with depth estimation (Sec.
+6.2), unsupervised methods with depth estimation (Sec.
+6.3) and methods without depth estimation (Sec.
+6.4).
 Moreover, we discuss potential directions for future research (Sec.
-V-E).
-TABLE II: Quantitative results of supervised online MVS methods on ScanNet [26] and 7-Scenes [27].
+6.5).
+TABLE I: Quantitative results of supervised online MVS methods on ScanNet [26] and 7-Scenes [179].
 Methods ScanNet [26] 7-Scenes [179]
-abs $\downarrow$ abs-rel $\downarrow$ abs-inv $\downarrow$ $\eta<1.25$uparrow abs $\downarrow$ abs-rel $\downarrow$ abs-inv $\downarrow$ $\eta<1.25$uparrow
-MVDepthNet [25] 0.167 0.087 0.054 0.925 0.201 0.117 0.071 0.877
-DPSNet [97] 0.219 0.119 0.071 0.868 0.249 0.149 0.085 0.826
-Neural-RGBD [87] 0.236 0.122 0.075 0.850 0.214 0.131 0.076 0.865
-GP-MVS [58] 0.149 0.076 0.049 0.940 0.174 0.100 0.064 0.903
-Long et al. [94] 0.151 0.081 - 0.931 0.253 0.147 - 0.804
-DeepVideoMVS [31] 0.119 0.060 0.038 0.965 0.145 0.038 0.054 0.938
-MaGNet [88] 0.147 0.081 - 0.930 0.213 0.126 - 0.855
-RIAV-MVS [111] 0.139 0.075 - 0.938 0.178 0.100 - 0.897
-SimpleRecon [35] 0.089 0.043 - 0.981 0.105 0.058 - 0.974
-TABLE III: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for unsupervised depth map-based MVS methods.“Finetuned with BlendedMVS” denotes whether the methods are finetuned on BlendedMVS [24] before evaluating on Tanks and Temples [22].
+abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$uparrow abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$uparrow
+MVDepthNet [25] 0.167 0.087 0.925 0.201 0.117 0.877
+DPSNet [96] 0.219 0.119 0.868 0.249 0.149 0.826
+Neural-RGBD [87] 0.236 0.122 0.850 0.214 0.131 0.865
+GP-MVS [60] 0.149 0.076 0.940 0.174 0.100 0.903
+DeepVideoMVS [30] 0.119 0.060 0.965 0.145 0.038 0.938
+MaGNet [88] 0.147 0.081 0.930 0.213 0.126 0.855
+RIAV-MVS [109] 0.139 0.075 0.938 0.178 0.100 0.897
+SimpleRecon [34] 0.089 0.043 0.981 0.105 0.058 0.974
+DoubleTake [89] 0.077 0.037 0.984 0.099 0.053 0.970
+TABLE II: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for unsupervised depth map-based MVS methods.“Finetuned with BlendedMVS” denotes whether the methods are finetuned on BlendedMVS [24] before evaluating on Tanks and Temples [22].
 Methods DTU [21] Finetuned with Tanks and Temples [22]
@@
 Multi-stage Self-supervised CVP-MVSNet [128] 0.308 0.418 0.363 43.48 -
-U-MVSNet [44] 0.354 0.354 0.354 57.15 30.97
+U-MVS [43] 0.354 0.354 0.354 57.15 30.97
 KD-MVS [44] 0.359 0.295 0.327 64.14 37.96
-TABLE IV: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for supervised offline MVS methods.
-“Finetuned with BlendedMVS” denotes whether the learning-based methods are finetuned on BlendedMVS [24] before evaluating on Tanks and Temples [22] and ETH3D [23].
+TABLE III: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for supervised offline MVS methods.
+“Finetuned with BlendedMVS” denotes whether methods are finetuned on BlendedMVS dataset [24] before evaluation on Tanks and Temples [22] and ETH3D [23].
 Methods DTU [21] Finetuned with Tanks and Temples [22] ETH3D [23]
@@
 PointMVSNet [97] 0.342 0.411 0.376 - - - - -
-VA-Point-MVSNet [116] 0.359 0.358 0.359 48.70 - - -
 PVA-MVSNet [91] 0.379 0.336 0.357 54.46 - - -
@@
 ET-MVSNet [86] 0.329 0.253 0.291 65.49 40.41 - -
+GoMVS [187] 0.347 0.227 0.287 66.64 43.07 79.16 85.91
 Iterative update PatchmatchNet [39] 0.427 0.277 0.352 53.15 32.31 64.21 73.12
@@
 IGEV-MVS [110] 0.331 0.326 0.324 - - - - -
-V-A Supervised Online MVS with Depth Estimation
+6.1 Supervised Online MVS with Depth Estimation
 Typically, online MVS methods are trained on ScanNet [26] and then evaluated on ScanNet [26] and 7-Scenes [179].
 We summarize the quantitative results of online MVS methods in Tab.
-II.
+I.
 By incorporating temporal information in depth estimation,  [60, 188, 30] achieve better performance than MVDepthNet [25].
@@
 Instead of temporal fusion, SimpleRecon [34] injects cheaply available metadata into the cost volume and achieves state-of-the-art performance on both ScanNet and 7-Scenes.
-V-B Supervised Offline MVS with Depth Estimation
+6.2 Supervised Offline MVS with Depth Estimation
 Offline MVS methods are usually trained on DTU [21] and then evaluated on DTU [21], Tanks and Temples [22] and ETH3D [23].
 Recently, many methods further finetune the DTU-pretrained model on BlendedMVS [24] before evaluating on Tanks and Temples and ETH3D since it contains large-scale scenes.
-V-B1 Benchmark Performance
+6.2.1 Benchmark Performance
 Tab.
-IV summarizes quantitative results of offline MVS methods.
+III summarizes quantitative results of offline MVS methods.
 Compared with those methods that use direct 3D CNN for regularization, the methods that use RNN and coarse-to-fine regularization perform much better.
@@
 Therefore, it is valuable to pay more attention to ETH3D to evaluate the robustness of learning-based methods in real-world scenes.
-V-B2 Memory Consumption and Run-time
-Low run-time and memory consumption are essential in most industrial applications with limited computational power and storage, e.g., autonomous driving and AR/VR.
+6.2.2 Memory Consumption and Run-time
+Low run-time and memory consumption are crucial in most industrial applications with limited computational power and storage, e.g., robotics and AR/VR.
 Because of the low-resolution input and the simplicity of network structures, online MVS methods [25, 30, 34] usually achieve high efficiency in both memory and run-time.
 In contrast, offline MVS methods focus on high-resolution images and usually use computationally expensive network modules, e.g., 3D convolutions, to improve reconstruction quality.
-These increase run-time and GPU memory.
-Recently, many researchers try to improve the efficiency while maintaining the reconstruction accuracy as other offline MVS methods.
-We compare and summarize the efficiency of existing efficient methods [117, 32, 39, 40, 102, 41, 99] in Fig.
-13.
-All the methods share similar coarse-to-fine structures to improve efficiency.
-To further reduce run-time and memory, [40, 102, 41, 99] further get rid of costly 3D convolution and use 2D convolution instead.
-It can be found many methods [40, 102, 41, 99] can achieve near real-time estimation (about 10 Hz) with relatively high resolution images, e.g., $1152\times 864$.
-In conclusion, it is promising to further improve both the performance and efficiency for practical applications.
-Figure 13: GPU memory consumption and run-time of efficient offline methods [117, 32, 39, 40, 102, 41, 99] on DTU dataset [21].
-The image resolution and the number of images are set to $1152\times 864$ and 5 respectively.
-Experiments are done on a NVIDIA 2080Ti GPU.
-V-C Unsupervised MVS with Depth Estimation
+These explicitly increase run-time and GPU memory.
+Recently, many researchers tried to improve the efficiency while maintaining the reconstruction accuracy as other offline MVS methods.
+We compare the efficiency of state-of-the-art offline MVS methods [31, 39, 104, 40, 100, 79, 111, 76, 86] in Fig.
+5.
+All these methods share similar coarse-to-fine structures to improve efficiency.
+However, we observe that some methods [31, 79, 111, 76] still require lots of computation resource and are not efficient in either memory or run-time.
+To further reduce run-time and memory, [39, 104, 40, 100] carefully simplify the architecture, e.g., replacing costly 3D convolution with 2D convolution.
+Moreover, these methods still achieve competitive performance when compared with the top-performing methods.
+Therefore, it is promising to further improve both the performance and efficiency for practical applications.
+Figure 5: GPU memory consumption and run-time of state-of-the-art offline methods [31, 39, 104, 40, 100, 79, 111, 76, 86] on DTU dataset [21].
+The image resolution and the number of images are set to $1600\times 1152$ and 5 respectively.
+For fair comparison, experiments are done on one workstation with a NVIDIA 2080 Ti GPU.
+6.3 Unsupervised MVS with Depth Estimation
 We summarize the results of unsupervised MVS methods in Tab.
-III.
+II.
 Based on the conventional photometric consistency loss (Sec.
-III-A), unsupervised MVS methods may have difficulties with many challenging situations, e.g., varying lighting conditions or occlusions, and suffer from downgraded reconstruction results.
+4.1), unsupervised MVS methods may have difficulties with many challenging situations, e.g., varying lighting conditions or occlusions, and suffer from downgraded reconstruction results.
 Therefore, many methods propose to employ different strategies to improve the robustness, e.g., segmentation consistency [41], optical flow consistency [43], neural rendering consistency [42], pseudo-label generation with geometric filtering [128, 44], multi-iteration training [128, 44] and featuremetric loss [44].
@@
 Employing state-of-the-art architectures is also a potential direction that can further boost the performance.
-V-D Learning-based MVS without Depth Estimation
-V-D1 Voxel-based Methods
+6.4 Learning-based MVS without Depth Estimation
+6.4.1 Voxel-based Methods
 We summarize the quantitative results on ScanNet [26] in Tab.
-V.
+IV.
 Since the voxel representation increases the computation overhead explicitly when the scale of scene increases, voxel-based methods mainly focus on indoor scenes, e.g., ScanNet [26].
@@
 This further shows the advantages of depth map-based methods over voxel-based methods.
-TABLE V: Quantitative results of voxel-based MVS methods on ScanNet [26].
+TABLE IV: Quantitative results of voxel-based MVS methods on ScanNet [26].
 Methods Prec. $\uparrow$ Recall $\uparrow$ F-score $\uparrow$
@@
 VoRTX [135] 0.767 0.651 0.703
-TABLE VI: Quantitative results of NeRF-based MVS methods on DTU [21].
+TABLE V: Quantitative results of NeRF-based MVS methods on DTU [21].
 Methods VolSDF [48] NeuS [49] NeuralWarp [129] HF-NeuS [190] RegSDF [191] PET-NeuS [192]
 Overall $\downarrow$ 0.86 0.87 0.68 0.77 0.72 0.71
-Methods GeoNeuS [156] Voxurf [193] NeuS2 [157] PermutoSDF [159] Neuralangelo [158] UniSDF [160]
+Methods Geo-NeuS [50] Voxurf [193] NeuS2 [148] PermutoSDF [150] Neuralangelo [149] UniSDF [151]
 Overall $\downarrow$ 0.51 0.72 0.70 0.68 0.61 0.64
-V-D2 NeRF-based Methods
+6.4.2 NeRF-based Methods
 The quantitative results of nerf-based methods on DTU dataset [21] are summarized in Tab.
-VI.
-We visualize the qualitative reconstructions of two representative methods, NeuS [50] and VolSDF [49], in Fig.
-14.
+V.
 The most outstanding features of NeuS [49] and VolSDF [48] are that they are self-supervised, i.e., without depth supervision, and produce smooth and complete mesh surfaces.
@@
 In addition, the scalability of NeRF-based methods remains a problem when attempting to apply on large-scale scenes [23].
-Figure 14: Reconstruction results (rendered surface normal map and surface mesh) of NeuS [50] and VolSDF [49] on DTU dataset [21]. NeuS and VolSDF apply different approximation of density with SDF.
-V-D3 3D Gaussian Splatting-based Methods
-Because of its highly efficient rasterization, 3D Gaussian Splatting [170] becomes popular in novel view synthesis and gradually replaces NeRF [28] in the last half year.
-Recently, some researchers start exploring the potential of 3D Gaussian Splatting in surface reconstruction.
-However, there are many challenges.
+6.4.3 3D Gaussian Splatting-based Methods
+Because of its efficient rasterization, 3D Gaussian Splatting [160] becomes popular in novel view synthesis in the last half year.
+However, there are many challenges to use reconstruct surfaces with 3DGS.
 For example, the 3D Gaussians do not correspond well to the actual surface of the scene since the 3D Gaussian has a 3D covariance matrix [51].
 Moreover, the surface may contain noisy undulations [51].
-However, it is worthwhile to further explore in this direction since 3D Gaussian Splatting is much faster to train than those NeRF-based MVS methods.
-Moreover, instead of directly extracting the surface from 3D space [52], rendering accurate depth maps and then fuse them like depth map-based MVS methods is a potential direction.
-V-D4 Large Feed-forward Methods
-As a new trend of 3D reconstruction, large feed-forward methods demonstrate their impressive performance in challenging settings that other MVS methods cannot handle, e.g., single image input and un-posed sparse images.
+However, it is worthwhile to further explore in this direction since 3DGS is much faster to train than those NeRF-based MVS methods.
+Moreover, instead of directly extracting the surface from 3D space [51], rendering accurate depth maps and then fuse them like depth map-based MVS methods is a promising direction [52, 53].
+Additionally, generalizable feed-forward 3DGS methods [194, 195] are interesting since they are proven to improve depth accuracy with rendering loss and do not be trained for each scene.
+6.4.4 Large Feed-forward Methods
+As a new trend of 3D reconstruction, large feed-forward methods demonstrate impressive performance in challenging settings that other MVS methods cannot handle, e.g., single image input and un-posed sparse images.
 Employing a large model that is able to get scaled up, these methods manage to learn strong priors of multi-view geometry from large datasets [29].
-However, these methods [53, 179, 54] face limitations when it comes to extending the scale of the scene.
-For example, the 3D representation employed in LRM [53] is a tri-plane representation [180], which has a small resolution and struggles with scaling to larger sizes due to memory constraints.
-In addition, the tri-plane representation is difficult to model more complex geometry, e.g., with strong self-occlusions.
-In addition to tri-plane-based 3D representations, new methods [194, 195] have been proposed that directly reconstruct 3D Gaussians in a feed-forward manner.
-These new methods demonstrate improved capabilities in reconstructing both objects and scenes, but require posed sparse images as inputs.
-It is worth noting that the generalization ability of these methods is questionable since they learn priors of multi-view geometry completely from the training dataset, while the depth map-based methods utilize feature matching from plane sweep algorithm from traditional methods.
-Moreover, these methods typically need high-end GPUs since they are usually GPU memory consuming.
-It is worth noting that, since large feed-forward methods usually do not enforce geometric constraint explicitly, e.g., epipolar geometry, the reconstructed geometry is by default placed in a local coordinate system without absolute scales, making it not feasible for applications requiring precise measurement.
-V-E Future Research Directions
-V-E1 Datasets and Benchmarks
+Since these large feed-forward methods [54, 174, 55] usually do not enforce geometric constraint explicitly, i.e., epipolar geometry, the reconstructed geometry is by default placed in a local coordinate system without absolute scales, making it not feasible for applications requiring precise measurement.
+Also, these methods reconstruct the target under a globally shared representation, while depth-based MVS is typically performed without global data dependency, making conventional MVS still more suitable for large-scale reconstruction and parallel processing.
+The massive and costly computation of large feed-forward methods also limit the feasibility of applications.
+Benefited from the deep prior rooted in the large datasets, large feed-forwards methods are able to handle some of the typical corner cases, e.g., non-Lambertian surfaces.
+However it also makes their generalization ability questionable since they learn priors of multi-view geometry completely from the training dataset, while the depth map-based methods utilize feature matching via plane sweep algorithm from traditional methods that are more explainable.
+6.5 Future Research Directions
+6.5.1 Datasets and Benchmarks
 For learning-based MVS methods, ScanNet [26] and DTU [21] are two main training datasets, while ScanNet [26], DTU [21], Tanks & Temples [22] and ETH3D [23] are the main evaluation benchmarks.
@@
 TartanAir dataset [197] is a large-scale synthetic dataset that is collected in photo-realistic simulation environments.
-Precise multi-modal sensor data and ground truth labels are provided, including the stereo RGB image, depth image, segmentation, optical flow, camera poses, and LiDAR point cloud.
 BlendedMVS [24] introduces more large-scale scenes and improves the performance of many MVS methods on real-world scenes, shown in Tab.
-IV.
+III.
 ArKitScenes [198] consists of 5,048 RGB-D sequences, which is more than three times the size of the current largest available indoor dataset, ScanNet.
@@
 For evaluation, the main benchmarks [26, 21, 22, 23] are all introduced before 2018.
-Some of these benchmarks, e.g.,  [21, 22], are currently saturated and the learning-based methods are difficult to further improve the performance on them.
+Some of these benchmarks are currently saturated and the methods are difficult to further improve the performance on them.
 Therefore, it is meaningful to introduce new benchmarks to evaluate the robustness and performance of learning-based methods in challenging real-world scenes.
@@
 These make LaMAR a potential benchmark for MVS.
-V-E2 View Selection
+6.5.2 View Selection
 As discussed in Sec.
-I-B, view selection is crucial for triangulation quality.
+2.2, view selection is crucial for triangulation quality.
 Picking neighboring views that are suitable for triangulation can not only improve the reconstruction accuracy but also reduce useless computation for the bad views, e.g., views with strong occlusions.
@@
 For example, all offline depth map-based MVS methods [31, 39, 79, 100] follow MVSNet [15] and use the same simple heuristic strategy to compute scores for neighboring views and sort them.
-Online depth map-based MVS methods [58, 31, 35] and voxel-based methods [48] also adopt heuristic strategies to choose views with enough pose-distance.
-Though it is intractable to incorporate non-differentiable view selection into deep learning, it is worth exploring new view selection strategies since it may improve the reconstruction without changing the model design of existing methods.
+Online depth map-based MVS methods [60, 30, 34] and voxel-based methods [47] adopt heuristic strategies to choose views with enough pose-distance.
+Though it is intractable to incorporate non-differentiable view selection into deep learning, it is worth exploring new view selection strategies since it may improve the reconstruction without changing the model design.
 Current view selection strategies simply pick the same set of neighboring views for all the pixels in the reference image, and it is probable that different reference pixels have different optimal choices of neighboring views.
-If it is possible to choose the best neighboring views for each pixel, reconstruction quality can be further improved and computation reduced.
 The learned pixel-wise view weight [94, 39, 40] is one solution for this since it weights the source views differently for each pixel based on its visibility across neighboring views.
-V-E3 Depth Fusion
+6.5.3 Depth Fusion
 As discussed in Sec.
-I-D, both online and offline MVS methods with plane-sweep use traditional TSDF fusion or depth filtering methods to reconstruct mesh or point cloud from the estimated depth maps.
+2.4, both online and offline MVS methods with plane-sweep use traditional TSDF fusion or depth filtering methods to reconstruct mesh or point cloud from the estimated depth maps.
 However, TSDF fusion [65, 66] is not robust enough to the outliers in the depth maps and may have memory issues due to the dense volumetric representation.
@@
 However, they are limited to small scenes due to the dense volume representation and large memory consumption.
-V-E4 Features
+6.5.4 Features
 One not well-studied topic is what kind of feature extractors are suitable for MVS, as mentioned in Sec.
-II-A.
+3.1.
 So far, most of the offline depth map-based methods [15, 31, 78] mainly apply simple 2D CNN and FPN [72] structure as a feature extractor.
@@
 Finding a good balance between performance and efficiency is an interesting direction, e.g., distilling the large ViT into a relatively small transformer [206].
-V-E5 Real-time and Memory Efficient Models
+6.5.5 Real-time and Memory Efficient Models
 Though reconstruction accuracy is considered the most important factor in MVS, it is also critical to have models that can run in near real-time and with low memory, e.g., AR/VR and autonomous driving.
@@
 For offline MVS methods, as discussed in Sec.
-V-B2, some recent methods try to improve efficiency with carefully simplified network architectures.
+6.2.2, some recent methods try to improve efficiency with carefully simplified network architectures.
 However, this usually limits the performance when compared with state-of-the-art methods.
@@
 For example, using model compression techniques and knowledge distillation are potential directions to not only keep the good performance and also improve efficiency.
-V-E6 Prior Assistance
-MVS mainly relies on measurements of local photometric consistency to find the optimal matching across reference and source images.
-Accordingly, it usually encounters difficulties when estimating the geometry for regions where the photometric measurement becomes unreliable or invalid, e.g., textureless areas and non-Lambertian surfaces, which are common in human-made environments, e.g., ETH3D [23].
-Therefore, leveraging prior information to guide the MVS algorithm in these challenging regions is a promising research direction.
-We elaborate several typical examples of prior assistance as follows.
-  • Surface Normal: It is a non-local representation of the geometry compared with the depth map.
+6.5.6 Prior Assistance
+MVS mainly relies on assessing local photometric consistency to find the optimal matching across reference and source images.
+Accordingly, it usually encounters difficulties when estimating the geometry for regions where the photometric measurement becomes unreliable or invalid, e.g., textureless areas and non-Lambertian surfaces, which are common in human-made scenes [23].
+Therefore, using prior information to guide the MVS algorithm in these challenging regions is a promising research direction.
+We elaborate 3 typical examples of prior assistance as follows.
+Surface Normal: As a non-local representation of the geometry compared with the depth map, surface normal has been proven effective in improving reconstruction performance in recent works.
 To enforce the constraints of normal maps in depth estimation, Kusupati et al.
-[206] integrate the multi-view normal estimation network into the MVS pipeline.
+[207] integrate a multi-view normal estimation network into the MVS pipeline.
 Long et al.
@@
 Liu et al.
-[209] perform graph-based depth map optimization as post-processing.
-Since normal monocular normal predictions [210] provide high-quality priors for the full scene, [154, 155] use monocular normal to improve the optimization of neural implicit surfaces.
+[210] apply graph-based depth map optimization as post-processing.
+Since monocular normal network [211] is trained on large-scale dataset and can provide high-quality priors, [146, 187] use monocular normal priors to improve the surface accuracy.
 Shape Prior: For indoor scenes where common textureless areas, e.g., walls, planes are suitable choices of the geometric primitives and are exploited in traditional methods [20].
@@
 Some attempts [215, 216] employ a rule-based protocol to utilize semantic segmentation for better matching and depth quality.
-Manhattan-SDF [216] adopts Manhattan world priors to handle low-textured planar regions corresponding to walls and floors with 2D segmentation.
-VI Conclusion
-In this survey, we have provided a comprehensive review of the learning-based MVS methods until 2023, which are categorized into: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
+Manhattan-SDF [217] relies on 2D segmentation results to enforce Manhattan world priors to handle low-textured planar regions.
+Similarly, Shvets et al.
+[218] also employ SAM (Segment Anything Model) encoder to generate semantic features to enhance MVS.
+7 Conclusion
+In this survey, we have provided a comprehensive review of the learning-based MVS methods, which are categorized into: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
 Particularly, we have devoted significant attention to depth map-based methods because of their conciseness, flexibility and scalability.

### v3

@@
 Given a set of calibrated images, MVS aims to reconstruct dense 3D geometry for an observed scene.
-Based on the scene representations, traditional MVS methods can be divided into three categories: volumetric, point cloud, and depth map.
-Volumetric methods [9, 10, 11, 12] discretize the 3D space into voxels and label each as inside or outside of the surface.
-They are limited to scenes of small scale due to the large memory consumption.
-Point cloud methods [13, 14] operate directly on 3D points and often employ propagation to gradually densify the reconstruction.
-As point cloud propagation occurs sequentially, these methods are difficult to parallelize, leading to longer processing times.
-[15].
+Based on the scene representations, traditional MVS methods can be mainly divided into three categories: voxel, point cloud, and depth map [9].
+Voxel-based methods [10, 11, 12, 13] discretize the 3D space into voxels and label each as inside or outside of the surface.
+They are limited to small-scale scenes due to large memory consumption.
+Point cloud-based methods [14, 15] operate directly on 3D points and often employ propagation to gradually densify the reconstruction.
+As point cloud propagation occurs sequentially, these methods are difficult to parallelize, leading to longer processing times [9].
 In addition, the irregularity and large size of point cloud are also not very suitable for deep learning, especially in large-scale scenes.
-In contrast, methods relying on depth maps [16, 17, 18, 19, 20] use patch matching with photometric consistency to estimate the depth maps for individual images.
-Subsequently, these depth maps are fused into a dense representation, e.g., point cloud or mesh, as a post-processing step.
-Such design decouples the reconstruction problem into per-view depth estimation and depth fusion, which explicitly improves the flexibility and scalability.
+In contrast, methods that are based on depth maps [16, 17, 18, 19, 20] use patch matching with photometric consistency to estimate the depth maps for individual images.
+Subsequently, these 2D depth maps are fused into a dense 3D representation, e.g., point cloud or mesh.
+This design decouples the reconstruction task into per-view depth estimation and depth fusion, which explicitly improves flexibility and scalability.
+In addition, storing a set of 2D depth maps to represent 3D geometry is more convenient [21, 17].
 Although MVS has been studied extensively for several decades, traditional MVS methods rely on hand-crafted matching metrics and thus encounter challenges in handling various conditions, e.g., illumination changes, low-textured areas, and non-Lambertian surfaces [22, 23, 24, 25].
@@
 These methods have significantly outperformed traditional methods on various benchmarks [22, 23, 24, 27].
-In this survey, we categorize existing learning-based MVS methods based on their characteristics as follows: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
-Voxel-based methods estimate the geometry with volumetric representation and implicit function, e.g., Signed Distance Functions (SDF).
-They are limited to small-scale scenes due to the high memory consumption.
-NeRF and 3D Gaussian Splatting-based methods adapt NeRF [27] and 3D Gaussian Splatting [28], which are originally used for novel view synthesis, to extract surface from the implicit field or point cloud.
-They typically need to optimize the geometry for each new scene, which needs lots of run-time and memory consumption.
-Large feed-forward methods typically use large transformer models to directly learn the 3D representation from given images.
-They require massive computation because of the huge network, and are mainly limited to object-level scenes [29].
-Depth map-based methods utilize deep learning in depth estimation and then fuse depth maps with traditional fusion algorithm.
-Comparatively, depth map-based methods [15, 25] are the main family of learning-based methods since they inherit the advantages from those traditional methods [17, 18, 19, 20] by decoupling 3D reconstruction into depth estimation and depth fusion.
-Therefore, we focus more on the depth map-based methods and discuss them in details.
-For clarity, we further categorize depth map-based methods into online and offline methods, shown in Fig.
+In this survey, motivated by previous works [21, 15, 9], we categorize learning-based MVS methods based on their underlying representations as follows: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based, and large feed-forward methods.
+Voxel-based methods estimate geometry with volumetric representation and Signed Distance Functions (SDF).
+They are limited to small-scale scenes because of the high memory storage of voxel grids.
+NeRF and 3D Gaussian Splatting-based methods adapt NeRF [28] and 3D Gaussian Splatting [29], which are used for novel view synthesis, and extract the surface from the volumetric implicit field and explicit 3D Gaussians points, respectively.
+They typically need to optimize the geometry for each scene.
+Large feed-forward methods typically use large models to directly learn the 3D representation from given images.
+They require massive computation because of the large network.
+Depth map-based methods introduce deep learning in depth estimation and then fuse depth maps with the traditional fusion algorithm.
+Similar to traditional MVS, depth map-based methods [9, 26] are the main family of learning-based MVS since they inherit the advantages of those traditional methods [30] by decoupling 3D reconstruction into depth estimation and depth fusion.
+Therefore, we focus more on depth map-based MVS in this survey.
+For clarity, we further categorize depth map-based MVS into online and offline methods, shown in Fig.
 1.
-Specifically, online MVS methods [25, 30] usually perform multi-view depth estimation with a video sequence of low-resolution images, e.g., ScanNet [26].These methods typically rely on low-resolution inputs and lightweight network structures to ensure fast inference, prioritizing speed and simplicity.
-The main goal of online MVS is to deliver real-time reconstruction for time-sensitive applications such as augmented reality and live video processing.
-In contrast, offline MVS methods [15, 31] prioritize better quality reconstruction at the cost of computation.
-These methods excel in multi-view depth estimation and point cloud fusion using high-resolution image sets, e.g., DTU [21], Tanks and Temples [22] and ETH3D [23].
-By operating on higher-resolution inputs and employing complex network, they thoroughly analyze the multi-view information in the images.
-The emphasis on achieving high accuracy and capturing intricate details often comes at the cost of requiring considerable computational resources and time.
-As a result, offline MVS is frequently employed in applications demanding precise and photorealistic scene representations, including 3D modeling and archaeological reconstruction.
-In summary, our survey covers the most recent literature on learning-based MVS methods, including four main families: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
+Online methods [26, 31] reconstruct scenes in an online manner with local temporal information, while offline methods [9, 32] reconstruct scenes in an offline manner with global information.
+To estimate depth for a given reference view, online methods select its best views in a local buffer that stores previous views.
+In contrast, offline methods select best views among all possible views in the database.
+In summary, our survey covers the most recent literature on learning-based MVS methods, including five main families: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based, and large feed-forward methods.
 We provide a comprehensive review and insights on different aspects, including the pipelines and algorithmic intricacies.
-Moreover, we summarize the performance of the reviewed methods on different benchmarks, and discuss the potential future directions for deep learning-based MVS methods.
+Moreover, we evaluate the performance and efficiency of the reviewed methods, and discuss the potential future directions for deep learning-based MVS.
 Figure 1: An overall illustration of both online and offline depth map-based MVS pipelines.
@@
 In this section, we introduce these components to provide readers a clear picture of the MVS.
-Note that camera calibration and view selection are components for other learning-based methods as well.
+Note that camera calibration and view selection are generally required by other methods as well.
 Sec.
@@
 2.5 lists common datasets and benchmarks for MVS and Sec.
-2.6 summarizes the common evaluation metrics.
-{forest}
-for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, draw=black, rounded corners,align=left, minimum width=2.5em, minimum height=1.2em, s sep=6pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=8em,font=, where level=3font=, where level=4font=, where level=5font=, [MVS [Traditional Methods,text width=10 em [COLMAP [33], P-MVS [14], Gipuma [18]…,leaf,text width=16.3 em] ] [Datasets & Benchmarks, text width=10 em [ScanNet [26], DTU [21], Tanks and Temples [22]…, leaf,text width=16.3 em] ] [Learning-based Methods
-with Depth Estimation,text width=10 em [Pipeline,text width=6 em [Camera Calibration, View Selection, Multi-View Depth Estimation, Depth Fusion,leaf,text width=27.4 em] ] [Supervised,text width=6em [Online Methods [MVDepthNet [25], DeepvideoMVS [30], SimpleRecon [34]…,leaf,text width=20.2 em] ] [Offline Methods [Direct 3D CNN,text width=5.4 em [MVSNet [15], CIDER [35]…,leaf,text width=13 em] ] [RNN,text width=5.4 em [R-MVSNet [36], $D^{2}$-RMVSNet [37]…,leaf,text width=13 em] ] [Coarse-to-fine,text width=5.4 em [CasMVSNet [31], UCSNet [38]…,leaf,text width=13 em] ] [Iterative Update,text width=5.4 em [PatchmatchNet [39], IterMVS [40]…,leaf,text width=13 em] ] ] ] [Unsupervised,text width=6em [End-to-end,text width=5 em [ JDACS [41], RC-MVSNet [42]…,leaf,text width=10.5 em] ] [Multi-stage,text width=5 em [U-MVSNet [43], KD-MVS [44]…,leaf,text width=10.5 em] ] ] [Semi-Supervised,text width=6em [ SGT-MVSNet [45],leaf,text width=6.2 em] ] ] [Learning-based Methods
-without Depth Estimation,text width=10em [Voxel-based,text width=10 em [Atlas [46], NeuralRecon [47]…, leaf,text width=13 em] ] [NeRF-based,text width=10 em [VolSDF [48], NeuS [49], Geo-NeuS [50]…, leaf,text width=13 em] ] [3D Gaussian Splatting-based,text width=10 em [ SuGaR [51], 2DGS [52], PGSR [53]…, leaf,text width=13 em] ] [Large feed-forward,text width=10 em [LRM [54], DUSt3R [55], MASt3R [56]…, leaf,text width=13 em] ] ] [Future Research Directions, text width=10 em [Dataset & Benchmarks, View Selection, Depth Fusion, Features, Run-time & Memory Efficiency, Prior Assistance, leaf,text width=37.5 em] ] ]
+2.6 summarizes common evaluation metrics.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+MVS
+Traditional Methods COLMAP [30], P-MVS [15], Gipuma [18] Datasets & Benchmarks ScanNet [27], DTU [22], Tanks and Temples [23] Learning-based MVS with Depth Representation Pipeline Camera Calibration, View Selection, Multi-View Depth Estimation, Depth Fusion Supervised
+Online Methods
+MVDepthNet [26], DeepvideoMVS [31], SimpleRecon [34]
+Offline Methods
+Direct 3D CNN MVSNet [9], CIDER [35] RNN R-MVSNet [36], $D^{2}$HC-RMVSNet [37] Coarse-to-fine CasMVSNet [32], UCSNet [38] Iterative Update PatchmatchNet [39], IterMVS [40] Unsupervised End-to-end JDACS [41], RC-MVSNet [42] Multi-stage U-MVSNet [43], KD-MVS [44] Semi-supervised SGT-MVSNet [45] Learning-based MVS with Other Representations Voxel Direct 3D CNN Atlas [46] Coarse-to-fine NeuralRecon [47], TransformerFusion [48] NeRF Optimization NeuS [49], Neuralangelo [50], UniSDF [51] Generalizable SparseNeuS [52], VolRecon [53], LRM [54] 3D Gaussian Splatting Optimization SuGaR [55], 2DGS [56], PGSR [57] Generalizable MVSplat [58], DepthSplat [59], LGM [60] Large Feed-forward Point-based DUSt3R [61], VGGT [62], MapAnything [63] Future Directions Dataset & Benchmarks, View Selection, Feature Extraction, Generative Reconstruction, Efficiency, Prior Assistance
 Figure 2: Taxonomy of Multi-View Stereo.
 2.1 Camera Calibration
-[…]
+Camera calibration is a process of determining the intrinsic and extrinsic parameters of a camera to understand its geometry and characteristics accurately [64].
+It serves as the foundational step in MVS, ensuring that the subsequent reconstruction process is built on accurate and consistent geometric information, ultimately leading to a more reliable and precise 3D representation of the scene.
+Typically, obtaining calibrated camera parameters is usually achieved by running off-the-shelf SfM algorithms [17, 65] or SLAM [66], which jointly optimize sparse triangulated 3D points and camera parameters.
+The camera parameters include the extrinsic matrix $\mathbf{T}=[\mathbf{R}|\mathbf{t}]$ and intrinsic matrix $\mathbf{K}$.
+Depth map-based MVS methods [9, 32, 39] require a bounded depth range $d_{\textrm{min}},d_{\textrm{max}}]$ to improve the estimation accuracy.
+For offline methods [9, 32], the depth range can be estimated by projecting the sparse point cloud from SfM to each viewpoint and compute the minimum and maximum $z$ values [9].
+In contrast, online methods [26, 31] usually set constant depth ranges, e.g., [0.25mm, 20.00mm], since the scene scale is usually fixed as room.
 2.2 View Selection
@@
 Currently, there are two main strategies for view selection.
-First, for most online MVS depth estimation methods [25, 60, 30], a frame is selected as a keyframe when its pose has sufficient difference compared with the previous keyframe.
-Then each keyframe adopts several previous keyframes to estimate depth.
-GP-MVS [60] proposes a heuristic pose-distance measure as:
-$\text{disc}(\mathbf{T}_{ij})=\sqrt{||\mathbf{t}_{ij}||^{2}+\frac{2}{3}\text{tr}(\mathbbm{I}-\mathbf{R}_{ij})},$ (1)
-where $\mathbf{T}_{ij}=[\mathbf{R}_{ij}|\mathbf{t}_{ij}]$ is the relative transformation between view $i$ and $j$. This strategy is used by many following methods [30, 34].
-Second, for most offline MVS methods [15, 36, 31], view selection is done with the sparse point cloud obtained by Structure-from-Motion [33, 58].
-For a reference view $i$, MVSNet [15] computes a score $s(i,j)=\sum_{\mathbf{P}}\eta(\theta_{ij}(\mathbf{P}))$ for the neighboring view $j$, where $\mathbf{P}$ is a 3D point observed by both view $i$ and $j$.
-$\theta_{ij}(\mathbf{P})=(180/\pi)\arccos{((\mathbf{c}_{i})-\mathbf{P})\cdot((\mathbf{c}_{j})-\mathbf{P}))}$ represents the baseline angle for $\mathbf{P}$ and $\mathbf{c}_{i},\mathbf{c}_{j}$ are the camera centers.
-$\eta(\cdot)$ is a piece-wise Gaussian function [61] to favor a certain baseline angle $\theta_{0}$:
-$\eta(\theta)=\left\{\begin{aligned} \exp{(-\frac{(\theta-\theta_{0})^{2}}{2\sigma_{1}^{2}})},\theta\leq\theta_{0}\\ \exp{(-\frac{(\theta-\theta_{0})^{2}}{2\sigma_{2}^{2}})},\theta>\theta_{0}\end{aligned}\right.$ (2)
-where $\theta_{0}$, $\sigma_{0}$, $\sigma_{1}$ are hyper-parameters.
+First, for most online MVS methods [26, 67, 31, 34], a frame is selected as a keyframe when its pose has sufficient difference compared with the previous keyframe.
+Within a keyframe buffer, each new keyframe selects several old keyframes with smallest pose difference to estimate depth [31].
+To measure this pose difference, GP-MVS [67] proposes a heuristic measure as:
+[…]
+where $\mathbf{T}_{ij}=[\mathbf{R}_{ij}|\mathbf{t}_{ij}]$ is the relative transformation between view $i$ and $j$.
+A new keyframe is added if the pose-distance $\text{disc}(\mathbf{T}_{ij})$ from the most recent keyframe is above a certain threshold [31, 34].
+Second, for most offline MVS methods [9, 36, 32], view selection is done with the sparse point cloud obtained by SfM [30, 65].
+For a reference view $i$, MVSNet [9] computes a score $s(i,j)=\sum_{\mathbf{P}}\eta(\mathbf{\Theta}_{ij}(\mathbf{P}))$ for the neighboring view $j$, where $\mathbf{P}$ is a 3D point observed by both view $i$ and $j$.
+$\mathbf{\Theta}_{ij}(\mathbf{P})$ is the baseline angle for $\mathbf{P}$, which is computed as:
+$\mathbf{\Theta}_{ij}(\mathbf{P})=(180/\pi)\arccos{((\mathbf{c}_{i})-\mathbf{P})\cdot((\mathbf{c}_{j})-\mathbf{P}))},$ (2)
+where $\mathbf{c}_{i},\mathbf{c}_{j}$ are the camera centers. $\eta(\cdot)$ is a piece-wise Gaussian function [68] to favor a certain baseline angle $\theta_{0}$:
+$\eta(\theta)=\left\{\begin{aligned} \exp{(-\frac{(\theta-\theta_{0})^{2}}{2\sigma_{1}^{2}})},\theta\leq\theta_{0}\\ \exp{(-\frac{(\theta-\theta_{0})^{2}}{2\sigma_{2}^{2}})},\theta>\theta_{0}\end{aligned},\right$.
+(3)
+where $\theta_{0}$ is the favored baseline angle, and $\sigma_{0}$ and $\sigma_{1}$ are weights.
 Then the view selection is done by choosing neighboring views with highest scores.
@@
 2.3 Multi-view Depth Estimation with Plane Sweep
-To form a structured data format that is more suitable for convolution operations, most learning-based MVS methods rely on the plane sweep algorithm [32] to calculate matching costs.
+To form a structured data format that is more suitable for convolution operations, most depth map-based MVS methods rely on the plane sweep algorithm [33] to calculate matching costs.
 The plane sweep algorithm discretizes the depth space with a set of fronto-parallel planes along the depth direction.
@@
 It operates by sweeping a conceptual plane through the object space and evaluating the spatial distribution of geometric surfaces.
-In practice, we divide the depth range, which is either manually set [25] or estimated by SfM [15], into discrete samples and assign hypotheses at these values.
-To map coordinates at depth hypothesis $d$, homography transformation [15] $\mathbf{p}_{i}(d)\sim\mathbf{H}_{i}(d)\cdot\mathbf{p}$, is applied, where $\mathbf{p}_{i}(d)$ is the corresponding pixel in the $i$-th source view for pixel $\mathbf{p}$ of the reference view.
+In practice, we divide the depth range, which is manually set [26] or estimated by SfM [9], into discrete samples and assign hypotheses at these values.
+To map coordinates with depth hypothesis $d$, homography transformation [9] $\mathbf{p}_{i}(d)\sim\mathbf{H}_{i}(d)\cdot\mathbf{p}$, is applied, where $\mathbf{p}_{i}(d)$ is the corresponding pixel in the $i$-th source view for the pixel $\mathbf{p}$ of the reference view.
 The homography $\mathbf{H}_{i}(d)$ between the $i$-th source view and the reference view 0 can be computed as:
 $\mathbf{H}_{i}(d)=\mathbf{K}_{i}\mathbf{R}_{i}\left(\mathbbm{I}-\frac{(-\mathbf{R}_{0}^{\top}\mathbf{t}_{0}+\mathbf{R}_{i}^{\top}\mathbf{t}_{i})\mathbf{n}^{\top}\mathbf{R}_{0}}{d}\right)\mathbf{R}_{0}^{\top}\mathbf{K}_{0}^{-1},$ (4)
-where $\mathbf{K}_{0}$, $\mathbf{K}_{i}$ denote camera intrinsics, $[\mathbf{R}_{0}|\mathbf{t}_{0}]$, and $[\mathbf{R}_{i}|\mathbf{t}_{i}]$ denote camera extrinsics, $\mathbf{n}$ denote the principle axis of the reference view.
-Equivalently, we can also project a reference pixel into source views with depth hypothesis [35, 39, 40].
-We compute $\mathbf{p}_{i}(d)$ in the $i$-th source view for pixel $\mathbf{p}$ in the reference and depth hypothesis $d$ as follows:
+where $\mathbf{K}_{0}$, $\mathbf{K}_{i}$ denote camera intrinsics, $\mathbf{R}_{0}|\mathbf{t}_{0}]$, $\mathbf{R}_{i}|\mathbf{t}_{i}]$ denote camera extrinsics, $\mathbf{n}$ denotes the principal axis of the reference view.
+Equivalently, we can also project a reference pixel in the source views with depth hypothesis [35, 39, 40].
+We compute $\mathbf{p}_{i}(d)$ in the $i$-th source view for pixel $\mathbf{p}$ in the reference and depth $d$ as follows:
 $\mathbf{p}_{i}(d)=\mathbf{K}_{i}\cdot\left(\mathbf{R}_{i}\mathbf{R}_{0}^{\top}\cdot(\mathbf{K}_{0}^{-1}\cdot\mathbf{p}\cdot d)-\mathbf{R}_{i}\mathbf{R}_{0}^{\top}\cdot\mathbf{t}_{0}+\mathbf{t}_{i}\right).$ (5)
-The warped source feature maps are then obtained via differentiable interpolation.
-To better elaborate the process of plane sweep, we include a toy example in the supplementary.
+The warped source feature is then obtained via differentiable bilinear interpolation of the source feature map.
+To better elaborate the process of plane sweep, we include a toy example in the supplement.
 The photometric similarity (or matching cost) is measured in a one-to-many manner between reference and warped source features for the following depth estimation, which will be discussed in Sec.
 3.2.
-By performing plane sweep, we obtain a cost volume with a regular spatial shape, making it easy for CNNs to process in parallel.
 2.4 Depth Fusion
 For depth map-based MVS, after estimating all the depth maps, we need to fuse them into a dense 3D representation, e.g., point cloud or mesh.
-Online MVS methods [30, 34] usually adopt TSDF (Truncated Signed Distance Function) fusion [65, 66] to fuse the depth maps into a TSDF volume and then use Marching Cube [67] to extract the mesh.
-However, there usually exist outliers in the depth maps, which may reduce the reconstruction accuracy.
-To overcome this problem and improve the accuracy, offline MVS methods [15, 31, 39] typically filter the depth maps before fusing into a point cloud, which is motivated by Galliani et al.
-[18].
-There are two main filtering steps: photometric consistency filtering and geometric consistency filtering [15].
-For more details, please refer to the supplementary.
+Online MVS methods [31, 34] usually adopt TSDF (Truncated Signed Distance Function) fusion [72, 6] to sequentially fuse depth maps into a TSDF volume and then use Marching Cube [73] to extract the mesh.
+However, there usually exist outliers in the depth maps, which may reduce the accuracy of reconstruction.
+To adress this problem, offline MVS methods [9, 32, 39] filter out outliers before converting the depth maps to point clouds.
+Depth filtering involves photometric and geometric filtering [9] (for more details, please refer to the supplement).
+Note that online MVS methods do not perform depth filtering before TSDF fusion.
+For depth fusion, offline MVS methods use the visibility-based fusion algorithm [74] to minimize depth occlusions and violations across different viewpoints.
+The fused depth maps are then directly reprojected to space to generate the 3D point cloud [9].
 2.5 Datasets and Benchmarks
-Along with this survey, we also include a brief summary of commonly used public MVS datasets and benchmarks for training and evaluation. For more details, please refer to supplementary.
+High-quality and large-scale 3D datasets and benchmarks are important for both training and evaluation. We briefly summarize the commonly used MVS datasets and benchmarks in  Tab. I.
+SUN3D [75] is a RGB-D video dataset for large indoor scenes, which provides RGB-D images, camera poses, segmentations, and point clouds.
+ScanNet [27] is a large RGB-D dataset that contains 1613 indoor scenes with ground-truth camera poses, depth maps, surface reconstruction, and semantic segmentation labels.
+Online MVS methods [26, 31, 34] mainly use ScanNet for training and testing.
+7-Scenes [76] is a RGB-D dataset captured in indoor scenes with a handheld Kinect RGB-D camera.
+Since it is relatively small, 7-Scenes is usually used to test the generalization performance of the models trained on ScanNet [27] without finetuning [31].
+DTU [22] is an object-centric MVS dataset collected under well-controlled laboratory conditions with known camera trajectory.
+It contains 128 scans with 49 or 64 views under 7 different lighting conditions.
+Since DTU dataset provides scanned ground truth point clouds instead of depth maps, it is required to generate mesh models with surface reconstruction, e.g., screened Poisson surface reconstruction algorithm [77], and then render depth maps [9] for training.
+Tanks and Temples [23] is a large-scale benchmark captured in more complex real indoor and outdoor scenarios.
+It is divided into intermediate and advanced sets.
+Different scenes have different scales, surface reflection and exposure conditions.
+Note that Tanks and Temples does not provide ground truth camera parameters, which are usually estimated with Structure-from-Motion [17, 65].
+ETH3D [24] contains 25 large-scale indoor and outdoor scenes with high-resolution RGB images.
+The scenes typically contain many low-textured regions and non-Lambertian surfaces, e.g., white walls and reflective floor.
+In addition, the images are usually sparse and have strong viewpoint variations and occlusions.
+Therefore, ETH3D is considered as a very challenging benchmark.
+BlendedMVS [25] is a large-scale synthetic dataset for MVS training.
+It contains a variety of scenes, such as cities, sculptures, and shoes.
+The dataset consists of more than 17kk high-resolution images rendered with reconstructed models and is split into 106 training scenes and 7 validation scenes.
+Since images are rendered through virtual cameras, the camera parameters are accurate enough for training.
+ScanNet++ [78] is a large-scale and high-resolution indoor dataset with 3D reconstructions, high-quality RGB images, commodity RGB-D video.
+Compared to ScanNet [27], the quality of color and geometry is higher.
+TABLE I: An overview of commonly used datasets and benchmarks for learning-based Multi-View Stereo.
+Dataset Year Num. Scene Complexity Provided Ground Truth Evaluation
+Scenes Indoor Outdoor Camera Pose Depth Point Cloud Mesh Target
+SUN3D [75] 2013 415 Depth
+7-Scenes [76] 2013 7 Depth
+DTU [22] 2016 124 Point Cloud
+ScanNet [27] 2017 1503 Depth / Mesh
+Tanks and Temples [23] 2017 21 Point Cloud1
+ETH3D [24] 2017 25 Point Cloud1
+BlendedMVS [25] 2020 113 Depth
+ScanNet++ [78] 2023 460 Depth / Mesh
+  • 1
+    For datasets with online benchmark, the point cloud ground truth of test set is not released.
 2.6 Evaluation Metrics
 Based on the ground truth, e.g., depth maps or point clouds, evaluation can be categorized into 2D and 3D metrics.
-2D Metrics: 2D depth-based metrics are commonly used by online MVS methods [25, 30] to evaluate the depth maps [68, 69].
-The common metrics are: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel) and inlier ratio with threshold 1.25 ($\delta<1.25$).
-For more details, please refer to supplementary.
-3D Metrics: 3D point cloud evaluation is widely used by offline MVS methods [15, 36].
-Note that the reconstructed point clouds should be aligned to ground truth point clouds before evaluation, e.g., by Iterative Closest Point (ICP), if their extrinsics are differently calibrated.
+2D Metrics: 2D metrics are used to evaluate depth maps [79, 80].
+The common metrics are: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel), and inlier ratio with threshold 1.25 ($\delta<1.25$).
+For more details, please refer to the supplement.
+3D Metrics: 3D metrics are used to evaluate the accuracy of point cloud.
+Note that the reconstructed point clouds should be aligned with ground truth point clouds before evaluation, e.g., by Iterative Closest Point (ICP), if their extrinsics are differently calibrated.
 The metrics are Precision/Accuracy, Recall/Completeness and F-score.
-For more details, please refer to supplementary.
-3 Supervised Method with Depth Estimation
-This section mainly introduces supervised learning-based MVS methods with depth estimation.
-A typical depth map-based MVS pipeline mainly consists of the feature extraction (Sec.
+For more details, please refer to the supplement.
+3 Supervised MVS with Depth Representation
+This section mainly introduces supervised MVS methods with depth representation.
+As shown in Fig.
+3, a typical depth map-based MVS pipeline mainly consists of feature extraction (Sec.
 3.1), cost volume construction (Sec.
 3.2), cost volume regularization (Sec.
-3.3) and depth estimation (Sec.
+3.3), and depth estimation (Sec.
 3.5).
-The pipelines of MVDepthNet [25] and MVSNet [15] are illustrated in Fig.
-3 and Fig.
-4, which are the representative methods of online and offline MVS methods respectively.
 3.1 Feature Extraction
-Considering efficiency, most methods use simple CNN structures to extract deep features from images, e.g., ResNet [70], U-Net [71] and FPN [72].
-For online MVS methods, feature extraction networks are usually used in line with the real-time operation goal.
+Considering efficiency, most methods use simple CNN structures to extract deep features $\left\{\mathbf{F}_{i}\right\}_{i=0}^{N-1}$ from images, e.g., ResNet [81], U-Net [82] and FPN [83].
+For online MVS methods, feature extraction networks are usually chosen in line with the real-time operation goal.
 DeepVideoMVS [31] combines MNasNet [84], which is lightweight and has low latency, with FPN.
-SimpleRecon [34] proposes utilizing the first two blocks from ResNet18 [70] and an EfficientNet-v2 [74] encoder, which maintains efficiency and yields a sizeable improvement in depth map accuracy.
-For offline MVS methods, MVSNet [15] uses a stacked 8-layer 2D CNN to extract deep features for all the images.
+SimpleRecon [34] and DoubleTake [85] use the first two blocks of ResNet18 [81] for the cost volume and an EfficientNet-v2 [86] encoder for the monocular image prior, which maintains efficiency and yields a significant improvement in depth map accuracy.
+For offline MVS methods, MVSNet [9] uses a stacked 8-layer 2D CNN to extract deep features for all images.
 Coarse-to-fine methods further extract multi-scale features for estimation on multiple scales with FPN [32, 38, 39] or multi-scale RGB images [87, 88].
-Recently, many following works have been paying more attention to feature extraction to improve the representation power of deep features.
+Recently, many following works pay more attention to feature extraction to improve the representation power of deep features.
 [89, 90, 91] introduce deformable convolutions to adaptively learn receptive fields for areas with varying richness of texture.
-CDS-MVSNet [80] uses a dynamic scale network, CDSFNet, to extract the discriminative features by analyzing the normal curvature of the image surface.
-GeoMVSNet [76] proposes a geometry fusion network that fuses the image features with coarse depth maps.
-With FPN as the main feature extractor, WT-MVSNet [81] and MVSFormer [82], further introduces Vision Transformers [83, 84, 85] for image feature enhancement.
-ET-MVSNet [86] integrates the Epipolar Transformer into the feature extraction, which performs non-local feature aggregation on the epipolar lines.
+TransMVSNet [91] adpots a Feature Matching Transformer for robust long-range global context aggregation within and across images.
+With FPN as the main feature extractor, WT-MVSNet [92] and MVSFormer++ [93], further introduce Vision Transformers [94, 95, 96] for image feature enhancement.
 3.2 Cost Volume Construction
-For both online and offline MVS methods, the cost volume is constructed with the plane sweep algorithm as discussed in Sec. 2.3.
+For both online and offline MVS, the cost volume is constructed with the plane sweep algorithm as discussed in Sec. 2.3 with depth samples $\mathbf{D}\in\mathbb{R}^{H\times W\times D}$.
 3.2.1 Online MVS
-To reduce computation and improve the efficiency for online applications, online MVS methods usually construct 3D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D}$, which store a single value as matching cost for each pixel $\mathbf{p}$ and depth sample $d_{i}$.
-MVDepthNet [25] and GP-MVS [60] compute the per-pixel intensity difference between the reference and each source view as matching cost.
-If there is more than one source view, the cost volumes are averaged.
-Instead of intensity difference, Neural RGB-D [87] computes deep feature difference.
-DeepVideoMVS [30] and MaGNet [88] calculate the cost from pixel-wise correlation between the reference feature and the warped source feature.
-In addition to the dot product between reference image features and warped source image features, SimpleRecon [34] further adds metadata, e.g., ray direction and relative pose, into the 4D cost volume.
-Then an MLP is used to reduce the dimension of 4D cost volume to a 3D cost volume.
+To reduce computation and improve the efficiency for online applications, online MVS methods usually construct 3D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D}$.
+The cost volume stores a single value as matching cost for each pixel $\mathbf{p}$ and depth sample $d$, which we denote as $\mathbf{C}(\mathbf{p},d)$.
+For example, MVDepthNet [26] and GP-MVS [67] compute the per-pixel intensity difference between the reference and each source view and average the matching costs from all source views:
+$\mathbf{C}(\mathbf{p},d)=\frac{1}{N-1}\sum_{i=1}^{N-1}AD(\mathbf{I}_{0}(\mathbf{p}),\mathbf{I}_{i}(\mathbf{p}_{i}(d))),$ (6)
+where $AD(\cdot,\cdot)$ denotes the absolute intensity difference, $\mathbf{p}_{i}(d)$ is the reprojected pixel computed from Eq.
+5.
+Instead of intensity difference, Neural RGB-D [97] computes the $L_{2}$ norm of the difference between the reference and warped source features as the matching cost.
+DeepVideoMVS [31] computes the dot product of the reference and warped source features:
+$\mathbf{C}(\mathbf{p},d)=-\frac{1}{C\times(N-1)}\sum_{i=1}^{N-1}\langle\mathbf{F}_{0}(\mathbf{p}),\mathbf{F}_{i}(\mathbf{p}_{i}(d))\rangle,$ (7)
+where $\mathbf{F}_{0},\mathbf{F}_{i}$ are image features, and $C$ is the number of feature channels.
+Based on DeepVideoMVS, SimpleRecon [34] further introduces metadata, e.g., ray direction and relative pose, into the cost volume, which is found to increase the accuracy.
 DoubleTake [85] further includes depth hint, the depth map rendered from the previously predicted geometry, into the cost volume.
 3.2.2 Offline MVS
-Offline MVS methods focus on reconstructing high-quality dense geometry with high-resolution images.
-To encode more matching information and improve the quality, offline methods [15, 31] usually construct 4D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D\times C}$, where each pixel $\mathbf{p}$ and depth sample $d_{i}$ is associated with a matching cost of dimension $C$.
-Since there may exist an arbitrary number of source views and serious occlusion [23], robustly aggregating matching information from all the source views is an important step.
-MVSNet [15] proposes a variance-based cost metric as follows:
+To encode more matching information and improve reconstruction quality, offline MVS methods [9, 32] usually construct 4D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D\times C}$, where each pixel $\mathbf{p}$ and depth sample $d$ is associated with a matching cost of dimension $C$.
+Since there may exist an arbitrary number of source views and serious occlusion [24], robustly aggregating matching information from all source views is an important step.
+MVSNet [9] proposes a variance-based cost volume as follows:
 $\mathbf{C}={\text{Var}}\left(\mathbf{V}_{0},\cdots,\mathbf{V}_{N-1}\right)=\frac{\sum_{i=0}^{N-1}\left(\mathbf{V}_{i}-\overline{\mathbf{V}}\right)^{2}}{N},$ (8)
-where $\{\mathbf{V}_{i}\}_{i=0}^{N-1}$ are (warped) feature volumes, $\overline{\mathbf{V}}$ is the average feature volume.
-To further reduce dimension, CIDER [35] adopts group-wise correlation [90] to compute a lightweight cost volume between reference and each warped source view.
-Then $N-1$ cost volumes are averaged for regularization.
+where $\mathbf{V}_{0}$ is the reference feature volume, and $\{\mathbf{V}_{i}\}_{i=i}^{N-1}$ are warped feature volumes of source views, $\overline{\mathbf{V}}=1/N\sum_{i=0}^{N-1}\mathbf{V}_{i}$ is the average feature volume.
+Specifically, $\mathbf{V}_{0}(\mathbf{p},d)=\mathbf{F}_{0}(\mathbf{p})$, $\mathbf{V}_{i}(\mathbf{p},d)=\mathbf{F}_{i}(\mathbf{p}_{i}(d))$.
+To reduce computation, CIDER [35] adopts group-wise correlation [71] to compute a lightweight cost volume between reference and warped features from N1N-1 source views.
+Then the individual cost volumes for different source images are averaged as the final cost volume $\mathbf{C}$.
 Without considering occlusions, these methods consider all source views equally with the averaging operation.
@@
 SimpleRecon [34] fuses multi-scale image features, extracted from the pretrained EfficientNetv2 [86], into the cost volume encoder to improve the performance.
-Figure 3: Pipeline of MVDepthNet [25].
-Multiple image frames are encoded in the cost volume.
-MVDepthNet takes the reference image and the cost volume as the input and outputs inverse depth maps of four different resolutions.
-Skip connections between the encoder and decoder of the same resolution are used for the per-pixel depth estimation.
-Figure 4: Pipeline of MVSNet [15].
-Reference and source images go through the feature extraction network, followed by the differentiable homograph warping to construct the cost volume.
-A 3D U-Net is used to regularize the cost volume into a probability volume.
-The final depth map is estimated from the probability volume and refined with the reference image features.
+Figure 3: Pipeline of depth map-based MVS, which usually consists of feature extraction, cost volume construction via plane sweep, cost volume regularization, depth estimation, and depth refinement.
 3.3.2 Offline MVS
-Among most learning-based offline MVS methods that use 4D cost volumes, there are three main categories to perform cost volume regularization: direct 3D CNN, coarse-to-fine and RNN.
+Among most offline MVS methods that use 4D cost volumes, there are three main categories of cost volume regularization: direct 3D CNN, RNN, and coarse-to-fine.
 Direct 3D CNN: Similar to stereo estimation [69, 70, 71], 3D CNN is widely used for cost volume regularization in MVS [9, 103, 104, 105, 35].
-As the blueprint of learning-based offline MVS methods, MVSNet [15] adopts a 3D U-Net [71] to regularize the cost volume, which aggregates context information from a large receptive field with relatively low computation cost.
+While this concept has been explored for stereo matching in works like DPSNet [103], MVSNet [9] serves as the blueprint for learning-based offline MVS methods.
+It adopts a 3D U-Net [82]to regularize the cost volume, which aggregates context information from a large receptive field with relatively low computation cost.
 It is found that the 3D regularization can capture better geometry structures, perform photometric matching in 3D space, and alleviate the influence of image distortion caused by perspective transformation and occlusions [104].
@@
 RNN: Instead of 3D CNN, R-MVSNet [36] sequentially regularizes 2D slices of the cost volume along the depth dimension with a convolutional GRU [106], which is able to gather spatial and uni-directional context information in the depth dimension.
-$D^{2}$-RMVSNet [37] augments R-MVSNet [36] with a complex convolutional LSTM [95].
+$D^{2}$HC-RMVSNet [37] augments R-MVSNet [36] with a complex convolutional LSTM [102].
 AA-RMVSNet [89] further introduce an intra-view feature aggregation module for feature extraction and an inter-view cost volume aggregation module to adaptively aggregate cost volumes of different views.
-With sequential process of 2D slices, these methods improve the scalability for high-resolution reconstruction as well as large-scale scenes and reduce memory, however, at the cost of run-time.
+With sequential process of 2D slices, these methods improve the scalability for high-resolution reconstruction as well as large-scale scenes and reduce memory, however, at the cost of run-time [107].
 Coarse-to-Fine: Predicting depth in a coarse-to-fine manner [32, 38, 87, 39, 90, 101, 108] is another solution of reducing both memory consumption and running-time.
 A coarse depth map is first predicted and then upsampled and refined during finer stages to construct fine details.
-CasMVSNet [31] constructs cascade cost volumes by warping features with reduced depth ranges around the previous coarse maps.
+CasMVSNet [32] constructs cascade cost volumes by warping features with reduced depth ranges around the previous coarse depth maps.
 Based on Eq.
-3, the homography of differential warping at stage $k+1$ is
+4, the homography of differential warping at stage $k+1$ is:
 $\mathbf{H}_{i}^{(k+1)}(d^{(k)}+\Delta^{(k+1)})=(d^{(k)}+\Delta^{(k+1)})\mathbf{K}_{0}\mathbf{T}_{0}\mathbf{T}^{-1}_{i}\mathbf{K}^{-1}_{i},$ (9)
@@
 Following MVSNet [9], 3D CNN is used on each stage to regularize the cost volume.
-UCS-Net [38] estimates uncertainty from the coarse prediction to adaptively adjust search ranges in finer stages.
+UCS-Net [38] estimates the uncertainty from coarse prediction to adaptively adjust search ranges in the finer stages.
 CVP-MVSNet [87] constructs the cost volume with a proposed optimal depth resolution of half pixel to narrow depth range in finer stages.
 EPP-MVSNet [109] introduces an epipolar-assembling module to assemble high-resolution information into cost volume and an entropy-based process to adjust depth range.
-TransMVSNet [79] uses a Feature Matching Transformer for robust long-range global context aggregation within and across images.
 WT-MVSNet [92] uses a window-based Epipolar Transformer for enhanced patch-to-patch matching, and a window-based Cost Transformer to better aggregate global information.
@@
 Recently, PatchmatchNet [39] proposes learned adaptive propagation and cost aggregation modules, which enables PatchMatch to converge faster and deliver more accurate depth maps.
-Instead of propagating depth samples naively from a static set of neighbors as done traditionally [18], PatchmatchNet adaptively gathers samples from pixels of the same surface with the guidance of image features.
 PatchMatch-RL [110], jointly estimates depth, normal and visibility with a coarse-to-fine structure.
@@
 DELS-MVS [117] proposes Epipolar Residual Network to search for the corresponding point in the source image directly along the corresponding epipolar line and follows an iterative manner to narrow down the search space.
-Based on IterMVS [40], RIAV-MVS [109] iteratively refines index fields with a 3-level GRUs.
 IGEV-MVS [118] iteratively updates a disparity map regressed from geometry encoding cost volumes and pairs of correlation volumes.
+Wang et al.
+[119] introduce diffusion model in MVS and propose a conditional diffusion network that iteratively denoises the depth map, achieving high accuracy and efficiency.
 3.5 Depth Estimation
 3.5.1 Online MVS
-Many methods [25, 60, 30, 34] apply encoder-decoder on the cost volume $\mathbf{C}$ and reduce the feature channel to 1, i.e., $\mathbf{C}^{\prime}\in\mathbb{R}^{H\times W\times 1}$.
+Many methods [26, 67, 31, 34] apply encoder-decoder on the cost volume $\mathbf{C}$ and reduce the feature channel to 1 as $\mathbf{C}^{\prime}\in\mathbb{R}^{H\times W}$.
 Then the sigmoid activation $\sigma$ is applied for normalization.
-Together with the predefined depth range $[d_{\textrm{min}},d_{\textrm{max}}]$ (mostly manually set), the depth map can be computed.
+Together with the predefined depth range $d_{\textrm{min}},d_{\textrm{max}}]$ (mostly manually set), the depth map can be computed.
 For example, DeepVideoMVS [31] estimates depth as:
@@
 For other methods, Neural-RGBD [97] directly models the cost volume as probability volume $\mathbf{P}$ and adopts soft argmax [69] to predict depth, Eq.
-8.
+11.
 MaGNet [116] takes the mean of Gaussian distribution at last iteration as depth.
@@
 Specifically, the expectation value along the depth direction of probability volume $\mathbf{P}$ is computed as the final prediction:
-$\hat{d}(\mathbf{p})=\sum_{i=1}^{D}d_{i}\cdot\mathbf{P}(i,\mathbf{p}),$ (8)
-where $\mathbf{p}$ is the pixel coordinate, $d_{i}$ is $i$-th depth sample and $P(i,\mathbf{p})$ is the predicted depth probability.
+$\hat{d}(\mathbf{p})=\sum_{i=1}^{D}d_{i}\cdot\mathbf{P}(\mathbf{p},i),$ (11)
+where $\mathbf{p}$ is the pixel coordinate, $d_{i}$ is $i$-th depth sample and $\mathbf{P}(\mathbf{p},i)$ is the predicted depth probability.
 For coarse-to-fine methods [32, 38, 87], soft argmax is applied on each stage to regress the depth maps.
@@
 [120] first use classification to get the optimal hypothesis and then regress the proximity for it.
-DMVSNet [112] estimates two depth maps on multi-stages and composes the final depth map by alternating between selecting the maximum and minimum predicted depth values, which generates an oscillating depth map for improved geometry.
 3.6 Depth Refinement
 Given that the raw depth estimation from MVS may be noisy, refinement is usually used to improve the accuracy.
+DeepMVS [121] apply DenseCRF [122] to the coarse disparity map to encourage the pixels which are spatially close and with similar colors to have closer disparity predictions.
 R-MVSNet [36] enforces multi-view photo-consistency to alleviate the stair effect and achieve sub-pixel precision.
@@
 The cross entropy loss function is defined as:
-$L=\sum_{\mathbf{p}}\left(\sum_{i=1}^{D}-G(i,\mathbf{p})\cdot\log[P(i,\mathbf{p})]\right),$ (11)
-where $G(i,\mathbf{p})$ is the ground truth one-hot vector of depth at pixel $\mathbf{p}$, and $P(i,\mathbf{p})$ is the predicted depth probability.
+$L=\sum_{\mathbf{p}}\left(\sum_{i=1}^{D}-G(i,\mathbf{p})\cdot\log[\mathbf{P}(\mathbf{p},i)]\right),$ (14)
+where $G(\mathbf{p},i)$ is the ground truth one-hot vector of depth at pixel $\mathbf{p}$, and $\mathbf{P}(\mathbf{p},i)$ is the predicted depth probability.
 For methods that predict depth with a hybrid strategy of classification and regression, Wang et al.
@@
 [120] use focal loss [134].
-4 Unsupervised & Semi-supervised Methods with Depth Estimation
-In this section, we introduce unsupervised learning-based MVS methods [125, 126, 42, 43, 41] and semi-supervised method [45].
+4 Unsupervised & Semi-supervised MVS with Depth Representation
+In this section, we introduce unsupervised and semi-supervised MVS with depth representation.
 Supervised MVS methods mentioned in Sec.
 3 depend extensively on the availability of accurate ground truth depth maps obtained through depth-sensing equipment.
-To make MVS practical in more general real-world scenarios, it is vital to consider alternative unsupervised learning-based methods that can provide competitive accuracy compared to the supervised ones without any ground truth depth.
-Existing unsupervised methods are built upon the assumption of photometric consistency (Sec.
+To make MVS practical in more general real-world scenarios, it is vital to consider alternative unsupervised methods that can provide competitive accuracy compared to the supervised ones without any ground truth.
+Existing unsupervised methods are built on the assumption of photometric consistency (Sec.
 4.1), and are categorized into end-to-end (Sec.
@@
 4.3).
-SGT-MVSNet [45] is the only semi-supervised method so far, and we introduce it in Sec.
+To the best of our knowledge, SGT-MVSNet [45] is the only semi-supervised method so far, and we introduce it in Sec.
 4.4.
 4.1 Photometric Consistency Assumption
-In the realm of unsupervised depth map prediction, extant methods [127, 43, 128, 129] endeavor to establish photometric consistency between reference and source perspectives.
-This pivotal notion revolves around the augmentation of similarity between the reference image $\mathbf{I}_{0}$ and individual source images $\mathbf{I}_{i}$ after their warping to align with the reference view.
-In relation to the depth estimation denoted as $\hat{d}$ for the initial image $\mathbf{I}_{0}$, the process involves the projection of reference pixels into the subsequent image $\mathbf{I}_{i}$ using the formulation presented in Eq.
-4.
-Subsequently, the distorted version of the source image denoted as $\hat{\mathbf{I}}_{0}^{i}$ is created by interpolating the RGB values of the source image at the displaced pixel positions through bilinear sampling.
-This interpolation is carried out at the locations where the pixels have been transformed due to the warping process.
+In the realm of unsupervised depth map prediction, extant methods [135, 43, 136, 137] endeavor to establish photometric consistency between reference and source views.
+This pivotal notion revolves around the augmentation of similarity between the reference image $\mathbf{I}_{0}$ and source image $\mathbf{I}_{i}$ after warping to align with the reference view.
+Specifically, for each pixel $\mathbf{p}_{0}$ in $\mathbf{I}_{0}$, we use the estimated depth value $\hat{d}(\mathbf{p}_{0})$ to project $\mathbf{p}_{0}$ into $\mathbf{I}_{i}$ with intrinsic and extrinsic matrices, following Eq.
+5.
+To get the warped source image $\hat{\mathbf{I}}_{0}^{i}$, we interpolate the RGB values from the source image $\mathbf{I}_{i}$ at the coordinates $\mathbf{p}_{i}$ with bilinear interpolation.
 Additionally, alongside the warped image $\hat{\mathbf{I}}_{0}^{i}$, a binary mask $M_{i}$ is commonly generated.
@@
 The photometric consistency loss $L_{\textrm{PC}}$ can be written as:
-$\begin{split}L_{\textrm{PC}}=&\sum_{i=1}^{N-1}\frac{1}{\left\|M_{i}\right\|_{1}}(\left\|\left(\hat{\mathbf{I}}_{0}^{i}-\mathbf{I}_{0}\right)\odot M_{i}\right\|_{2}\\ &+\left\|\left(\nabla\hat{\mathbf{I}}_{0}^{i}-\nabla\mathbf{I}_{0}\right)\odot M_{i}\right\|_{2}),\end{split}$ (12)
+$\begin{split}L_{\textrm{PC}}=&\sum_{i=1}^{N-1}\frac{1}{\left\|M_{i}\right\|_{1}}(\left\|(\hat{\mathbf{I}}_{0}^{i}-\mathbf{I}_{0})\odot M_{i}\right\|_{2}\\ &+\left\|(\nabla\hat{\mathbf{I}}_{0}^{i}-\nabla\mathbf{I}_{0})\odot M_{i}\right\|_{2}),\end{split}$ (15)
 where $\nabla$ denotes the gradient at the pixel level, while $\odot$ symbolizes element-wise Hadamard multiplication.
-In most cases [127, 41, 43, 128, 42], the incorporation of structural similarity loss and depth smoothness loss into the computation is commonplace.
-This practice aims to enhance the training process’s stability and speed up the convergence.
-The computation of the structural similarity loss occurs between a synthesized image and the initial reference image, aiming to uphold contextual congruity.
-More precisely, the assessment of contextual similarity often involves the Structural Similarity Index (SSIM) proposed [130], defined as:
+In most cases [135, 41, 43, 136, 42], the incorporation of structural similarity loss, $L_{\textrm{SSIM}}$, and depth smoothness loss, $L_{\textrm{SM}}$, into the computation is common, which improves the stability of training process and speeds up the convergence.
+Based on the Structural Similarity Index (SSIM) [138], structural similarity loss is computed between a synthesized image and the reference image as:
 $L_{\textrm{SSIM}}=\sum_{i=1}^{N-1}\left[1-\operatorname{SSIM}\left(\mathbf{I}_{0},\hat{\mathbf{I}}_{0}^{i}\right)\right]\odot M_{i},$ (16)
@@
 This continuity is evaluated based on the color intensity gradient present in the input reference image.
-The computation of the smoothness loss, $L_{\textrm{SM}}$, is defined as follows:
+The smoothness loss, $L_{\textrm{SM}}$, is defined as:
 $L_{\textrm{SM}}=\sum_{\mathbf{x}}\left|\nabla_{u}\tilde{d}(\mathbf{x})\right|e^{-\left|\nabla_{u}\mathbf{I}_{0}(\mathbf{x})\right|}+\left|\nabla_{v}\tilde{d}(\mathbf{x})\right|e^{-\left|\nabla_{v}\mathbf{I}_{0}(\mathbf{x})\right|},$ (18)
@@
 They are based on photometric consistency, structural similarity, and smoothness constraints for loss terms.
-Unsup_MVS [127] adopts view synthesis supervision and dynamically selects the $X$ best (lowest loss) values out of $Y$ loss maps.
+Khot et al.
+[135] adopts view synthesis supervision and dynamically selects the $X$ best (lowest loss) values out of $Y$ loss maps.
 However, photometric correspondences can be inaccurate due to non-Lambertian surfaces, camera exposure variations, and occlusions, leading to ambiguous supervision (Sec.
@@
 However, this approach can struggle to converge and lacks detailed information.
-RC-MVSNet [42] enhances supervision through neural rendering by combining Neural Radiance Fields and cost volumes.
-It uses a CasMVSNet backbone to generate depth maps supervised by photometric consistency, along with rendering-consistency and view synthesis losses to handle occlusions and varying lighting.
+RC-MVSNet [42] enhances supervision through neural rendering by combining NeRF [28] and cost volumes.
+Based on a CasMVSNet [32] backone that is supervised by photometric consistency, it uses rendering consistency and view synthesis loss to handle occlusions and varying lighting.
 ElasticMVS [139] addresses issues in photometric loss-based geometry by introducing a part-aware patch-match framework, using an elastic part representation to guide the depth map prediction.
@@
 CL-MVSNet [140] improves proximity between positive pairs through contrastive consistency between a regular CasMVSNet branch and two contrastive branches.
-It also introduces $L0.5$ photometric consistency to improve the precision of the reconstruction.
+It also introduces $L_{0.5}$ photometric consistency loss, which is more robust to outliers because it penalizes large errors less severely than traditional $L_{1}$ or $L_{2}$ norms.
+By growing sub-linearly with the error, it prevents the model’s training from being dominated by outliers with large errors.
+This allows the network to focus on well-matched regions.
 These end-to-end methods train from scratch without pre-processing, reducing training time and complexity in real-world applications.
 4.3 Multi-stage Unsupervised Methods
-Multi-stage methods require pre-training or data pre-processing and are based on pseudo-label generation.
-Self-supervised CVP-MVSNet [128] generates pseudo ground truth depth using photometric consistency, refining depth maps with cross-view consistency checks and point cloud fusion [15].
-The meshes are then reconstructed for further training iterations.
+Multi-stage methods require pre-training or data pre-processing and are typically based on pseudo-label generation.
+Self-supervised CVP-MVSNet [136] generates pseudo ground truth depth using photometric consistency, geometric consistency and point cloud fusion [9].
 U-MVS [43] pre-trains an optical-flow network, PWC-Net [141], and uses flow-depth consistency to generate pseudo labels, reducing supervision ambiguity in foregrounds and backgrounds with an uncertainty-aware self-training consistency.
@@
 A coarse-to-fine depth propagation module improves accuracy at edges and boundaries.
-5 Learning-based MVS without Depth Estimation
-Though the learning-based methods that predict individual depth maps with plane-sweep are the main family of learning-based multi-view stereo, there are many methods of other families that achieve impressive 3D reconstruction quality in the recent years.
-In this section, We discuss four main families: voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
+5 Learning-based MVS with Other Representations
+n this section, we discuss learning-based MVS with the following representations: voxel grids, NeRF, 3D Gaussian Splatting, and direct point-based representations.
+These approaches have demonstrated impressive 3D reconstruction capability in recent years.
 5.1 Voxel-based Methods
-These methods [46, 47, 134, 135] estimate the scene geometry with volumetric representation by leveraging implicit function, e.g., SDF.
-Specifically, Atlas [46] and NeuralRecon [47] attempt to predict the TSDF volume from the 3D feature volume constructed by lifting 2D image features.
-Atlas uses 3D CNN to regress the TSDF volume based on the feature volume accumulated from all images of the scene, which exhibits great completeness of reconstruction.
-The TSDF reconstruction is supervised using $L_{1}$ loss to the ground truth TSDF vaules.
-Since dense feature volume brings lots of computational overhead, NeuralRecon further improves the efficiency by incrementally reconstructing the scene in a fragment-wise and coarse-to-fine manner.
-The 3D features from different fragments are passed through the whole incremental reconstruction process with a RNN.
-TransformerFusion [134] fuses coarse and fine image features in a voxel grid with two transformers and then predicts an occupancy field to represent the scene geometry.
-VoRTX [135] uses a similar design to TransformerFusion and the scene geometry is obtained by passing three different-level features output from transformers through a 3D CNN.
+These models [46, 47, 48, 142] estimate the scene geometry with volumetric representation by leveraging implicit function, e.g., SDF.
+Specifically, these methods predict the TSDF volume from the 3D feature volume constructed by lifting 2D image features and extract the mesh with marching cube [73].
+5.1.1 Direct 3D CNN
+As a seminal work, Atlas [46] employs a 3D CNN with four scale resolution pyramid to regress the TSDF volume from the feature volume accumulated from all images, as shown in Fig.
+4.
+The TSDF reconstructions are supervised using $L_{1}$ loss to the ground truth TSDF vaules.
+However, constructing a dense 3D feature volume is a huge computational overhead for large scenes.
+5.1.2 Coarse-to-fine
+It is a common practice to adopt a coarse-to-fine paradigm [47, 48, 142] for the sake of efficiency.
+One of its core steps is voxel sparsification, which eliminates empty voxels at the coarse level so that memory consumption is reduced in the fine level.
+NeuralRecon [47] incrementally reconstructs the scene in a coarse-to-fine and fragment-wise manner.
+First, in order to reduce memory, NeuralRecon predicts TSDF with a three-level coarse-to-fine approach that gradually increases the density of sparse voxels.
+Second, it reconstructs the TSDF volume of a local fragment and then fuses it with the global volume by GRU [143] Fusion.
+Based on the coarse-to-fine framework of NeuralRecon, TransformerFusion [48] introduces transformer [144] that enables the network to learn to attend to the most relevant image frames for each 3D location in the scene.
+Similarly, VoRTX [142] uses a transformer to fuse image features at each voxel to produce the multi-view feature volume.
+Although state-of-the-art voxel-based methods use coarse-to-fine framework to explicitly improve the efficiency, they are still limited to indoor scenes that are strictly bounded, e.g., ScanNet [27], since the memory storage explicitly increases when the scale of scene increases, e.g., unbounded outdoor scenes.
+Figure 4: Pipeline of Atlas [46]. 2D image features are back-projected into 3D volumes, which are aggregated and passed through a 3D CNN to directly regress a TSDF volume.
 5.2 NeRF-based Methods
@@
 The model is trained by minimizing the loss between the predicted and ground truth color:
-$\mathcal{L}_{\text{color}}=\mathbb{E}[(||\mathbf{C}-\mathbf{C}_{gt}||^{2}].$ (17)
+[…]
+The pipeline of NeRF is shown in Fig.
+5.
 Many subsequent endeavors [145, 146, 147, 148, 149, 150, 151, 152, 153] further improve NeRF in quality, fast training, memory efficiency and real-time rendering.
+Figure 5: Pipeline of NeRF [28].
+Given a 3D position and 2D viewing direction (a), an MLP produces the color and volume density (b).
+Then volume rendering is used to composite these values into an image (c).
+The optimization is minimizing the rendering loss (d).
+Figure 6: Pipeline of 3DGS [29].
+3DGS initializes 3D Gaussian primitives from scene geometry priors (e.g., SfM points) and computes pixel colors by compositing their projected attributes via 3D-to-2D splatting.
+Figure 7: Pipeline of DUSt3R [61].
+Two views of a scene $(I^{1},I^{2})$ are first encoded in a Siamese manner with a shared ViT encoder.
+The resulting token representations $F^{1}$ and $F^{2}$ are then passed to two transformer decoders that constantly exchange information via cross-attention.
+Finally, two regression heads output the two corresponding pointmaps and associated confidence maps.
+Importantly, the two pointmaps are expressed in the same coordinate frame of the first image $I^{1}$.
 Though the initial purpose of NeRF is to perform novel view synthesis, VolSDF [154] and NeuS [49] integrate NeRF with SDF for surface reconstruction.
 The SDF, denoted as $f$, is transformed into the density $\sigma$ for volume rendering.
-Specifically, for a point $\mathbf{p}(t)$, VolSDF computes the volume density $\sigma(\mathbf{p}(t))$ from the signed distance $f(\mathbf{p}(t))$ as:
-$\sigma(\mathbf{p}(t))=\frac{1}{\beta}\Psi_{\beta}(-f(\mathbf{p}(t))),$ (18)
-where $\Psi_{\beta}$ denotes the Cumulative Distribution Function (CDF) of a zero-mean Laplace distribution with learnable scale parameter $\beta>0$. NeuS computes the density in a different way as:
-$\sigma(\mathbf{p}(t))=\max(\frac{-f^{\prime}(\mathbf{p}(t))\Phi^{\prime}_{s}(f(\mathbf{p}(t)))}{\Phi_{s}(f(\mathbf{p}(t)))},0),$ (19)
-where $\Phi_{s}$ is the sigmoid function with learnable scale parameter $s$. After training, the mesh can be extracted from the SDF field with Marching Cubes [67].
-Note that without ground truth geometry supervision, e.g., depth map or TSDF volume, these methods are trained in a self-supervised way as NeRF, Eq.
-17.
-Since training with rendering loss only like NeRF has ambiguity in geometry [145], some following methods improve the reconstruction with explicit geometry supervision, e.g., monocular depth/normal priors [146, 147] and sparse SfM point cloud [50].
-Motivated by Instant-NGP [138] that accelerates training with hash grids, other methods [148, 149, 150, 151] use hash grids to speed up training and improve surface details.
-In addition, to overcome the drawback that NeRF typically needs to be trained for each new scene, some recent methods [152, 153, 154] propose generalizable pipelines for implicit reconstruction even under sparse-view settings.
-Specifically, these methods project 3D points on the image planes and aggregate the corresponding image features as a feature volume like Atlas [46], which is used to estimate the surface location.
+After training, the mesh can be extracted from the SDF field with Marching Cubes [73].
+Since the main focus of this survey is 3D reconstruction, we discuss these NeRF-based methods with SDF representation in the following.
+5.2.1 Optimization-based NeRF
+Without ground truth geometry supervision, e.g., depth map or TSDF volume, training is mainly done in a self-supervised manner with rendering loss, Eq.
+20.
+To enforce SDF function $f$ approximate a valid SDF, existing methods [154, 49] use an eikonal loss [155] for regularization:
+[…]
+where $\mathbf{x}$ denotes the 3D point.
+Motivated by Instant-NGP [147] that accelerates training with hash grids, many methods [156, 50, 157, 51] use hash grids to speed up training and improve surface details.
+However, these methods still mainly use pixel-wise rendering loss for training, which may result in geometric ambiguity, especially in textureless areas [158].
+To alleviate this issue, some methods [159, 160] use patch-wise loss to reduce ambiguity.
+Specifically, based on rendered depth and normal, they warp reference image patches to neighboring views and measure photometric consistency, which is similar to unsupervised depth map-based MVS.
+In addition, other methods introduce explicit geometry supervision, e.g., monocular depth/normal priors [161, 162] and sparse SfM point cloud [160].
+For example, MonoSDF [161] uses a pre-trained Omnidata model [163] to predict monocular depth and normal, which are used to enforce consistency with rendered depth and normal.
+Recently, Ref-NeRF [164], a variant of original NeRF, reparameterizes the appearance with separate diffuse and reflective components by using the reflected view direction, which improves the rendering of specular surfaces.
+Therefore, some follow-up methods [165, 166, 167, 168, 51] adopt this representation in reconstruction and can successfully reconstruct specular surfaces.
 Recall that the depth map-based MVS methods predict depth with photometric consistency across multiple views.
-However, this assumption fails for glossy surfaces with reflections and thus they cannot reconstruct them accurately.
-In contrast, some NeRF methods can handle reflections well.
-Recently, Ref-NeRF [155] reparameterizes the appearance with separate diffuse and reflective components by using the reflected view direction, which improves the rendering of specular surfaces.
-Therefore, recent methods [156, 157, 158, 159, 151] adopt this representation in reconstruction and can successfully reconstruct specular surfaces.
-Specifically, based on VolSDF/NeuS, these methods mainly replace the view direction with the reflected view direction following Ref-NeRF.
+However, this assumption fails for glossy surfaces, and thus these methods cannot reconstruct them accurately.
+Therefore, we believe that NeRF-based MVS is more suitable than depth map-based MVS for reconstructing specular surfaces.
+5.2.2 Generalizable NeRF
+NeRF usually requires a lengthy optimization for each scene and cannot generalize to unseen scenes in a zero-shot manner.
+To address this issue, recent methods [169, 52, 53, 170] propose generalizable NeRF pipelines for both rendering and reconstruction.
+Interestingly, we observe that the majority of these methods are inspired by depth map-based and voxel-based MVS.
+For example, MVSNeRF [169] and C2F2NeUS [171] construct cost volumes with plane sweep to encode global features, following depth map-based offline MVS methods [9, 32].
+Similar to voxel-based MVS [46], some methods [52, 53, 170] use image features to construct feature volumes for geometry estimation.
+More recently, large reconstruction models (LRMs) [54] leverage transformer backbones to directly predict NeRF representations from single or sparse images in a feed-forward manner.
+These models support inputs ranging from single image [54, 172, 173] to posed [174, 175] and un-posed multi-view imagery [176], significantly improving practicality beyond per-scene optimization.
 5.3 3D Gaussian Splatting-based Methods
@@
 During training, 3DGS minimizes the rendering loss like NeRF, as in Eq.
-17.
+20.
 3DGS can be initialized with SfM [178, 17] or MVS [30], which performs better than random initialization [177].
-Motivated by the NeRF-based MVS methods, researchers try to adapt 3DGS for reconstruction.
+The pipeline of 3DGS is shown in Fig.
+6.
+5.3.1 Optimization-based 3DGS
+Motivated by NeRF-based MVS methods, researchers try to adapt 3DGS for reconstruction task.
 Though 3DGS achieves high-quality novel-view synthesis, it is challenging to recover high quality geometry since no explicit geometry constraint is used and 3D Gaussians do not correspond well to the actual surface because of the 3D covariance [55].
-SuGaR [51] introduces a geometry regularization term encouraging the 3D Gaussians to be well-aligned over the scene surfaces so that the Gaussians can contribute to better scene geometry.
+In order to make 3D Gaussians correspond well to the actual surface, SuGaR [55] introduces a geometry regularization term that encourages the 3D Gaussians to be well-aligned over the scene surfaces so that the Gaussians can contribute to better geometry.
 To reconstruct the mesh, SuGaR samples 3D points on a level set of the density computed from 3D Gaussians and then runs Poisson Reconstruction [179] on these points.
-NeuSG [163] jointly trains Neuralangelo [149], a NeRF-based method, and 3DGS.
-During training, the normals from the SDF field of Neuralangelo regularize the rotation of 3D Gaussians, while the SDF field is regularized to ensure that the SDF values at 3D Gaussians’ positions are close to zero.
-2DGS [52] further improves surface accuracy by modeling the 3D primitives as surfels ( tangent plane clipped) and fuses rendered depth maps into mesh with TSDF fusion.
-PGSR [53] elicits unbiased depth maps as strong geometry prior from rendered distances and rendered normals, and enforces multi-view geometric consistency by regularizing such a prior with local plane assumption and edge constraint.
-RaDe-GS [164] also improves the quality of geometry by precisely modeling the depth and normal, instead of the error-prone approximate projection in the original 3DGS.
-However, retrieving good geometry from Gaussian-based representations still needs to be explored.
-5.4 Large Feed-forward Methods
-In 3D reconstruction and 3D generation, a recent prominent trend is to directly learn the 3D representation from large-scale 3D datasets, e.g.Objaverse [29].
-These methods typically adopt large-scale transformers [165] to generate a 3D representation based on various inputs, including single image [54, 166, 167], textual description [168, 169], posed multi-view images [168, 170, 171, 172, 173], and un-posed multi-view images [174, 55].
-LRM [54] uses a transformer-based feed-forward network to regress the 3D object from a single image.
-Specifically, it predicts features of a tri-plane representation [175], which is subsequently converted into a radiance field similar to NeRF.
-Large feed-forward models can also be combined with 3D Gaussian Splatting, where they directly generate 3D Gaussian points from posed multi-view inputs [168, 170, 171, 172, 173, 176].
-Instead of predicting 3D Gaussian locations directly, which is a highly ill-posed task, the common practice is to use per-pixel aligned Plücker ray embeddings and predict the depth for each pixel [176] of input views.
-As such, this design requires the input views to cover a large range of the object.
-Recently, DUSt3R [55] predicts pixel-aligned point coordinates directly from uncalibrated image pairs.
-After extracting image token representations with Vision Transformer [177], two transformer decoders exchange information of the token representations via cross-attention and then regress the per-pixel 3D points in the coordinate frame of the first image.
-Many downstream tasks can be performed with the point clouds, e.g., point matching, localization, intrinsic and relative pose estimation.
-Directly extended from DUSt3R, MASt3R [56] focuses on the specific application of feature description and matching across multi-view images.
-It leverages an additional matching head on top of DUSt3R’s architecture to boost the performance of feature-based matching.
-With the dense correspondences established, MASt3R manages to achieve a comparable reconstruction quality as MVSNet [15].
-Spann3R [178] manages an external spatial memory to keep track of all previous relevant 3D information yielded by DUSt3R and then queries this spatial memory to predict the 3D structure.
+Instead of performing regularization on 3D Gaussian, 2DGS [56] introduces a 2D Gaussian representation by modeling the 3D primitives as surfels.
+Similar to NeRF, relying solely on image rendering loss can easily fall into local minima, leading to Gaussian shapes inconsistent with the actual surface.
+Introducing explicit geometry constraints can alleviate this problem.
+For example, PGSR [57] enforces the local consistency of depth and normal in a local image patch, and multi-view geometric and photometric consistency as COLMAP [17].
+5.3.2 Generalizable 3DGS
+Following the paradigm of feed-forward NeRF, PixelSplat [180] predicts 3DGS representations from paired images by leveraging epipolar constraints to resolve scale ambiguity.
+MVSplat [58] further incorporates depth cues for 3DGS prediction by constructing a cost volume from multiple input images.
+DepthSplat [59] extends this idea by achieving more robust depth estimation through the combination of cost volume and monocular depth features [181].
+Instead of focusing on improving geometry and rendering quality, recent works such as SplatterImage [182], Flash3D [183], GS-LRM [184],LGM [60], and Bolt3D [185] explore the generalization capability of single-image-to-3DGS generation without per-scene optimization.
+These methods leverage multi-view diffusion models or transformers pretrained on large-scale multi-view datasets to directly predict per-pixel 3D Gaussian attributes from images.
+5.4 Large Feed-forward Point-based Methods
+In 3D reconstruction, a prominent trend is to directly learn 3D point geometry from large-scale multi-view data with transformer backbones [144], shifting the paradigm from per-scene optimization toward one-shot inference of geometry or renderable content.
+Unlike traditional MVS that relies on hand-crafted cost volumes, these methods shift the paradigm toward one-shot inference of dense point maps.
+Recent feed-forward point-based methods compress the classical MVS pipeline into a single forward pass: (1) feature extraction; (2) cross-view aggregation via attention in lieu of an explicit plane-sweep cost volume; and (3) prediction heads that jointly regress cameras and dense geometry (depth/point maps).
+We illustrate the pipeline of DUSt3R [61], a pioneering feed-forward method, in Fig.
+7.
+Below we describe how modern feed-forward methods instantiate each stage in a single forward pass.
+ViT feature extraction.
+A pre-trained transformer encoder extracts view-wise, geometry-aware features/tokens from few to many images, supporting downstream camera, depth, and pointmap prediction [62, 61, 186, 63].
+In practice, DINO models are often chosen as the ViT backbone because its self-supervised features transfer robustly across tasks and domains and exhibit emergent semantics useful for dense correspondences [187].
+Cross-view aggregation via attention.
+Instead of building explicit plane-sweep cost volumes, global or alternating attention patterns interleave per-frame and multi-frame aggregation, allowing long-range correspondence reasoning while remaining feed-forward [62, 61].
+The attention mechanism functions as a learnable and unconstrained feature matching in multi-view feed-forward 3D reconstruction.
+With no pose prior and unknown epipolar lines, the theoretical area to be searched is actually the whole image and the Transformer modules learn implicit correspondences in a data-driven manner, as is revealed in [188].
+While the more classical cost volumes are built with known poses and epipolar lines, where the area to be searched become a 1D segment.
+Prediction heads for geometry and pose.
+Rather than deriving geometry purely from pairwise matches, modern heads directly regress intrinsics/extrinsics together with dense depth or point maps.
+Early work showed that regressing point maps is effective even from two views and unifies monocular/binocular settings [61]; recent multi-view extensions output cameras, depth, point maps, and tracks in one pass [62].
+Emerging designs relax reference-view assumptions via permutation-equivariant architectures and scale to long sequences with spatial memory or causal/streaming transformers [189, 190, 191].
+For future work, our analysis points to two avenues for advancing feed-forward 3D reconstruction.
+First, to address the fundamental out-of-distribution problem, future research may focus on developing memory-efficient training paradigms that allow large-scale, many-view scenes to be incorporated directly during training.
+By aligning the training distribution more closely with real-world test scenarios, such approaches could substantially improve generalization to complex environments.
+Second, to overcome the accuracy limitations of single-pass prediction, a promising direction is to create hybrid architectures that merge feed-forward speed with iterative refinement.
 6 Discussions
-In this section, we summarize and discuss the performance of learning-based MVS methods, including supervised online methods with depth estimation (Sec.
-6.1), supervised offline methods with depth estimation (Sec.
-6.2), unsupervised methods with depth estimation (Sec.
-6.3) and methods without depth estimation (Sec.
-6.4).
-Moreover, we discuss potential directions for future research (Sec.
-6.5).
-TABLE I: Quantitative results of supervised online MVS methods on ScanNet [26] and 7-Scenes [179].
+In this section, we summarize and discuss the performance and efficiency of learning-based MVS methods. Moreover, we discuss potential directions for future research.
+TABLE II: Quantitative results of point cloud evaluation on Tanks & Temples [23].
+F-score is listed (the higher, the better).
+MapAnything [63] reconstructs point clouds directly from unposed images.
+All images are used for MapAnything, except for Courthouse (50% subset due to memory limitations).
+Representation Methods Barn Caterpillar Courthouse Ignatius Meetingroom Truck Mean
+Depth map DeepVideoMVS [31] - - - 0.01 - 0.01 -
+CasMVSNet [32] 0.52 0.42 0.36 0.71 0.35 0.59 0.49
+$D^{2}$HC-RMVSNet [37] 0.64 0.57 0.42 0.74 0.33 0.70 0.57
+AA-RMVSNet [89] 0.61 0.58 0.40 0.77 0.36 0.71 0.57
+IterMVS [40] 0.62 0.49 0.38 0.78 0.38 0.68 0.55
+TransMVSNet [91] 0.57 0.47 0.38 0.73 0.42 0.63 0.53
+MVSFormer++ [93] 0.66 0.59 0.42 0.78 0.40 0.69 0.59
+CasDiffMVS [119] 0.69 0.54 0.42 0.81 0.45 0.70 0.60
+Voxel NeuralRecon [47] - 0.01 - 0.07 - - -
+VisFusion [192] - 0.09 - 0.09 0.03 0.15 -
+NeRF NeuS [49] 0.29 0.29 0.17 0.83 0.24 0.45 0.38
+Geo-NeuS [160] 0.33 0.26 0.12 0.72 0.20 0.45 0.35
+Neuralangelo [50] 0.70 0.36 0.28 0.89 0.32 0.48 0.50
+3DGS SuGaR [55] 0.14 0.16 0.08 0.33 0.15 0.26 0.19
+2DGS [56] 0.36 0.23 0.13 0.44 0.16 0.26 0.30
+GOF [193] 0.51 0.41 0.28 0.68 0.28 0.58 0.46
+PGSR [57] 0.66 0.41 0.21 0.80 0.29 0.60 0.50
+Feed-forward Point-based MapAnything [63] 0.14 0.34 0.10 0.18 0.12 0.23 0.19
+6.1 Performance Comparison
+To provide a thorough comparison across different methods, we evaluate state-of-the-art methods on Tanks & Temples [23] and summarize the results in Tab. II.
+For depth map-based methods, we find that online methods [31] perform worse than offline methods [32, 91, 119, 93].
+Online methods typically sample sparse depth hypotheses from a pre-defined depth range to construct a lightweight cost volume, which limits the accuracy in large-scale scenes.
+In addition, online methods are limited by causal/temporal processing and cannot exploit geometric constraints from future frames as offline methods.
+For voxel-based methods, we observe that they fail on both outdoor and indoor scenes.
+From the representation’s perspective, we can attribute the performance gap to the following two aspects.
+(a) Voxel-based methods usually specify the spatial location of the potential voxels given the camera trajectory, with a metric voxel resolution.
+The inherent memory complexity and limited resolution of voxel grids are fundamental bottlenecks for large-scale, unbounded scenes.
+(b) Unlike depth-based representations, which estimate geometry primarily based on photometric consistency on the 2D image plane and allow for sharp discontinuities (suitable for complex occlusions in outdoor scenes), voxel-based SDFs impose a global 3D continuity constraint.
+While this constraint helps regularize geometry in texture-less indoor regions, it acts as a limiting factor in large-scale scenes by enforcing smoothness where sharp edges or fine details are expected.
+For NeRF/3DGS-based methods, we observe that they achieve high accuracy in small scenes, e.g., Ignatius, producing high-quality reconstruction with fine-grained details.
+However, the performance is notably worse than offline methods on large scenes, e.g., Courthouse.
+Scaling NeRF/3DGS to large-scale scenes is still an open topic in research.
+For large feed-forward point-based methods, although the reconstructed point clouds appear reasonable, their quantitative accuracy is typically low.
+This is due to several challenges: (1) they must estimate camera poses, significantly increasing task complexity; however, models like DUSt3R skip explicit camera projection, and VGGT predicts camera tokens in parallel, bypassing camera geometry during the forward pass.
+This flexibility in handling unposed data, while advantageous, means that metric evaluation is highly sensitive to pose alignment and can limit accuracy.
+(2) These models are typically not trained on scenes with hundreds of views, creating an out-of-distribution problem for larger datasets; and (3) huge memory consumption compels the use of low-resolution images or fewer input frames, directly limiting the fidelity of the final reconstruction.
+6.2 Efficiency Comparison
+Low run-time and memory consumption are crucial in most industrial applications with limited computational power and storage, e.g., robotics and AR/VR.
+Compared to other representations, depth-map based MVS is usually more suitable for these applications.
+Volume-based methods are difficult to scale to large scenes due to substantial memory storage.
+NeRF-based and 3D Gaussian Splatting-based methods typically require per-scene optimization for high-quality reconstruction.
+Large feed-forward methods have huge backbones and thus their GPU memory consumption is high.
+We compare the efficiency of state-of-the-art depth-map based MVS methods [32, 39, 31, 112, 40, 108, 91, 120, 88, 194] in Fig.
+8.
+Although online methods [31] achieve high efficiency in run-time with low-resolution images, we observe that their run-time and GPU memory consumption are higher than many offline methods [39, 112, 40, 108, 119] with high-resolution images.
+To improve efficiency, offline methods [39, 112, 40, 108, 119] carefully simplify the architecture, e.g., replacing costly 3D convolution with 2D convolution.
+Moreover, these methods, especially CasDiffMVS [119], achieve competitive performance when compared to the top-performing methods [194, 93].
+Therefore, it is promising to further improve both the performance and efficiency for practical applications.
+Figure 8: GPU memory consumption, run-time, and reconstruction error of state-of-the-art depth map-based methods [31, 39, 112, 40, 108, 91, 120, 194] on DTU [22].
+To fairly compare efficiency, we fix the image resolution and the number of input images, which are set to $1600\times 1152$ and 5 respectively.
+All experiments are done on the same workstation with a NVIDIA 2080 Ti GPU.
+6.3 Future Research Directions
+6.3.1 Datasets and Benchmarks
+For MVS methods, ScanNet [27] and DTU [22] are two main training datasets, while ScanNet [27], DTU [22], Tanks & Temples [23] and ETH3D [24] are the main evaluation benchmarks.
+For training, the scene scale of ScanNet and DTU is relatively small (room-scale for ScanNet and object-scale for DTU) and their quality is not satisfactory [195].
+Recently, many researchers have tried to improve the dataset quality.
+For example, TartanAir dataset [196] is a large-scale synthetic dataset that is collected in photo-realistic simulation environments.
+BlendedMVS [25] introduces more large-scale scenes and improves the performance of many MVS methods on real-world scenes.
+MVImgNet [197] and Objaverse [198] are large datasets of multi-view images collected by shooting videos of real-world objects or rendering 3D object models.
+However, both of them focus on object-level scenes, whereas large-scale scenes are essential for many applications.
+DL3DV [199] attempts to address this issue by capturing videos of large-scale real-world scenes.
+However, it lacks the ground truth geometry required for MVS training.
+Therefore, improving the scalability and quality of datasets is an important research direction.
+One direction is to use advanced sensors, e.g., depth camera and LiDAR, to capture high-quality data in diverse real-world scenario [78, 200].
+The other direction is to exploit large-scale video data, e.g., Youtube videos, and extract high-quality pseudo ground-truth with state-of-the-art models.
+For evaluation, the main benchmarks [27, 22, 23, 24] are all introduced before 2018.
+Some of these benchmarks are currently saturated and the methods are difficult to further improve the performance on them.
+Therefore, it is meaningful to introduce new benchmarks to evaluate the robustness and performance of learning-based methods in challenging real-world scenes.
+For example, the new benchmarks are expected to include challenging settings, such as dynamics, low-texture, reflections, and appearance changes [201].
+6.3.2 View Selection
+As discussed in Sec.
+2.2, view selection is crucial for triangulation quality.
+Picking neighboring views that are suitable for triangulation can not only improve the reconstruction accuracy but also reduce useless computation for the bad views, e.g., views with strong occlusions.
+However, view selection is often overlooked and not well studied.
+For example, all offline depth map-based MVS methods [32, 39, 91, 108] follow MVSNet [9] and use the same simple heuristic strategy to compute scores for neighboring views and sort them.
+Online depth map-based MVS methods [67, 31, 34] and voxel-based methods [47] adopt heuristic strategies to choose views with enough pose-distance.
+Though it is intractable to incorporate non-differentiable view selection into deep learning, it is worth exploring new view selection strategies since it may improve the reconstruction without changing the model design.
+Current view selection strategies simply pick the same set of neighboring views for all the pixels in the reference image, and it is probable that different reference pixels have different optimal choices of neighboring views.
+The learned pixel-wise view weight [101, 39, 40] addresses this issue by assigning different weights to source views for each pixel based on their visibility across neighboring views.
+For large feed-forward methods such as VGGT [62], cross-view attention provides an alternative mechanism for adaptive view selection with the attention weights.
+6.3.3 Feature Extraction
+Currently, most of the depth map-based [9, 31] and voxel-based [47] methods mainly use 2D CNN and FPN [83] structure to learn multi-scale features.
+To flexibly improve the receptive fields, Deformation Convolution [202] and attention mechanism [144] are applied  [98, 90, 89, 39, 91].
+However, they still cannot generalize well in challenging regions, such as reflective and low-textured areas.
+Motivated by the success of vision transformer [203], recent works [204, 93, 61, 62] introduce pre-trained vision transformers, e.g., ViT [203] and DINO [94, 187, 205], to capture global information, alleviating the issues of reflections and texture-less regions.
+However, vision transformers are usually trained and tested on low-resolution images due to their complexity.
+To accurately reconstruct the scene with high-resolution inputs, it is important to utilize fine-grained information for precise feature matching across views, which can be nicely captured with CNN.
+Therefore, an interesting direction is to combine vision transformer and CNN in a coarse-to-fine framework, where vision transformer is used in low-resolution and CNN is used in high-resolution for local refinement.
+6.3.4 Generative Reconstruction
+MVS tends to be sensitive to occlusions and sparse viewpoints because of their strict geometric requirements.
+However, these challenging conditions are quite common in practice, such as autonomous driving.
+To this end, while mainstream MVS methods follow a regression-based paradigm, it would be promising to leverage the inpainting power of generative models to produce reasonable reconstruction in unseen regions.
+For example, ReconFusion [206] explores to distill the generated novel views to a NeRF representation for sparse-view reconstruction.
+Difix3D+ [207] refines noisy 2D renderings from Gaussian Splatting with reference views and Stable Diffusion [208].
+Achieving visually pleasing 3D reconstruction while maintaining the reconstruction fidelity and geometric consistency with generative models is a significant research trend, which will also power the development of world models.
+6.3.5 Efficiency
+Though reconstruction accuracy is considered the most important factor in MVS, it is also critical to have models that can run in near real-time and with low memory, e.g., AR/VR and autonomous driving.
+Comparatively, depth map-based methods are most suitable for achieving efficiency because of their conciseness, flexibility, and scalability.
+Though online MVS methods can achieve real-time estimation with images of low resolutions, they may have issues with memory consumption since many large backbones are usually used for feature extraction [31, 34].
+In addition, the efficiency in both run-time and memory will drop when image resolution increases, as shown in Fig.
+8.
+For offline MVS methods, as discussed in Sec.
+6.2, some recent methods try to improve efficiency with carefully simplified network architectures.
+However, this usually limits the performance when compared to state-of-the-art methods.
+In addition to designing lightweight modules [119], using model compression techniques and knowledge distillation are potential directions to not only keep the good performance but also improve efficiency.
+Large feed-forward models heavily rely on attention mechanism to capture intra-frame and inter-frame relations.
+However, the high computational cost restricts the use of high-resolution or numerous input images, which can degrade performance [62, 63].
+To reduce computational cost of attention mechanism, one direction is to introduce token merging to reduce attention redundancy [209].
+6.3.6 Prior Assistance
+MVS mainly relies on assessing local photometric consistency to find the optimal matching across images.
+Accordingly, it usually encounters difficulties when estimating the geometry for regions where the photometric measurement becomes unreliable or invalid, e.g., textureless areas and non-Lambertian surfaces, which are common in human-made scenes [24].
+Therefore, using prior information to guide the MVS algorithm in these challenging regions is a promising research direction.
+We elaborate 4 typical examples of prior assistance as follows.
+Monocular Depth: Recently, significant progress has been made in monocular depth estimation [210, 181], with existing methods achieving accurate results on diverse in-the-wild datasets.
+Leveraging monocular depth as a strong prior can enhance robustness in challenging scenarios such as low-textured regions and reflective surfaces [161, 59].
+For example, monocular depth can be used to regularize or inpaint depth estimation in these challenging regions.
+Surface Normal: As a non-local representation of the geometry compared with the depth map, surface normal has been proven effective in improving reconstruction performance in recent works [211, 212].
+Similar to monocular depth [210, 181], monocular normal estimation [163] produces high-quality normal priors on in-the-wild data, which can improve reconstruction quality in low-textured regions[161, 213].
+Shape Prior: For indoor scenes where common textureless areas, e.g., walls, planes are suitable choices of the geometric primitives and are exploited in traditional methods [20].
+Recently, PlaneMVS [214] and PlanarRecon [215] explicitly estimate plane parameters for depth refinement or holistic reconstruction.
+In addition, predicting object-level attributes simultaneously when estimating depth [216] is also a feasible solution for specific applications like urban modeling.
+Semantic Segmentation: Intuitively, points assigned with the same semantic labels may be more likely to lie on the same 3D plane.
+Some attempts [217, 218] employ a rule-based protocol to utilize semantic segmentation for better matching and depth quality.
+Manhattan-SDF [219] relies on 2D segmentation results to enforce Manhattan world priors to handle low-textured planar regions.
+Similarly, Shvets et al.
+[220] also employ SAM (Segment Anything Model) encoder to generate semantic features to enhance MVS.
+7 Conclusion
+In this survey, we have provided a comprehensive review of the learning-based MVS methods, which are categorized into: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
+Particularly, we have devoted significant attention to depth map-based methods because of their conciseness, flexibility and scalability.
+We explain key aspects such as datasets utilized, general working pipelines, and algorithmic intricacies.
+Furthermore, we have summarized and compared the quantitative performance of these methods across popular benchmarks, providing valuable insights into their efficacy and applicability.
+Finally, our discourse extends to an in-depth exploration of potential research directions for the future of MVS, highlighting avenues for further investigation and innovation in the field.
+Supplementary Material
+1 Preliminaries
+1.1 Plane Sweep
+We take a toy example of three cameras viewing the same object as illustrated in Fig.
+9.
+Plane sweep is then performed on the frustum of cam_x (termed as the reference camera) by creating a series of fronto-parallel hypothesized planes within a given range (typically $d_{\textrm{min}},d_{\textrm{max}}]$), with each plane corresponding to a depth value w.r.t.
+cam_x.
+Let’s then examine two 3D points, $M$ and $M^{\prime}$, as examples of occupied and unoccupied positions on hypothesized planes, respectively.
+For $M$, all of the three cameras capture the identical point lying on the geometry surface with photometric consistency.
+In contrast, the photometric consistency of the observations of $M^{\prime}$ are poor, indicating that $M^{\prime}$ is an invalid hypothesis.
+To evaluate the similarity of the observations towards $M$ and $M^{\prime}$, the images of cam_1 and cam_2 (termed as the source cameras) are warped to each plane by homography.
+These warped images are then compared against the scaled images of cam_x.
+Depth hypotheses with high similarity measures are considered reliable.
+Figure 9: Illustration of plane sweep algorithm [33].
+To estimate the depth map for reference image (cam_x), neighboring source images (cam_1, cam_2) are projected with homography to fronto-parallel planes of the reference view frustum.
+Regarding selecting $D$ depth hypotheses from the triangulated depth range $d_{\textrm{min}},d_{\textrm{max}}]$, there are two main schemes, namely the forward depth sampling and the inverse depth sampling.
+The naive forward sampling divides the depth range into D1D-1 depth intervals with identical lengths.
+Given a depth index $k$, we have
+$d_{k}=d_{\textrm{min}}+\frac{k}{D-1}(d_{\textrm{max}}-d_{\textrm{min}}),k=0,\ldots,D-1,$ (25)
+where the depth hypotheses distribute uniformly between the two ends. While for the inverse sampling scheme [35], we sample uniformly in the multiplicative inverse of $d$, such that
+$\frac{1}{d_{k}}=\frac{1}{d_{\textrm{max}}}+\frac{k}{D-1}(\frac{1}{d_{\textrm{min}}}-\frac{1}{d_{\textrm{max}}}),k=0,\ldots,D-1.$ (26)
+In this way, the distribution of sampled depth values becomes sparser with $d$ approaching to $d_{\textrm{max}}$.
+When reconstructing unbounded outdoor scenes, where the depth range is rather large, the inverse sampling will be a reasonable choice since it samples more densely at the foreground.
+In addition, as revealed in [101], the inverse sampling leads to a uniform sampling on the projected epipolar lines of source images.
+1.2 Depth Fusion
+To remove outliers and improve reconstruction quality, offline MVS methods filter out outliers before converting the depth maps to point clouds.
+Depth filtering mainly involves photometric and geometric filtering [9].
+For photometric consistency filtering, a per-pixel confidence is estimated to measure the confidence of depth estimation, i.e., probability that the ground truth depth is within a small range near the estimation.
+Then a threshold can be set to filter depth values with low confidence.
+To estimate this confidence, existing methods either follow MVSNet [9] and estimate photometric confidence from the probability volume [36, 37, 89, 9, 32, 39] or use a network to predict the confidence of estimation [40, 119].
+For geometric consistency filtering, the consistency of depth estimations are measured among multiple views.
+For a pixel $\mathbf{p}$ in the reference view 0, we project it to pixel $\mathbf{p}^{\prime}$ in its $i$-th neighboring view through its depth prediction $d_{0}(\mathbf{p})$.
+After looking up the depth for $\mathbf{p}^{\prime}$, $d_{i}(\mathbf{p}^{\prime})$, we re-project $\mathbf{p}^{\prime}$ back to the reference view at pixel $\mathbf{p}^{\prime\prime}$ and look up its depth $D_{0}(\mathbf{p}^{\prime\prime})$.
+We consider pixel $\mathbf{p}$ and its depth as consistent to the $i$-th neighboring view, if the distances, in image space and depth, between the original estimate and its re-projection satisfy:
+$\xi_{d}=\|\mathbf{p}-\mathbf{p}^{\prime\prime}\|_{2}\leq\tau_{1},$ (27)
+$\xi_{p}=\frac{\|d_{0}(\mathbf{p}^{\prime\prime})-d_{0}(\mathbf{p})\|}{d_{0}(\mathbf{p})}\leq\tau_{2},$ (28)
+where $\tau_{1}$ and $\tau_{2}$ are thresholds.
+The pixels are considered to be reliable estimations if they are consistent in at least $N$ neighboring views.
+Recently, instead of using predefined thresholds for $\tau_{1}$ and $\tau_{2}$, dynamic consistency checking [37] is proposed to dynamically aggregate geometric matching error among all views and improve the robustness of reconstruction.
+Specifically, the dynamic multi-view geometric consistency $C_{\textrm{geo}}(\mathbf{p})$ is computed as:
+$C_{\textrm{geo}}(\mathbf{p})=\sum_{i}\exp{(-(\xi_{p}+\lambda\xi_{d}))},$ (29)
+where $\lambda$ is a weight to balance the reprojection error in two metrics. Then the outliers with $C_{\textrm{geo}}(\mathbf{p})$ smaller than a threshold are filtered.
+1.3 Evaluation Metrics
+1.3.1 2D Metrics
+We list the commonly used metrics as follows: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel) and inlier ratio with threshold 1.25 ($\delta<1.25$):
+$\text{Abs}=$ $\frac{1}{N}\sum_{\mathbf{p}}|d(\mathbf{p})-\hat{d}(\mathbf{p})|,$ (30)
+$\text{Abs Rel}=$ $\frac{1}{N}\sum_{\mathbf{p}}\frac{|d(\mathbf{p})-\hat{d}(\mathbf{p})|}{d(\mathbf{p})},$
+$\text{Inlier Ratio}=$ $\frac{1}{N}\sum_{\mathbf{p}}\mathbbm{1}\left[\frac{d(\mathbf{p})}{1.25}<\hat{d}(\mathbf{p})<1.25d(\mathbf{p})\right],$
+where $d(\mathbf{p})$ denotes the ground truth depth, $\hat{d}(\mathbf{p})$ denotes the estimation, $N$ denotes the number of pixels with valid depth measurements and $\mathbbm{1}[\cdot]$ denotes the indicator function.
+1.3.2 3D Metrics
+Precision/Accuracy: Precision/Accuracy measures the percentage of predicted points that can be matched to the ground truth point cloud.
+Considering a point $\mathbf{P}_{p}$ in the predicted point cloud, it is considered to have a good match in the ground truth point cloud $\{\mathbf{P}_{g}\}$ if
+$\|\mathbf{P}_{p}-\mathop{\arg\min}\limits_{\mathbf{P}\in\{\mathbf{P}_{g}\}}\|\mathbf{P}-\mathbf{P}_{p}\|_{2}\|_{2}\leq\tau,$ (31)
+where $\tau$ is a scene-dependent threshold assigned by datasets, usually set to a large value for large-scale scenes.
+Note that in some datasets, instead of being measured by percentage [23, 24], precision/accuracy is measured by mean or median absolute distance [22].
+Recall/Completeness: Recall/Completeness measures the percentage of ground truth points that can be matched to the predicted point cloud.
+For a point $\mathbf{P}_{g}$ in the ground truth point cloud, it is considered a good match in the predicted point cloud $\{\mathbf{P}_{p}\}$ if
+$\|\mathbf{P}_{g}-\mathop{\arg\min}\limits_{\mathbf{P}\in\{\mathbf{P}_{p}\}}\|\mathbf{P}-\mathbf{P}_{g}\|_{2}\|_{2}\leq\tau.$ (32)
+Similar to precision/accuracy, recall/completeness can be measured by the percentage of points [23, 24] or measured by the absolute distance [22], similar to Chamfer distance.
+F-Score: The two aforementioned metrics measure the accuracy and completeness of predicted point clouds.
+However, each of these metrics alone cannot present the overall performance since different MVS methods adopt different assumptions.
+A stronger assumption usually leads to higher accuracy but lower completeness.
+If only precision/accuracy is reported, it would favor MVS algorithms that only include estimated points of high certainty.
+On the other hand, if only recall/completeness is reported it would favor MVS algorithms that include everything, regardless of accuracy.
+Therefore, F-score, an integrated metric, is introduced [23, 24].
+F-score is the harmonic mean of precision and recall.
+It is sensitive to extremely small values and tends to get more affected by smaller values, which means that F-score does not encourage imbalanced results.
+However, in most cases, F-score still suffers from unfairness due to the limitations of ground truth, e.g., sparse and incomplete point clouds may penalize for filling in the areas that are not present in the ground truth [46].
+2 Supervised Method with Depth Representation
+2.1 Cost Volume Regularization
+In Fig. 10, we visualize the three main categories to perform cost volume regularization for offline MVS methods: direct 3D CNN, coarse-to-fine and RNN.
+Figure 10: Illustration of typical cost regularization schemes for offline MVS methods that use 4D cost volumes.
+(a) direct 3D CNN regularization [9, 104, 105, 35] applies a 3D CNN to aggregate contextual information; (b) coarse-to-fine regularization [32, 38, 87, 101] constructs multi-scale cost volumes based on coarse prediction and uses 3D CNN regularization on each scale; (c) RNN regularization [36, 37, 89] sequentially regularizes 2D slices of the cost volume to reduce memory consumption.
+TABLE III: Quantitative results of voxel-based MVS methods on ScanNet [27].
+Methods Prec. $\uparrow$ Recall $\uparrow$ F-score $\uparrow$
+Atlas [46] 0.565 0.598 0.578
+NeuralRecon [47] 0.684 0.479 0.562
+TransformerFusion [48] 0.648 0.547 0.591
+VoRTX [142] 0.688 0.607 0.644
+VisFusion [192] 0.695 0.527 0.598
+3 Benchmark Performance
+In this section, we summarize the performance of different methods on the common benchmarks they use.
+Specifically, we summarize the quantitative results of online depth map-based MVS in Tab.
+IV, offline depth map-based MVS in Tab.
+V, voxel-based MVS in Tab.
+III, and NeRF/3D Gaussian Splatting-based MVS in Tab.
+VI.
+TABLE IV: Quantitative results of online MVS methods on ScanNet [27] and 7-Scenes [76].
 Methods ScanNet [27] 7-Scenes [76]
-abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$uparrow abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$uparrow
+abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$ $\uparrow$ abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$ $\uparrow$
 MVDepthNet [26] 0.167 0.087 0.925 0.201 0.117 0.877
-DPSNet [96] 0.219 0.119 0.868 0.249 0.149 0.826
 Neural-RGBD [97] 0.236 0.122 0.850 0.214 0.131 0.865
@@
 DoubleTake [85] 0.077 0.037 0.984 0.099 0.053 0.970
-TABLE II: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for unsupervised depth map-based MVS methods.“Finetuned with BlendedMVS” denotes whether the methods are finetuned on BlendedMVS [24] before evaluating on Tanks and Temples [22].
-Methods DTU [21] Finetuned with Tanks and Temples [22]
-Acc. $\downarrow$ Comp. $\downarrow$ Overall $\downarrow$ BlendedMVS [24] Intermediate $F_{1}$uparrow Advanced $F_{1}$uparrow
-End-to-end Unsup_MVSNet [127] 0.881 1.073 0.977 - -
-MVS2 [125] 0.760 0.515 0.637 37.21 -
-M3VSNet [126] 0.636 0.531 0.583 37.67 -
-JDACS-MS [41] 0.398 0.318 0.358 45.48 -
-RC-MVSNet[42] 0.396 0.295 0.345 55.08 30.82
-ElasticMVS[131] 0.374 0.325 0.349 57.88 37.81
-CL-MVSNet[132] 0.375 0.283 0.329 59.39 37.03
-Multi-stage Self-supervised CVP-MVSNet [128] 0.308 0.418 0.363 43.48 -
-U-MVS [43] 0.354 0.354 0.354 57.15 30.97
-KD-MVS [44] 0.359 0.295 0.327 64.14 37.96
-TABLE III: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for supervised offline MVS methods.
+TABLE V: Quantitative results of point cloud evaluation on MVS benchmarks [22, 23, 24] for offline MVS methods.
 “Finetuned with BlendedMVS” denotes whether methods are finetuned on BlendedMVS dataset [25] before evaluation on Tanks and Temples [23] and ETH3D [24].
 Methods DTU [22] Finetuned with Tanks and Temples [23] ETH3D [24]
-Acc. $\downarrow$ Comp. $\downarrow$ Overall $\downarrow$ BlendedMVS [24] Intermediate $F_{1}$uparrow Advanced $F_{1}$uparrow Training $F_{1}$uparrow Test $F_{1}$uparrow
+Acc. $\downarrow$ Comp. $\downarrow$ Overall $\downarrow$ BlendedMVS [25] Intermediate $F_{1}$ $\uparrow$ Advanced $F_{1}$ $\uparrow$ Training $F_{1}$ $\uparrow$ Test $F_{1}$ $\uparrow$
 Direct 3D CNN MVSNet [9] 0.396 0.527 0.462 43.48 - - -
@@
 RNN R-MVSNet [36] 0.383 0.452 0.417 48.40 24.91 - -
-$D^{2}$-RMVSNet [37] 0.395 0.378 0.386 59.20 - - -
+$D^{2}$HC-RMVSNet [37] 0.395 0.378 0.386 59.20 - - -
 AA-RMVSNet [89] 0.376 0.339 0.357 61.51 - - -
@@
 PVSNet [101] 0.337 0.315 0.326 59.11 35.51 76.57 82.62
-MVSFormer [82] 0.327 0.251 0.289 66.37 40.87 - -
 IS-MVSNet [224] 0.351 0.359 0.355 62.82 34.87 73.33 83.15
@@
 GoMVS [213] 0.347 0.227 0.287 66.64 43.07 79.16 85.91
+MVSFormer++ [93] 0.309 0.252 0.281 67.03 41.70 - 82.99
 Iterative update PatchmatchNet [39] 0.427 0.277 0.352 53.15 32.31 64.21 73.12
@@
 IGEV-MVS [118] 0.331 0.326 0.324 - - - - -
-6.1 Supervised Online MVS with Depth Estimation
-Typically, online MVS methods are trained on ScanNet [26] and then evaluated on ScanNet [26] and 7-Scenes [179].
-We summarize the quantitative results of online MVS methods in Tab.
-I.
-By incorporating temporal information in depth estimation,  [60, 188, 30] achieve better performance than MVDepthNet [25].
-However, this usually increase the network complexity.
-Instead of temporal fusion, SimpleRecon [34] injects cheaply available metadata into the cost volume and achieves state-of-the-art performance on both ScanNet and 7-Scenes.
-6.2 Supervised Offline MVS with Depth Estimation
-Offline MVS methods are usually trained on DTU [21] and then evaluated on DTU [21], Tanks and Temples [22] and ETH3D [23].
-Recently, many methods further finetune the DTU-pretrained model on BlendedMVS [24] before evaluating on Tanks and Temples and ETH3D since it contains large-scale scenes.
-6.2.1 Benchmark Performance
-Tab.
-III summarizes quantitative results of offline MVS methods.
-Compared with those methods that use direct 3D CNN for regularization, the methods that use RNN and coarse-to-fine regularization perform much better.
-Since RNN based methods have bad time efficiency, we can find that coarse-to-fine methods become main stream of the community because of their impressive performance and high efficiency in both memory and run-time.
-In addition, iterative methods based on traditional PatchMatch [39, 102] or RAFT [40, 104, 103] also become popular since they achieve comparable performance as state-of-the-art methods with more lightweight structure.
-For benchmarks, most learning-based methods perform explicitly better than traditional methods on DTU and Tanks and Temples.
-However, on ETH3D, traditional methods [189, 123] perform better than all existing learning-based methods.
-We conjecture that the strong viewpoint variations and large textureless regions in human-made environments still make learning-based methods struggle.
-Therefore, it is valuable to pay more attention to ETH3D to evaluate the robustness of learning-based methods in real-world scenes.
-6.2.2 Memory Consumption and Run-time
-Low run-time and memory consumption are crucial in most industrial applications with limited computational power and storage, e.g., robotics and AR/VR.
-Because of the low-resolution input and the simplicity of network structures, online MVS methods [25, 30, 34] usually achieve high efficiency in both memory and run-time.
-In contrast, offline MVS methods focus on high-resolution images and usually use computationally expensive network modules, e.g., 3D convolutions, to improve reconstruction quality.
-These explicitly increase run-time and GPU memory.
-Recently, many researchers tried to improve the efficiency while maintaining the reconstruction accuracy as other offline MVS methods.
-We compare the efficiency of state-of-the-art offline MVS methods [31, 39, 104, 40, 100, 79, 111, 76, 86] in Fig.
-5.
-All these methods share similar coarse-to-fine structures to improve efficiency.
-However, we observe that some methods [31, 79, 111, 76] still require lots of computation resource and are not efficient in either memory or run-time.
-To further reduce run-time and memory, [39, 104, 40, 100] carefully simplify the architecture, e.g., replacing costly 3D convolution with 2D convolution.
-Moreover, these methods still achieve competitive performance when compared with the top-performing methods.
-Therefore, it is promising to further improve both the performance and efficiency for practical applications.
-Figure 5: GPU memory consumption and run-time of state-of-the-art offline methods [31, 39, 104, 40, 100, 79, 111, 76, 86] on DTU dataset [21].
-The image resolution and the number of images are set to $1600\times 1152$ and 5 respectively.
-For fair comparison, experiments are done on one workstation with a NVIDIA 2080 Ti GPU.
-6.3 Unsupervised MVS with Depth Estimation
-We summarize the results of unsupervised MVS methods in Tab.
-II.
-Based on the conventional photometric consistency loss (Sec.
-4.1), unsupervised MVS methods may have difficulties with many challenging situations, e.g., varying lighting conditions or occlusions, and suffer from downgraded reconstruction results.
-Therefore, many methods propose to employ different strategies to improve the robustness, e.g., segmentation consistency [41], optical flow consistency [43], neural rendering consistency [42], pseudo-label generation with geometric filtering [128, 44], multi-iteration training [128, 44] and featuremetric loss [44].
-It is inspiring that KD-MVS [44] already outperforms its backbone, CasMVSNet [31], and many other fully supervised methods.
-However, since unsupervised methods mainly test on relatively simple scenes [21, 22], their scalability to more complex and large-scale scenes, e.g., ETH3D [23], remains a question.
-This good generalization capability across various scenes and datasets is crucial for practical applications.
-Therefore, evaluating unsupervised MVS methods on various scenes is an important task for future research.
-Most unsupervised methods use simple MVS backbones [15, 31, 75] and mainly focus on the training strategy.
-Employing state-of-the-art architectures is also a potential direction that can further boost the performance.
-6.4 Learning-based MVS without Depth Estimation
-6.4.1 Voxel-based Methods
-We summarize the quantitative results on ScanNet [26] in Tab.
-IV.
-Since the voxel representation increases the computation overhead explicitly when the scale of scene increases, voxel-based methods mainly focus on indoor scenes, e.g., ScanNet [26].
-We find that depth map-based MVS methods, e.g., SimpleRecon [34], fuse depth maps with traditional TSDF fusion [65] and can achieve comparable performance as these voxel-based methods with lower complexity.
-This further shows the advantages of depth map-based methods over voxel-based methods.
-TABLE IV: Quantitative results of voxel-based MVS methods on ScanNet [26].
-Methods Prec. $\uparrow$ Recall $\uparrow$ F-score $\uparrow$
-Atlas [46] 0.675 0.605 0.636
-NeuralRecon [47] 0.630 0.612 0.619
-TransformerFusion [134] 0.728 0.600 0.655
-VoRTX [135] 0.767 0.651 0.703
-TABLE V: Quantitative results of NeRF-based MVS methods on DTU [21].
-Methods VolSDF [48] NeuS [49] NeuralWarp [129] HF-NeuS [190] RegSDF [191] PET-NeuS [192]
-Overall $\downarrow$ 0.86 0.87 0.68 0.77 0.72 0.71
-Methods Geo-NeuS [50] Voxurf [193] NeuS2 [148] PermutoSDF [150] Neuralangelo [149] UniSDF [151]
-Overall $\downarrow$ 0.51 0.72 0.70 0.68 0.61 0.64
-6.4.2 NeRF-based Methods
-The quantitative results of nerf-based methods on DTU dataset [21] are summarized in Tab.
-V.
-The most outstanding features of NeuS [49] and VolSDF [48] are that they are self-supervised, i.e., without depth supervision, and produce smooth and complete mesh surfaces.
-However, many recent methods [50, 146, 147] find that incorporating more explicit geometric supervision explicitly improves the reconstruction quality.
-Compared with depth map-based MVS methods [25, 15], the main drawback of NeRF-based methods is that they are mostly per-scene optimization problems and need lots of time to train the model on each new scene.
-Recently, some methods [152, 153, 154] try to make the pipeline generalizable so that there is no need to train the model for new scenes.
-However, they perform worse than depth map-based methods [15].
-Another way to improve training time efficiency is to employ hash grids [138], which also improves surface details.
-However, these methods [148, 149, 150, 151] need lots of GPU memory and space to train and store the large model.
-In addition, the scalability of NeRF-based methods remains a problem when attempting to apply on large-scale scenes [23].
-6.4.3 3D Gaussian Splatting-based Methods
-Because of its efficient rasterization, 3D Gaussian Splatting [160] becomes popular in novel view synthesis in the last half year.
-However, there are many challenges to use reconstruct surfaces with 3DGS.
-For example, the 3D Gaussians do not correspond well to the actual surface of the scene since the 3D Gaussian has a 3D covariance matrix [51].
-Moreover, the surface may contain noisy undulations [51].
-However, it is worthwhile to further explore in this direction since 3DGS is much faster to train than those NeRF-based MVS methods.
-Moreover, instead of directly extracting the surface from 3D space [51], rendering accurate depth maps and then fuse them like depth map-based MVS methods is a promising direction [52, 53].
-Additionally, generalizable feed-forward 3DGS methods [194, 195] are interesting since they are proven to improve depth accuracy with rendering loss and do not be trained for each scene.
-6.4.4 Large Feed-forward Methods
-As a new trend of 3D reconstruction, large feed-forward methods demonstrate impressive performance in challenging settings that other MVS methods cannot handle, e.g., single image input and un-posed sparse images.
-Employing a large model that is able to get scaled up, these methods manage to learn strong priors of multi-view geometry from large datasets [29].
-Since these large feed-forward methods [54, 174, 55] usually do not enforce geometric constraint explicitly, i.e., epipolar geometry, the reconstructed geometry is by default placed in a local coordinate system without absolute scales, making it not feasible for applications requiring precise measurement.
-Also, these methods reconstruct the target under a globally shared representation, while depth-based MVS is typically performed without global data dependency, making conventional MVS still more suitable for large-scale reconstruction and parallel processing.
-The massive and costly computation of large feed-forward methods also limit the feasibility of applications.
-Benefited from the deep prior rooted in the large datasets, large feed-forwards methods are able to handle some of the typical corner cases, e.g., non-Lambertian surfaces.
-However it also makes their generalization ability questionable since they learn priors of multi-view geometry completely from the training dataset, while the depth map-based methods utilize feature matching via plane sweep algorithm from traditional methods that are more explainable.
-6.5 Future Research Directions
-6.5.1 Datasets and Benchmarks
-For learning-based MVS methods, ScanNet [26] and DTU [21] are two main training datasets, while ScanNet [26], DTU [21], Tanks & Temples [22] and ETH3D [23] are the main evaluation benchmarks.
-For training, the scene scale of ScanNet and DTU is relatively small (room-scale for ScanNet and object-scale for DTU) and their quality is not satisfactory.
-For example, the camera calibration of ScanNet is not very accurate.
-For DTU, the ground truth depth maps are rendered from mesh, which contain some outliers and holes.
-This is because the mesh is reconstructed from the sparse ground truth point cloud with Screened Poisson surface reconstruction [196] and thus contains incomplete regions [182].
-Therefore, improving the scalability and quality of training datasets is an important research direction.
-Recently, there are many researchers trying to solve this problem.
-TartanAir dataset [197] is a large-scale synthetic dataset that is collected in photo-realistic simulation environments.
-BlendedMVS [24] introduces more large-scale scenes and improves the performance of many MVS methods on real-world scenes, shown in Tab.
-III.
-ArKitScenes [198] consists of 5,048 RGB-D sequences, which is more than three times the size of the current largest available indoor dataset, ScanNet.
-As an extension of ScanNet, ScanNet++ [199] is a large-scale dataset that captures high-quality geometry and color of indoor scenes with high-end laser scanner, DSLR camera, and RGB-D streams from an iPhone.
-Motivated by ImageNet [200] that drives a remarkable trend of learning from large-scale data in 2D visual tasks, Yu et al.
-[201] propose MVImgNet, a large-scale dataset of multi-view images collected by shooting videos of real-world objects.
-MVImgNet demonstrates the potential of various 3D visual tasks, including radiance field reconstruction, multi-view stereo, and view-consistent image understanding.
-Similarly, Objaverse [29] is a large dataset of objects with 800K+ (and growing) 3D models with descriptive captions, tags, and animations.
-Both MVImgNet [200] and Objaverse [29] can be used to train large feed-forward models, e.g., LRM [54], for reconstruction, since these models need massive data to learn the strong prior.
-However, both of them mainly focus on object-level scenes.
-Large datasets for larger-scale scenes, e.g., rooms, are valuable to be explored.
-For evaluation, the main benchmarks [26, 21, 22, 23] are all introduced before 2018.
-Some of these benchmarks are currently saturated and the methods are difficult to further improve the performance on them.
-Therefore, it is meaningful to introduce new benchmarks to evaluate the robustness and performance of learning-based methods in challenging real-world scenes.
-For example, LaMAR [202] is a recent large-scale dataset captured using multiple modern AR devices in diverse environments.
-It contains challenging short-term appearance and structural changes and high quality LiDAR point cloud ground truth.
-These make LaMAR a potential benchmark for MVS.
-6.5.2 View Selection
-As discussed in Sec.
-2.2, view selection is crucial for triangulation quality.
-Picking neighboring views that are suitable for triangulation can not only improve the reconstruction accuracy but also reduce useless computation for the bad views, e.g., views with strong occlusions.
-However, view selection is often overlooked and not well studied.
-For example, all offline depth map-based MVS methods [31, 39, 79, 100] follow MVSNet [15] and use the same simple heuristic strategy to compute scores for neighboring views and sort them.
-Online depth map-based MVS methods [60, 30, 34] and voxel-based methods [47] adopt heuristic strategies to choose views with enough pose-distance.
-Though it is intractable to incorporate non-differentiable view selection into deep learning, it is worth exploring new view selection strategies since it may improve the reconstruction without changing the model design.
-Current view selection strategies simply pick the same set of neighboring views for all the pixels in the reference image, and it is probable that different reference pixels have different optimal choices of neighboring views.
-The learned pixel-wise view weight [94, 39, 40] is one solution for this since it weights the source views differently for each pixel based on its visibility across neighboring views.
-6.5.3 Depth Fusion
-As discussed in Sec.
-2.4, both online and offline MVS methods with plane-sweep use traditional TSDF fusion or depth filtering methods to reconstruct mesh or point cloud from the estimated depth maps.
-However, TSDF fusion [65, 66] is not robust enough to the outliers in the depth maps and may have memory issues due to the dense volumetric representation.
-Depth filtering following Galliani et al.
-[18] introduces lots of hyper-parameters.
-Researchers may carefully finetune these hyper-parameters for each scene since it may have a great impact on the evaluation metrics.
-Therefore, it is valuable to improve the depth fusion step so that the reconstruction quality can be further improved.
-Recently, there have been some learning-based depth map fusion methods [203, 204].
-However, they are limited to small scenes due to the dense volume representation and large memory consumption.
-6.5.4 Features
-One not well-studied topic is what kind of feature extractors are suitable for MVS, as mentioned in Sec.
-3.1.
-So far, most of the offline depth map-based methods [15, 31, 78] mainly apply simple 2D CNN and FPN [72] structure as a feature extractor.
-To improve the receptive fields flexibly, Deformation Convolution [205] is used  [78, 77, 39, 79]).
-The attention mechanism is also applied in feature learning, where works such as  [91, 79, 77] use intra-attention or inter-attention to capture long-range feature dependencies.
-Furthermore, MVSFormer [82] digs deeper into using patch-wise ViT [177] as a feature extractor in MVS.
-The insight is that ViT works better to formulate global feature correlations, and FPN can learn detailed ones.
-In contrast, online depth map-based methods [30, 34] usually adopt efficient pretrained backbones [73, 70, 74].
-Recently, large feed-forward methods [54, 55, 166] adopt large pretrained feature encoders, e.g., ViT [177], DINO [83], to learn strong prior from large-amount of data.
-However, using a powerful feature extraction network usually increases the computation overhead and reduces the efficiency.
-Finding a good balance between performance and efficiency is an interesting direction, e.g., distilling the large ViT into a relatively small transformer [206].
-6.5.5 Real-time and Memory Efficient Models
-Though reconstruction accuracy is considered the most important factor in MVS, it is also critical to have models that can run in near real-time and with low memory, e.g., AR/VR and autonomous driving.
-Comparatively, depth map-based methods are most suitable for achieving efficiency because of their conciseness, flexibility, and scalability.
-Though online MVS methods can achieve real-time estimation with images of low resolutions, they may have issues with memory consumption since many large backbones are usually used for feature extraction [30, 34].
-In addition, the efficiency in both run-time and memory will drop when image resolution increases.
-For offline MVS methods, as discussed in Sec.
-6.2.2, some recent methods try to improve efficiency with carefully simplified network architectures.
-However, this usually limits the performance when compared with state-of-the-art methods.
-There are many other directions for further improvement.
-For example, using model compression techniques and knowledge distillation are potential directions to not only keep the good performance and also improve efficiency.
-6.5.6 Prior Assistance
-MVS mainly relies on assessing local photometric consistency to find the optimal matching across reference and source images.
-Accordingly, it usually encounters difficulties when estimating the geometry for regions where the photometric measurement becomes unreliable or invalid, e.g., textureless areas and non-Lambertian surfaces, which are common in human-made scenes [23].
-Therefore, using prior information to guide the MVS algorithm in these challenging regions is a promising research direction.
-We elaborate 3 typical examples of prior assistance as follows.
-Surface Normal: As a non-local representation of the geometry compared with the depth map, surface normal has been proven effective in improving reconstruction performance in recent works.
-To enforce the constraints of normal maps in depth estimation, Kusupati et al.
-[207] integrate a multi-view normal estimation network into the MVS pipeline.
-Long et al.
-[208] introduce a Combined Normal Map, estimated with PlaneCNN [209] to enforce the normal consistency on the depth map.
-Liu et al.
-[210] apply graph-based depth map optimization as post-processing.
-Since monocular normal network [211] is trained on large-scale dataset and can provide high-quality priors, [146, 187] use monocular normal priors to improve the surface accuracy.
-Shape Prior: For indoor scenes where common textureless areas, e.g., walls, planes are suitable choices of the geometric primitives and are exploited in traditional methods [20].
-Recently, PlaneMVS [212] and PlanarRecon [213] explicitly estimate plane parameters for depth refinement or holistic reconstruction.
-In addition, predicting object-level attributes simultaneously when estimating depth [214] is also a feasible solution for specific applications like urban modeling.
-Semantic Segmentation: Intuitively, points assigned with the same semantic labels may be more likely to lie on the same 3D plane.
-Some attempts [215, 216] employ a rule-based protocol to utilize semantic segmentation for better matching and depth quality.
-Manhattan-SDF [217] relies on 2D segmentation results to enforce Manhattan world priors to handle low-textured planar regions.
-Similarly, Shvets et al.
-[218] also employ SAM (Segment Anything Model) encoder to generate semantic features to enhance MVS.
-7 Conclusion
-In this survey, we have provided a comprehensive review of the learning-based MVS methods, which are categorized into: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
-Particularly, we have devoted significant attention to depth map-based methods because of their conciseness, flexibility and scalability.
-We explain key aspects such as datasets utilized, general working pipelines, and algorithmic intricacies.
-Furthermore, we have summarized and compared the quantitative performance of these methods across popular benchmarks, providing valuable insights into their efficacy and applicability.
-Finally, our discourse extends to an in-depth exploration of potential research directions for the future of MVS, highlighting avenues for further investigation and innovation in the field.
+DiffMVS [119] 0.318 0.297 0.308 63.39 39.69 74.86 82.10
+CasDiffMVS [119] 0.310 0.286 0.298 65.87 41.81 76.76 85.11
+TABLE VI: Quantitative results of NeRF-based and 3D Gaussian Splatting-based MVS methods on DTU [22].
+Methods VolSDF [154] NeuS [49] NeuralWarp [137] RegSDF [228] PET-NeuS [229]
+Overall $\downarrow$ 0.86 0.87 0.68 0.72 0.71
+Methods Geo-NeuS [160] NeuS2 [156] PermutoSDF [157] Neuralangelo [50] UniSDF [51]
+Overall $\downarrow$ 0.51 0.70 0.68 0.61 0.64
+Methods SuGaR [55] 2DGS [56] GaussianSurfels [230] GOF [193] PGSR [57]
+Overall $\downarrow$ 1.33 0.80 0.88 0.74 0.49